### v1

# Improved energy barrier in higher-dimensional hypergraph product codes

Abstract
Single-shot error correction outperform conventional approaches by requiring only one round of stabilizer measurements for decoding, even in the presence of measurement errors. This capability relates to the confinement property of codes, which provides an energy barrier lower bound. Earlier research by Quintavalle et al. Quintavalle et al. (2021) established a confinement property for higher-dimensional hypergraph product (HHGP) codes, resulting in an energy barrier lower bound for these codes. In this work, we show an improved energy barrier lower bound for HHGP codes with low-density parity-check (LDPC) properties. This bound is derived by analyzing the structure of the logical operators. Our bound exceeds results derived from confinement alone, and unlike standard hypergraph product codes, these higher dimensional variants can possess macroscopic energy barriers even when the underlying classical codes lack this property. Specifically, our analysis shows that the energy barrier of LDPC HHGP codes is lower bounded by the distance of the underlying classical codes. This lower bound is tight if the underlying classical codes exhibit system size-dependent distances but constant energy barriers, like 3D and 4D toric codes.
I Introduction
Quantum error correction is crucial for fault tolerant quantum computing, yet it introduces significant overhead. Within the stabilizer code framework Gottesman (1997) implemented through the quantum circuit model Nielsen and Chuang (2001), a substantial portion of overhead stems from the stabilizer measurements required to detect and identify errors. Specifically, along the quantum circuit, stabilizer measurement and decoding must be performed at regular intervals to maintain the integrity of the code state. Moreover, to overcome measurement errors, conventional protocols require multiple rounds of stabilizer measurements for each decoding process, with the number of rounds typically scaling with the code distance Shor (1996); DiVincenzo and Shor (1996).
In response to these challenges, single-shot error correction emerges as a promising alternative by requiring only one round of stabilizer measurements per decoding process Bombín (2015). In the presence of measurement errors, although this approach still leaves a small residual error after decoding, the residual error remains bounded. Such controlled error propagation is sufficient for fault-tolerant quantum computation.
Single-shot error correction requires quantum codes with specialized structural properties. Bombín first demonstrated that codes with confinement property enable single-shot correction Bombín (2015). Later, Campbell proposed that good soundness is also sufficient for single-shot error correction Campbell (2019). He also showed that one can realize single-shot error correction for any code by measuring a carefully chosen set of stabilizers. Further advancing the theory, Quintavalle et al. provided a general definition of confinement, and proved that confinement is sufficient for single-shot decoding for adversarial errors Quintavalle et al. (2021).
Codes demonstrated to support single-shot error correction include the 3D gauge color code Bombín (2015); Brown et al. (2016a), 3D hypergraph product codes (only for either X or Z noise, depending on code structure)Kubica and Preskill (2019); Vasmer et al. (2021); Quintavalle et al. (2021); Higgott and Breuckmann (2023), 4D hypergraph product codesBreuckmann et al. (2017); Campbell (2019); Zeng and Pryadko (2019), 4D hyperbolic codes Breuckmann and Londe (2022), quantum expander codes Fawzi et al. (2020), and quantum Tanner codes Gu et al. (2024).
Among these codes, HHGP codes Zeng and Pryadko (2019) represent an important family of quantum codes that extend the original hypergraph product construction Tillich and Zemor (2014) into higher dimensions. While the standard hypergraph product combines only two classical codes to form a quantum code, this higher-dimensional approach incorporates additional classical codes. These additional codes enable the presence of meta-checks for X and Z stabilizers, effectively validating the stabilizers themselves, thereby providing an intuitive foundation for the single-shot error correction capability these codes exhibit.
Single-shot error correction shares fundamental connections with self-correcting Bombín (2015); Hong et al. (2025), which necessitates that the system preferentially remains within the code space at finite temperature Brown et al. (2016b); Terhal (2015). A thorough analysis of self-correcting requires evaluation of free energy dynamics. While entropic effects play a crucial role in determining the system’s stability Plischke and Bergersen (1994); Bravyi and Haah (2013, 2011), macroscopic energy barriers emerge as a necessary condition. The term “macroscopic” means these barriers scale with system size.
Previously, our author established a lower bound for the energy barrier of LDPC hypergraph product codes, which is directly determined by the energy barriers of their underlying classical codes Zhao et al. (2025). In this scenario, constructing quantum codes with macroscopic energy barriers therefore necessitates classical codes with the same property. However, determining energy barriers for classical codes remains challenging, raising the question of how to construct quantum codes that exhibit macroscopic energy barriers when their classical constituents do not.
The HHGP codes indeed meet this criteria. We will show that the energy barrier of LDPC HHGP codes is lower bounded not only by the energy barriers of the underlying classical codes (as in standard hypergraph products), but also by the distances of these underlying classical codes. Thus to construct quantum codes with macroscopic energy barrier, one can just use classical codes that have large distance.
Notably, Quintavalle et al. demonstrated that all 3D hypergraph product codes exhibit confinement in X-components (under their experimental setup), making them single-shot correctable against adversarial Z-type errors Quintavalle et al. (2021). This confinement induces a lower bound on the energy barrier. However, we demonstrate that this confinement-derived lower bound is not optimal. Our approach, based on analyzing logical operator structures, provides a more accurate characterization of the energy landscape of those LDPC HHGP codes.
To guide the reader, our paper is structured as follows. In Sec. II, we review the concepts of soundness, confinement, energy barrier, expander code and HHGP codes. We then show formally how confinement induces a lower bound of energy barrier in Sec. III, and use the confinement result in Ref. Quintavalle et al. (2021) to give an energy barrier lower bound for 3D hypergraph product codes. In Sec. IV, we provide an energy barrier lower bound for tensor product codes, as formally stated in Lemma 11. This lower bound serves as a foundation for proving the energy barrier lower bound of LDPC HHGP codes. Finally, by analyzing the structure of logical operators of HHGP codes and applying Lemma 11, we derive improved energy barrier lower bounds for 3D and 4D hypergraph product codes in Sec. V and Sec. VI, respectively. These bounds, formally presented in Theorem 16 and Theorem 17, are shown to relate directly to the distances of the underlying classical codes.
II Preliminaries
II.1 Quantum LDPC codes
Stabilizer codes, conceptualized by Daniel Gottesman in 1996 Gottesman (1997), encode quantum information in a protected subspace of a larger Hilbert space. Specifically, these codes store quantum states within the $+1$ co-eigenspace of a set of $m$ commuting Pauli operators, $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{m}\}$, defined on an $n$-qubit Hilbert space $\mathcal{H}_{2}^{\otimes n}$. Each operator $S_{i}\in\{I,X,Y,Z\}^{\otimes n}$ consists of a tensor product of the Pauli operators, with the requirement that all these operators commute with each other. $\mathcal{S}$ excludes the operator $-I$ to ensure a non-trivial code space. Logical operators are defined as the elements of $\mathcal{C}(S)\backslash S$.
Calderbank-Shor-Steane (CSS) codes Calderbank and Shor (1996); Steane (1996) are an important family of stabilizer codes with special structural properties. Their stabilizer generators are exclusively either X-type or Z-type Pauli operators. This separation allows all X-type stabilizers to be represented by a parity check matrix $H_{X}$, while all Z-type stabilizers are represented by a parity check matrix $H_{Z}$. Consequently, the parity check matrix for the quantum CSS code takes the form
$H=\left(\begin{array}[]{cc}0&H_{Z}\\ H_{X}&0\end{array}\right),$ (3)
with the condition $H_{X}H_{Z}^{T}=0$ ensures that all stabilizers commute with each other.
Quantum LDPC codes are stabilizer codes with sparse parity-check matrices (see Babar et al. (2015); Breuckmann and Eberhardt (2021a) for good reviews). The sparsity is defined by two parameters: $w_{c}$, the maximum weight amongst all stabilizer generators, and $w_{q}$, the maximum number of stabilizer generators associated with any single qubit. Formally, a code is classified as LDPC if both $w_{c}$ and $w_{q}$ remain constant as the code size increases, specifically $w_{c},w_{q}=O(1)$.
Quantum LDPC codes are promising because they often provide lower overhead and can be designed for fault-tolerance Gottesman (2014). Recent research has led to several families of quantum LDPC codes with enhanced parameters Freedman et al. (2002); Evra et al. ; Kaufman and Tessler (2021); Hastings et al. (2021); Panteleev and Kalachev (2022a), with some efforts even achieving good quantum LDPC codes Breuckmann and Eberhardt (2021b); Panteleev and Kalachev (2022b); Leverrier and Zemor (2022); Dinur et al. (2023).
Various methods exist for constructing quantum LDPC codes. From a mathematical perspective, the hypergraph product construction Tillich and Zemor (2014); Zeng and Pryadko (2019) represents a well-established approach, providing a systematic way to construct quantum codes from any classical codes. Alternative methods such as two-block group-algebra (2BGA) codes Wang et al. (2023); Lin and Pryadko (2024) offer different advantages, like suitability for two-dimensional layouts Bravyi et al. (2024), making them useful for practical quantum computing architectures.
II.2 Energy barrier of codes
Following Ref. Bravyi and Terhal (2009), we present a formal definition of the energy barrier for stabilizer codes. Let $\mathcal{C}$ be the code subspace of a stabilizer code, defined in terms of the stabilizer group $S$. The code subspace can be viewed as the ground state subspace of a Hamiltonian of the form $\hat{H}=\sum_{i=1}^{m}(I-s_{i})/2$, where $\{s_{1},\ldots,s_{m}\}\subset S$ is a set of generators. Note that the energy barrier depends on the choice of the generating set.
For an operator $P$ in the Pauli group $\mathcal{P}$, the energy of the state $P|\psi\rangle$ is given by $\bra{\psi}P^{\dagger}\hat{H}P\ket{\psi}=\epsilon(P)$. Here |$\ket{\psi}$ is any ground state with $\bra{\psi}\hat{H}\ket{\psi}=0$, and $\epsilon(P)$ is the energy cost of $P$. This is the number of $s_{i}$s that anti-commute with $P$, i.e., $\epsilon(P)=|\{i:s_{i}P=-Ps_{i}\}|$. Equivalently, one may define the energy barrier in terms of the parity check matrix $H$ of the stabilizer code. Let $v(P)$ be the binary (bit-string) representation of a Pauli $P$, then
$\epsilon(P)=\text{wt}(Hv(P)).$ (4)
A sequence $P_{0},P_{1},\ldots,P_{t}$ from the Pauli group $\mathcal{P}$ forms a path from $P_{0}$ to $P_{t}$ if for each index $i$, the operators $P_{i}$ and $P_{i+1}$ differ at no more than one qubit, we call it local error condition. The notation $w(P_{0},P_{t})$ represents the collection of all such paths from $P_{0}$ to $P_{t}$. For $r\in w(P_{0},P_{t})$, $\epsilon_{\max}(r)$ denotes the highest energy along path $r$, i.e., $\epsilon_{\max}(r)=\max_{P_{i}\in r}\epsilon(P_{i})$, as the energy barrier of $E$ along the path $r$.
The minimum energy associated with a Pauli $P$ is the smallest value of $\epsilon_{\text{max}}$ across all possible paths from 00 to $P$. This is the energy barrier of $P$, denoted as $\Delta(P)$,
$\Delta(P)=\min_{r\in w(I,P)}\epsilon_{\max}(r).$ (5)
The energy barrier of the quantum code is the minimum energy barrier over the set of nontrivial logical operators.
Definition 1.
Let $S$ be a stabilizer group and $L(S)$ be the set of nontrivial logical operators. The energy barrier is
$\Delta(H):=\min_{\ell\in L(S)}\Delta(\ell).$ (6)
This is the smallest energy the environment has to overcome to enact a logical operation on the encoded qubit. We can similarly define the energy barrier of classical codes by only considering the path formed by Pauli-Xs. We shall denote this energy barrier also as $\Delta(H)$, where $H$ in this case is the parity check matrix of the classical code.
There are two main obstacles for computing the energy barrier of quantum codes. First, numerous paths exist for implementing any logical operator, creating a vast solution space to explore. Second, in quantum codes, any stabilizer can be multiplied with a logical operator to produce an equivalent logical operator, while for those equivalent logical operators, the energy barriers may differ. That is to say, for a logical operator $L$ and stabilizer $S$, usually $\Delta(L)\neq\Delta(LS)$. This complicates the analysis as one must consider the full equivalence class of logical operators.
Quantum LDPC codes offer significant advantages for energy barrier computation by reducing analytical complexity. As shown in Ref. Zhao et al. (2025), in LDPC codes, two logical operators that are equivalent under stabilizers have identical energy barriers, provided at least one of their energy barriers equals or exceeds $w_{c}w_{q}$, the product of sparsity parameters. With $w_{c}w_{q}=O(1)$, this property yields a powerful simplification: if any non-trivial logical operator has an energy barrier of $\Omega(1)$, all equivalent logical operators share the same energy barrier.
Therefore, to determine the energy barriers of quantum LDPC codes, one needs only analyze a fixed complete set of logical operators. Once the energy barrier is established for this complete set, the energy barrier of the code follows directly from these results.
II.3 Soundness
Soundness originated in the study of locally testable codes (LTCs) Blum et al. (1990); Arora (1994); Rubinfeld and Sudan (1996); Friedl and Sudan (1995). LTCs are a class of error-correcting codes that allow efficient verification of whether a string is a valid codeword or is far from the code space. The soundness of a code, denoted by $R(\delta)$, measures the likelihood that a randomly selected local constraint is violated by a string that is close to $\delta n>0$ from the code space, where $n$ is the codeword length. The concept of locally testable has deeply influenced fields like PCP, combinatorial optimization, property testing, program verification, and cryptography.
Aharonov and Eldar introduced quantum locally testable codes (QLTCs) Aharonov and Eldar (2015), with interest surging after Eldar and Harrow showed Eldar and Harrow (2017) that QLTCs with constant soundness, locality, and relative distance could construct Hamiltonians lacking low-energy trivial states—addressing the NLTS conjecture Freedman and Hastings (2014). QLTCs are closely related to the quantum PCP conjecture Aharonov et al. (2013), a fundamental problem in quantum complexity theory related to the quantum analog of the classical PCP theorem.
Roughly speaking, the soundness property ensures that high-weight errors produce high-weight syndromes. For stabilizer codes, soundness can be characterized as follows. Consider a code $C$ defined on $n$ qubits, determined by a generating set $\mathcal{G}$ where each generator has support over $w$ qubits (representing the stabilizer size). For any error $E$ acting on the code with a syndrome weight of at least $\delta n$, a randomly chosen generator $g\in\mathcal{G}$ fails to commute with $E$ with probability at least $R(\delta)$, where $R(\delta)$ is the soundness function Aharonov and Eldar (2015). This property ensures that errors can be detected probabilistically by examining only a limited number of stabilizers.
To analysis single-shot property of quantum codes, Campbell provided the following alternative definition of soundness Campbell (2019).
Definition 2 (Soundness, Definition 3 in Campbell (2019)).
Let $t$ be an integer and $f:\mathbb{Z}\rightarrow\mathbb{R}$ be a function termed the soundness function, where $f(0)=$ 0. A set of Pauli stabilizers $\mathcal{M}$ is called $(t,f)$-sound if, for every Pauli error $E$ with syndrome weight $|\sigma(E)|=x<t$, there exists an $E^{\star}$ such that $\sigma\left(E^{\star}\right)=\sigma(E)$ and the weight wt $\left(E^{\star}\right)\leqslant f(x)$.
In essence, this definition requires that low-weight syndromes must correspond to low-weight errors that could produce them. This formulation represents the contrapositive of definition of locally testable codes, with the key distinction being its focus on scenarios where $|\sigma(E)|=x<t$. Under this definition, good soundness is characterized by function $f(x)$, which is a monotonically increasing polynomial function of X that remains independent of the size of the check set.
Campbell demonstrated that codes possessing good soundness naturally exhibit single-shot error correction capabilities, with the specific performance determined by the soundness parameters $t,f$ and the code’s parameters $n,k,d$ Campbell (2019). Moreover, he proved that good soundness in LDPC codes necessarily implies the existence of a macroscopic energy barrier, as formalized in the following lemma:
Lemma 3 (Lemma 3 in Campbell (2019)).
Consider a $\left[\left[n,k,d\right]\right]$ quantum code with checks $\mathcal{M}$ that is $(t,f)$-sound and where all qubits are involved in no more than $w_{c}$ checks. It follows that the energy barrier is at least $f^{-1}(c)$ where $c=\min\left[(t-1)/w_{c},\left(d-1\right)/2\right]$ and $f^{-1}$ is the inverse of the soundness function.
It’s important to note that the converse statement, any LDPC check family with a macroscopic energy barrier necessarily possesses good soundness, does not hold. The expander code provides a clear counterexample: despite exhibiting bad soundness due to its lack of check redundancy Ben-Sasson et al. (2010), it has been proven to support single-shot error correction with the small-set flip decoder Fawzi et al. (2020, 2018a). In this context, check redundancy refers to the ratio between the check set size (encompassing all checks, including dependent ones) and the check group size (comprising only independent checks).
The soundness of a stabilizer code depends specifically on the selection of the stabilizer set, not merely on the stabilizer generators. Campbell also demonstrated that for any stabilizer code, it is possible to strategically choose a stabilizer set that endows the code with good soundness properties Campbell (2019). The price is that the size of some checks could be excessively large. LDPC property may be lost due to the existence of high-weight checks.
To maintain the LDPC characteristic, Campbell discovered an elegant solution for achieving single-shot properties: applying the hypergraph product construction twice Campbell (2019). The resulting structures naturally possess meta-checks, a direct consequence of the iterative hypergraph product process. Through this approach, Campbell effectively combined soundness and LDPC properties by strategically leveraging check redundancy. The following lemma show this formally.
Lemma 4 (Lemma 5 and Lemma 6 in Campbell (2019)).
Let $C_{0}\xrightarrow{\partial_{0}}C_{1}$ be a chain complex. Applying the hypergraph product, we obtain a new chain complex $\tilde{C}_{-1}\xrightarrow{\tilde{\delta}_{-1}}\quad\tilde{C}_{0}\xrightarrow{\tilde{\delta}_{0}}\quad\tilde{C}_{1}$, where the maps $\tilde{\delta}_{0}^{T}$ and $\tilde{\delta}_{-1}$ are $(t,f)$-sound with $f(x)=x^{2}/4$ and $t=\min\left(d_{0},d_{0}^{T}\right)$, where $d_{0},d_{0}^{T}$ are the distances of codes defined by $\delta_{0},\delta_{0}^{T}$, respectively. Applying the hypergraph products again, we obtain a new chain complex $\breve{C}_{-2}\xrightarrow{\breve{\delta}_{-2}}\quad\breve{C}_{-1}\xrightarrow{\breve{\delta}_{-1}}\quad\breve{C}_{0}\xrightarrow{\breve{\delta}_{0}}\quad\breve{C}_{1}\xrightarrow{\breve{\delta}_{1}}\quad\breve{C}_{2}$, where the maps $\breve{\delta}_{0}$ and $\breve{\delta}_{-1}^{T}$ are $(t,g)$-sound with soundness function $g(x)=x^{3}/4$, and $t=\min\left(d_{0},d_{0}^{T}\right)$.
For a 4D hypergraph product code that constructed with four identical classical codes defined by the same parity check matrix $\delta$ with distance $d$ ($\delta^{T}$ with distance $d^{T}$), according to Lemma 3, the energy barrier for logical operators of the resulting quantum code then has a lower bound of $O\left({(\min(d,d^{T}))}^{\frac{1}{3}}\right)$. Later, we will demonstrate that the optimal lower bound for this energy barrier scales as $O(\min(d,d^{T}))$ in the given scenario.
II.4 Confinement
Confinement is another robustness measure of quantum codes. Bombín originally introduced the concept of confinement as a mechanism that allows quantum codes to achieve single-shot error correction Bombín (2015). However, in this discussion, we use a more developed definition provided in Quintavalle et al. (2021).
Definition 5 (Confinement, Definition 1 in Quintavalle et al. (2021)).
[…]
$f(|\sigma(e)|)\geq|e|^{\mathrm{red}}.$ (7)
In this definition, $|e|^{\text{red}}$ denotes the reduced weight of the error $e$, which represents the smallest weight of any error $e^{\prime}$ that produces the same error syndrome as $e$.
Unlike Bombín’s more restrictive formulation (Definition 16 in Bombín (2015)), the above definition permits non-linear functions $f(x)$. This broader scope proves significant, the 2D repetition code and 3D hypergraph product codes both exhibit superlinear confinement functions while supporting single-shot correction. These important cases would be excluded under Bombín’s original framework.
A code family is said to have good confinement if each code within the family satisfies $(t,f)$-confinement with the following criteria: (1) $t$ increases with the length of the code $n$, specifically $t\in\Omega\left(n^{b}\right)$ for some $b>0$, and (2) the confinement function $f(\cdot)$ increases monotonically and does not depend on $n$. Additionally, a code family is said to exhibit good X-confinement if this property applies only to Pauli Z errors Quintavalle et al. (2021).
Confinement is a weaker requirement than soundness. While good soundness requires low-weight syndromes to correspond to low-weight errors, good confinement demands that low-weight errors produce low-weight syndromes. For an LDPC code that is $(t,f)$-sound, if the maximum qubit degree is $\omega_{q}$ (the sparsity parameter), then the code has $(\frac{t}{\omega_{q}},f)$ confinement (see Lemma 2 in  Quintavalle et al. (2021)). This follows from a direct analysis: if $e$ is an error set with $|e|^{\text{red}}\leq\frac{t}{\omega_{q}}$, then its syndrome satisfies $|\sigma(e)|\leq\frac{t}{\omega_{q}}\cdot\omega_{q}=t$. By the soundness of the code, we have $f(|\sigma(e)|)\geq|e|^{\mathrm{red}}$.
Importantly, one can show that codes with good confinement possess single-shot error correction capabilities Quintavalle et al. (2021).
Lemma 6 (Theorem 1 in Quintavalle et al. (2021)).
Consider a family of $[[n,k,d]]$ quantum-LDPC codes with good confinement such that $d\geq an^{b}$ with $a,b>0$. This code family is single-shot for the adversarial noise model. If the code family only has good X-confinement then it is single-shot with respect to Pauli Z noise.
Confinement is also related to the energy barrier of a code. In Sec. III, we demonstrate formally that confinement naturally induces an energy barrier lower bound.
II.5 Expander code
An expander graph embodies a bipartite graph characterized by significant connectivity. Let $G=(V\cup C,E)$ be a bipartite graph. For a vertex subset $S\subseteq V$, a vertex $g\in C$ is a neighbor of $S$ if it is connected to at least one vertex in $S$, $N(S)$ is the set of neighbors of $S$. We say $g\in C$ is a unique neighbor of $S$ if it is only connected to a single vertex in $S$. We use $U(S)$ to denote the collection of unique neighbors of $S$. The expander graph can be defined as follows:
Definition 7 (Expander graph).
Let $G=(V\cup C,E)$ be a bipartite graph with left and right degrees bounded by $w$. Let $|V|=n$ and $|C|=m$. We say $G$ is $(\gamma_{1},\partial_{1})$-left-expanding with constants $\gamma_{1},\partial_{1}>0$, if for any subset $S\subseteq V$ with $|S|\leq\gamma_{1}n$, the neighbor $N(S)$ of $S$ in the subset $C$ satisfies
$|N(S)|\geqslant w(1-\partial_{1})|S|.$ (8)
Similarly, We say $G$ is $(\gamma_{2},\partial_{2})$-right-expanding with constants $\gamma_{2},\partial_{2}>0$, if for any subset $T\subseteq C$ with $|T|\leq\gamma_{2}n$, the neighbor $N(T)$ of $T$ in the subset $V$ satisfies $|N(T)|\geqslant w(1-\partial_{2})|T|$. We call $G$ an $\left(n,m,w,\gamma_{1},\partial_{1},\gamma_{2},\partial_{2}\right)$ expander graph if it is both left and right expanding.
Let $H_{G}$ be the parity check matrix of an expander code defined by this $\left(n,m,w,\gamma_{1},\partial_{1},\gamma_{2},\partial_{2}\right)$ expander graph (choosing $V$ and $C$ as the variable and check nodes of the tanner graph) with $\partial_{1}<1/2$. Then for any $S\subset V$ such that $|S|\leq\gamma_{1}n,\quad U(S)\geq w(1-2\partial_{1})|S|$. Furthermore, the distance $d_{C}$ of the code $\mathcal{C}(G)$ is greater than $\gamma_{1}n$ Sipser and Spielman (1996).
The above analysis directly establishes the confinement property of expander codes. For any error $e$ with reduced weight $|e|^{\mathrm{red}}<d$, we have $|\sigma(e)|\geqslant U(e)\geqslant w(1-2\partial_{1})|e|$. Noting that in classical codes, $|e|^{\mathrm{red}}=|e|$. We can then conclude that $({1}/({w(1-2\partial_{1})}))|\sigma(e)|\geqslant|e|^{\mathrm{red}}$. This relationship demonstrates that expander codes exhibit linear confinement.
Conversely, linear confinement also induces left-expansion. For $(t,f)$-confinement with $t=an$ and $f(x)=cx$, where $a$ and $c$ are constants, any error $S$ with $|S|<an$ satisfies $c|\sigma(S)|\geqslant|S|$. Thus $N(S)\geqslant|\sigma(S)|\geqslant(1/c)|S|$. With degree $w$, we have $N(S)\geqslant w(1-(1-\frac{1}{cw}))|S|$. Therefore, this code is an expander code with $(a,1-\frac{1}{cw})$-lift-expanding.
Furthermore, the energy barrier of expander codes can be characterized by the following lemma.
Lemma 8.
let $\Delta(H_{G})$ be the energy barrier of classical expander code defined by $H_{G}$ with parameters $(n,m,w,\gamma_{1},\partial_{1},\gamma_{2},\partial_{2})$ and $\partial_{1}<\frac{1}{2}$, then
$\Delta(H_{G})\geq cn,$ (9)
where $c>0$ is a constant.
Proof.
Without loss of generality, suppose that energy barrier $\Delta(H_{G})$ is equal to an energy barrier of a logical operator $L$, with path $r_{L}$. Since $|L|\geq\gamma n$, there is a point along the path (denoted as $P$) such that $\left|P\right|=\gamma n$. Because $U(P)\geq w(1-2\partial_{1})|P|$, we get $\Delta(H_{G})\geq U(P)\geq w(1-2\partial_{1})|P|=w(1-2\partial_{1})\gamma_{1}n$. Thus we have $\Delta(H_{G})\geq cn$ for $c=w(1-2\partial_{1})\gamma_{1}>0$. ∎
From Lemma 8, it follows that the classical expander code has an energy barrier scaling linearly with the number of bits. A similar conclusion follows the transpose code in which the variable and the check nodes are exchanged. Therefore, according to the results in Zhao et al. (2025), the quantum expander code defined by parity check matrix $H_{(H_{G},H_{G})}$ has $\Omega(\sqrt{n})$ energy barrier, where $n$ here is the number of qubits.
Research has established that soundness requires check redundancy in LDPC codes Ben-Sasson et al. (2010, 2003). Consequently, the expander code exhibits poor soundness due to its lack of check redundancy. However, despite this limitation, the expander code has been proven to support single-shot error correction when implemented with the small-set flip decoder Fawzi et al. (2020, 2018a). This makes confinement a more general and inclusive property related to single-shot error correction, compared to soundness.
II.6 Higher-dimensional hypergraph product
The hypergraph products provide a way to construct quantum codes with any two classical linear codes Tillich and Zemor (2014). It enables the application of extensive insights from classical coding theory to quantum coding. Those codes have good properties, such as circuit-level distance preservation Manes and Claes (2025); Tan and Stambler (2025).
Given two classical codes $C_{1}$ and $C_{2}$, with respective parity-check matrices $H_{1}$ and $H_{2}$, the hypergraph product defines a quantum code with parity-check matrices
$H_{X}$ $=\left(H_{1}\otimes I\quad I\otimes H_{2}^{T}\right),$ (10)
$H_{Z}$ $=\left(I\otimes H_{2}\quad H_{1}^{T}\otimes I\right).$
One can verify that $H_{X}H_{Z}^{T}=0$ within the field $\mathbb{F}_{2}$. Thus these two parity check matrices define a CSS code, with a quantum parity check matrix of the form
$H_{\left(H_{1},H_{2}\right)}=\left(\begin{array}[]{cc}H_{X}&0\\ 0&H_{Z}\end{array}\right).$ (11)
Hypergraph product codes can be interpreted through chain complexes. Classical codes correspond to length-2 chain complexes, where boundary maps define parity check matrices. A hypergraph product code arises from the tensor product of two such chain complexes, yielding a length-3 chain complex. This structure provides two boundary maps that serve as parity checks. The fundamental property that “a boundary has no boundary” ensures these checks commute, allowing the construction of CSS quantum codes directly from the boundary maps of the resulting length-3 chain complex.
The HHGP codes Zeng and Pryadko (2019) extends this concept by generalizing the construction to create a length-$n$ chain complex. This generalization has important connections to topological codes: $n$-dimensional toric codes on hypercubic lattices constitute an length-$n$ chain complex Bombin and Martin-Delgado (2007); Dennis et al. (2002), and can be perfectly described by higher-dimensional hypergraph products.
Following Ref. Zeng and Pryadko (2019), we give a short introduction to the chain complex interpretation of hypergraph product. A chain complex is a sequence of finite-dimensional vector spaces $\{\ldots,\mathcal{A}_{j-1},\mathcal{A}_{j},\ldots\}$ connected by boundary operators $\partial_{j}:\mathcal{A}_{j}\rightarrow\mathcal{A}_{j-1}$, satisfying $\partial_{j}\partial_{j+1}=0$ for all $j\in\mathbb{Z}$.
For qubit system, we consider vector spaces $\mathcal{A}_{j}=\mathbb{F}_{2}^{n_{j}}$, consisting of binary vectors of length $n_{j}$. $\mathcal{A}\equiv\mathcal{K}\left(A_{1},\ldots,A_{m}\right)$ can be defined as a length-$(m+1)$ chain complex with boundary map $\partial_{j}$ represented by $n_{j-1}\times n_{j}$ binary matrices $A_{j}$:
$\mathcal{A}:\{0\}\stackrel{{\scriptstyle\partial_{0}}}{{\leftarrow}}\mathcal{A}_{0}\stackrel{{\scriptstyle A_{1}}}{{\leftarrow}}\mathcal{A}_{1}\ldots\stackrel{{\scriptstyle A_{m}}}{{\leftarrow}}\mathcal{A}_{m}\stackrel{{\scriptstyle\partial_{\mathrm{m}+1}}}{{\leftarrow}}\{0\},$ (12)
where adjacent matrices satisfy the orthogonality condition $A_{j-1}A_{j}=0$ for $j\in\{1,\ldots,m\}$. The trivial operators $\partial_{0}:\{0\}\rightarrow\mathcal{A}_{0}$ and $\partial_{m+1}:\mathcal{A}_{m}\rightarrow\{0\}$ are treated as zero matrices of dimensions $0\times n_{0}$ and $n_{m}\times 0$, respectively.
The subspace $\operatorname{Im}\left(A_{j+1}\right)\subseteq\mathcal{A}_{j}$ consists of boundaries, which can be represented as linear combinations of the columns of $A_{j+1}$. This subspace defines a binary linear code with generator matrix $A_{j+1}^{T}$, denoted as $\operatorname{Im}\left(A_{j+1}\right)=\mathcal{C}_{A_{j+1}^{T}}$. For the special case where $j=m$, we have $\operatorname{Im}\left(\partial_{m+1}\right)={0}$, which forms a trivial vector space.
The subspace $\operatorname{Ker}\left(A_{j}\right)\subseteq\mathcal{A}_{j}$ consists of cycles, which are vectors $x\in\mathcal{A}_{j}$ orthogonal to the rows of $A_{j}$, satisfying $A_{j}x^{T}=0$. This subspace defines a binary linear code with parity check matrix $A_{j}$, expressed as $\operatorname{Ker}\left(A_{j}\right)=\mathcal{C}_{A_{j}}^{\perp}$. In the special case where $j=0$, we have $\operatorname{Ker}\left(\partial_{0}\right)=\mathcal{A}_{0}$, encompassing the entire vector space.
Within the binary vector space $\mathcal{A}_{j}$, we have defined two key subspaces: $\operatorname{Im}\left(A_{j+1}\right)$, consisting of boundaries, and $\operatorname{Ker}\left(A_{j}\right)$, comprising cycles. A fundamental property of chain complexes is that all boundaries are cycles, though not all cycles are boundaries. This relationship is mathematically expressed through the orthogonality condition $A_{j}A_{j+1}=0$.
Based on these subspaces, we can define the $j$-th homology group as
$H_{j}(\mathcal{A})\equiv\operatorname{Ker}\left(A_{j}\right)/\operatorname{Im}\left(A_{j+1}\right).$ (13)
This quotient space contains all elements that are cycles but not boundaries, which we refer to as homologically nontrivial cycles.
The rank of the $j$-th homology group, representing the dimension of the associated vector space, is given by
$k_{j}\equiv\operatorname{rank}H_{j}(\mathcal{A})=n_{j}-\operatorname{rank}A_{j}-\operatorname{rank}A_{j+1}.$ (14)
Additionally, the homological distance $d_{j}$ is defined as the minimum Hamming weight of a nontrivial element in $H_{j}(\mathcal{A})$, expressed as:
$d_{j}=\min_{0\neq x\in H_{j}(\mathcal{A})}\operatorname{wt}(x).$ (15)
A quantum CSS code can be derived from a length-3 chain complex $\mathcal{A}_{0}\stackrel{{\scriptstyle A_{1}}}{{\leftarrow}}\mathcal{A}_{1}\stackrel{{\scriptstyle A_{2}}}{{\leftarrow}}\mathcal{A}_{2}$ by setting $H_{Z}^{T}=A_{1}$ and $H_{X}=A_{2}$. In chain complex terminology, the code length equals $\operatorname{dim}\mathcal{A}_{1}$, while its dimension corresponds to the first homology group $\mathcal{H}_{1}=\operatorname{ker}A_{1}/\operatorname{Im}A_{2}$, or equivalently, the first cohomology group $\mathcal{H}_{1}^{*}=\operatorname{ker}A_{1}^{T}/\operatorname{Im}A_{2}^{T}$. The X and Z distances are determined by the minimum weights of non-zero vectors in $\mathcal{H}_{1}^{*}$ and $\mathcal{H}_{1}$ respectively.
Classical codes can be described by length-2 chain complexes. To construct quantum codes from classical codes, It just the question how to build a length-3 chain complex from two length-2 chain complexes. For this purpose, we employ the tensor product of chain complexes.
The tensor product $\mathcal{A}\times\mathcal{B}$ of two chain complexes $\mathcal{A}$ and $\mathcal{B}$ is defined as a new chain complex in the following way
$(\mathcal{A}\times\mathcal{B})_{l}=\bigoplus_{i+j=l}\mathcal{A}_{i}\otimes\mathcal{B}_{j}.$ (16)
The vector spaces in the new chain complex are direct sums of Kronecker products of vector spaces from the original chain complexes. Moreover, the boundary operators in this tensor product construction act as
$\partial_{i+j}(a\otimes b)=\partial_{i}^{\prime}a\otimes b+(-1)^{i}a\otimes\partial_{j}^{\prime\prime}b,$ (17)
where $a\in\mathcal{A}_{i},b\in\mathcal{B}_{j}$, and $\partial_{i}^{\prime}$ and $\partial_{j}^{\prime\prime}$ are the boundary operators in the corresponding space $\mathcal{A}$ and $\mathcal{B}$. If both $\mathcal{A}$ and $\mathcal{B}$ have finite dimension, then the dimension of $\mathcal{C}=\mathcal{A}\times\mathcal{B}$ is
$n_{j}(\mathcal{C})=\sum_{i}n_{i}(\mathcal{A})n_{j-i}(\mathcal{B}).$ (18)
By the Künneth theorem, the homology groups of the product complex $\mathcal{C}=\mathcal{A}\times\mathcal{B}$ are given by
$H_{j}(\mathcal{C})\cong\bigoplus_{i}H_{i}(\mathcal{A})\otimes H_{j-i}(\mathcal{B}).$ (19)
Consequently, the rank $k_{j}(\mathcal{C})$ of the $j$-th homology group $H_{j}(\mathcal{C})$ can be expressed as
$k_{j}(\mathcal{C})=\sum_{i}k_{i}(\mathcal{A})k_{j-i}(\mathcal{B}).$ (20)
The hypergraph product code is defined as the tensor product of two length-2 chain complexes. The higher-dimensional hypergraph product extends this concept to the tensor product of two chain complexes of arbitrary lengths. As demonstrated in Ref. Zeng and Pryadko (2019), for HHGP codes formed from a length-m chain complex $\mathcal{A}$ and a length-l chain complex $\mathcal{B}$ with $l=2$, there exists a lower bound on the distance of logical operators in the resulting quantum codes:
$d_{j}(\mathcal{A}\times\mathcal{B})=\min\left[d_{j-1}(\mathcal{A})d_{1}(\mathcal{B}),d_{j}(\mathcal{A})d_{0}(\mathcal{B})\right].$ (21)
In this work, we focus specifically on those 3D and 4D hypergraph product codes that can be regarded as tensor products of three or four classical codes, respectively. This allows us to apply the distance bounds given in by above equation.
III Confinement and energy barrier
The relation between confinement and energy barrier is well-know Bombín (2015); Hong et al. (2025). In this section, we formalize this connection through the following lemma.
Lemma 9.
For a $[[n,k,d]]$ quantum code $\mathcal{C}$ with $(t,f)$ confinement where $t\leqslant d$ and $f(\cdot)$ is increasing monotonically, the energy barrier of the code is at least $f^{-1}(w)$, where $w={(d-1)}/{2}$.
Proof.
Without loss of generality, suppose that the code energy barrier $\Delta(\mathcal{C})$ is given by logical operator $L$, with a path $r_{L}$. Since $|L|\geqslant d\geqslant t$, there always exist a step along the path $r_{L}$, denoted as $P_{c}$, such that $\left|P_{c}\right|^{\mathrm{red}}=(d-1)/2$. This can be argued as following Campbell (2019):
Suppose $r_{L}=\{P_{0},P_{1},\cdots,P_{L}\}$, the relation $P_{j}P_{j-1}=E_{j}$ ensures that each $E_{j}$ is a single qubit Pauli error. For each $P_{j}$, we define the reduced weight as:
$\mathrm{wt}_{\mathrm{red}}(P_{j}):=\min_{V}\{\mathrm{wt}(P_{j}V):V\in\mathcal{P},\sigma(V)=0\},$ (22)
where the minimization is over Pauli operators $V$ with trivial syndrome. Let $V_{j}$ denote the Pauli operators achieving this minimum. Since $\sigma\left(V_{j}\right)=0$, each $V_{j}$ is either a stabilizer or a nontrivial logical operator.
Trivially, we find $V_{0}=I$ and $V_{L}=P_{L}$, indicating that the sequence starts with a stabilizer and ends with a logical operator. Therefore, there exists an index $j^{\star}$ such that $V_{j^{*}}$ is a stabilizer, while $V_{j^{*}+1}$ is a nontrivial logical operator. Then $V_{j^{*}}V_{j^{*}+1}$ is also a nontrivial logical operator, leading to:
$d\leq\mathrm{wt}\left(V_{j^{*}}V_{j^{*}+1}\right).$ (23)
Notice that
$\operatorname{wt}\left(V_{j^{*}}V_{j^{*}+1}\right)$ $=$ $\operatorname{wt}\left(V_{j^{*}}V_{j^{*}+1}P_{j^{*}}P_{j^{*}}^{\dagger}P_{j^{*}+1}P_{j^{*}+1}^{\dagger}\right)$
$=$ $\operatorname{wt}\left(V_{j^{*}}P_{j^{*}}V_{j^{*}+1}P_{j^{*}+1}P_{j^{*}}^{\dagger}P_{j^{*}+1}^{\dagger}\right).$
Apply the triangle inequality twice gives:
$\mathrm{wt}\left(V_{j^{*}}V_{j^{*}+1}\right)\leq\mathrm{wt}_{\mathrm{red}}\left(P_{j^{*}}\right)+\mathrm{wt}_{\mathrm{red}}\left(P_{j^{*}+1}\right)+1.$ (25)
Thus, combining this result, we obtain:
$d\leq 2\max\left[\mathrm{wt}_{\mathrm{red}}\left(P_{j^{*}}\right),\mathrm{wt}_{\mathrm{red}}\left(P_{j^{*}+1}\right)\right]+1,$ (26)
which implies
$\frac{d-1}{2}\leq\max\left[\mathrm{wt}_{\mathrm{red}}\left(P_{j^{*}}\right),\mathrm{wt}_{\mathrm{red}}\left(P_{j^{*}+1}\right)\right]$ (27)
The sequence of reduced weights $\left\{\mathrm{wt}_{\mathrm{red}}\left(P_{0}\right),\mathrm{wt}_{\text{red}}\left(P_{1}\right),\cdots,\mathrm{wt}_{\text{red }}\left(P_{L}\right)\right\}$ starts and ends at zero. It must reach a value of $\left(d-1\right)/2$ or higher. Moreover, the local error condition ensures that $\left|\mathrm{wt}_{\mathrm{red}}\left(P_{j+1}\right)-\mathrm{wt}_{\text{red }}\left(P_{j}\right)\right|$ is either 0 or 1, then the sequence must include every integer from 0 to $\left(d-1\right)/2$.
Recall $(t,f)$ confinement means that for all errors $e$ with $|e|^{\text{red}}\leq t$, $f(|\sigma(e)|)\geq|e|^{\mathrm{red}}$. Then for $P_{c}$, set $w=(d-1)/2$, we have $f(|\sigma(P_{c})|)\geqslant|P_{c}|^{\mathrm{red}}\geqslant w$. Given $f(\cdot)$ is monotonically increasing, $|\sigma(P_{c})|\geqslant f^{-1}(w)$. This implies that the energy barrier of code $\mathcal{C}$ is at least $f^{-1}(w)$. ∎
Quintavalle et al. proved that all 3D hypergraph product codes exhibit $(t,f)$ X-confinement (according to their convention) where $t$ equals the minimum distance of classical codes and $f(x)=x^{3}/4$ (see Lemma 10 of Ref Quintavalle et al. (2021)). When the 3D hypergraph product code is constructed using three identical classical codes defined by the same parity check matrix $\delta$ with distance $d$, their result implies a lower bound for the energy barrier of Z-type logical operators as $O\left(d^{\frac{1}{3}}\right)$. In section V, we will demonstrate that the optimal lower bound for this energy barrier actually scales as $O(d)$ in this scenario.
The converse statement of Lemma 9—that any quantum code with a macroscopic energy barrier necessarily exhibits confinement—does not generally hold. This can be demonstrated with a straightforward counterexample.
Consider a quantum LDPC code $\mathcal{C}$ constructed as a composite system comprising two regions: Region A, based on a 2D quantum repetition code $[[n,1,n]]$ with energy barrier $O(\sqrt{n})$, and Region B, based on a 1D quantum repetition code $[[n,1,n]]$ with constant energy barrier $O(1)$, see Fig. 1. The overall code is a $[[2n,1,2n]]$ quantum repetition code with energy barrier $O(\sqrt{n})$ determined by Region A. However, for an error $e$ contained entirely within Region B, we have $|e|^{\mathrm{red}}=t$ that grows up to $(d-1)/2$ while maintaining $|\sigma(e)|\leq c$, where $c$ is constant. Consequently, the overall code fails to satisfy the confinement criterion.
Region ARegion B
Figure 1: A composite system consisting of a 2D quantum repetition code (Region A) and a 1D quantum repetition code arranged in a snake pattern (Region B). The overall code exhibits an energy barrier determined by the 2D quantum repetition code, which scales as the linear dimension of Region A. However, confinement properties are influenced by the 1D quantum repetition component, resulting in the absence of confinement for the composite system.
IV Tensor product code and its energy barrier
In this section, we analyze the energy barrier of tensor product codes — a specific classical code structure that builds new classical codes from two existing ones. We will demonstrate that the energy barrier of the resulting code is related to the distance of the underlying classical codes.
We need this result for our subsequent analysis about HHGP codes. Specifically, we will show that the logical operators of HHGP codes contain this tensor product structure. This connection allow us to establish a lower bound for the energy barrier of HHGP codes.
IV.1 Code construction
Given two classical codes with parity check matrix $\delta_{a}$ and $\delta_{b}$, the tensor product code $\delta_{c}$ is a classical code defined to have the parity check matrix
$\delta_{c}=\left(\begin{array}[]{c}\delta_{a}\otimes I_{n_{b}}\\ I_{n_{a}}\otimes\delta_{b}\end{array}\right).$ (30)
The tensor product code structure can be also understood from the perspective of code space. Let $\mathcal{C}_{a}$ with parameters $\left[n_{a},k_{a},d_{a}\right]$ and $\mathcal{C}_{b}=$ with parameters $\left[n_{b},k_{b},d_{b}\right]$ be two classical codes. The tensor product of these two codes, denoted $\mathcal{C}_{a}\otimes\mathcal{C}_{b}$, encodes messages as follows: First, represent a message of $\mathcal{C}_{a}\otimes\mathcal{C}_{b}$ as a $k_{a}\times k_{b}$ matrix $M$. Then, encode each row of $M$ using $\mathcal{C}_{a}$, resulting in an intermediate $k_{b}\times n_{a}$ matrix. Finally, encode each column of this intermediate matrix using $\mathcal{C}_{b}$, producing an $n_{b}\times n_{a}$ matrix representing the codeword of the tensor product code. Furthermore, encoding the columns with $\mathcal{C}_{b}$ first, followed by the rows with $\mathcal{C}_{a}$, yields the same result.
Let $L_{a}$ be the logical operator of $\delta_{a}$, and $L_{b}$ be the logical operator of $\delta_{b}$, one can verify that $L_{c}=L_{a}\otimes L_{b}$ is the logical operator of $\delta_{c}$ because
$\delta_{c}L_{c}$ $=$ $\left(\begin{array}[]{c}\delta_{a}\otimes I_{n_{b}}\\ I_{n_{a}}\otimes\delta_{b}\end{array}\right)(L_{a}\otimes L_{b})$ (33)
$=$ $\left(\begin{array}[]{c}(\delta_{a}\otimes I_{n_{b}})(L_{a}\otimes L_{b})\\ (I_{n_{a}}\otimes\delta_{b})(L_{a}\otimes L_{b})\end{array}\right)$ (36)
$=$ $0$ (37)
Here we used the relation $\delta_{a}L_{a}=\delta_{b}L_{b}=0$.
Given $k_{a}$ distinct logical operators $L_{a}$ and $k_{b}$ distinct logical operators $L_{b}$, one can construct $k_{a}k_{b}$ distinct logical operators $L_{c}$ through this tensor product construction.
The distance of the tensor product code is at least $d_{a}d_{b}$. A simple argument demonstrates this: consider two different message matrices of code $\mathcal{C}_{c}$, denoted as $M_{a}$ and $M_{b}$. After the first encoding step, we obtain intermediate matrices $N_{a}$ and $N_{b}$, and the second encoding step produces the final codewords $C_{a}$ and $C_{b}$. If $M_{a}$ and $M_{b}$ differ in the $i$-th row, then $N_{a}$ and $N_{b}$ differ in at least $d_{a}$ positions in that row, corresponding to $d_{a}$ columns. Let the indices of these differing columns be $\{j_{1},\ldots,j_{d_{a}}\}$. During the second encoding step, each of these columns generates differences in at least $d_{b}$ positions. Consequently, $C_{a}$ and $C_{b}$ differ in at least $d_{a}d_{b}$ positions.
In summary, the tensor product code $\delta_{c}$ has parameters $[n_{a}n_{b},k_{a}k_{b},d_{a}d_{b}]$. This construction bears strong similarity to the concatenated code, which shares the same code parameters $[n_{a}n_{b},k_{a}k_{b},d_{a}d_{b}]$, but employs a different parity check matrix $\delta_{c}=\delta_{a}\otimes\delta_{b}$.
IV.2 Energy barrier of tensor product code
Let $L_{a}$ and $L_{b}$ be the logical operators of $\delta_{a}$ and $\delta_{b}$. Our goal is to determine the energy barrier of the operator $L_{c}=L_{a}\otimes L_{b}$, which serves as a logical operator for $\delta_{c}$.
To compute the energy barrier of $L_{c}$, we consider a path $r=\{P_{0},P_{1},\cdots,P_{F}\}$ with $P_{0}=I$ and $P_{F}=L_{c}$. Then
$\Delta(L_{c})=\min_{r\in w(0,L_{c})}\epsilon_{\max}(r).$ (38)
Note that $r$ is a path that walks on the bit of code $\delta_{c}$. Given a $P_{\ell}\in r$, the energy (the number of violated checks) of $P_{\ell}$ is
$\epsilon(P_{\ell})$ $=$ $\text{wt}(\delta_{c}P_{\ell})$ (41)
$=$ $\text{wt}\left(\left(\begin{array}[]{c}\delta_{a}\otimes I_{n_{b}}\\ I_{n_{a}}\otimes\delta_{b}\end{array}\right)P_{\ell}\right)$
$=$ $\text{wt}((\delta_{a}\otimes I_{n_{b}})P_{\ell})+\text{wt}((I_{n_{a}}\otimes\delta_{b})P_{\ell}).$ (42)
Because $\text{wt}(\cdot)\geqslant 0$, then we have
$\epsilon(P_{\ell})\geqslant\text{wt}((\delta_{a}\otimes I_{n_{b}})P_{\ell}),$ (43)
or
$\epsilon(P_{\ell})\geqslant\text{wt}((I_{n_{a}}\otimes\delta_{b})P_{\ell}).$ (44)
Because the path $r$ contains paths for $L_{a}$ and $L_{b}$, the above inequalities indicate that the energy barrier of $L_{c}$ is lower bounded by the energy barrier of $L_{a}$ or $L_{b}$. Furthermore, the energy barrier of code $\delta_{c}$ is then lower bounded by energy barrier of codes $\delta_{a}$ and $\delta_{b}$. We conclude this as the following Lemma
Lemma 10.
Given two classical codes $\delta_{a}$ and $\delta_{b}$ with parameters $[n_{a},k_{a},d_{a},E_{a}]$ and $[n_{b},k_{b},d_{b},E_{b}]$, where $E_{a}$ and $E_{b}$ are energy barriers of codes. Then the energy barrier of the tensor product code $\delta_{c}$ is lower bounded by
$\Delta(\delta_{c})\geqslant\min(E_{a},E_{b}).$ (45)
However, this may not be the tightest bound, particularly when codes $\delta_{a}$ and $\delta_{b}$ have constant energy barriers. In the following, we establish a relationship between the energy barrier of $L_{c}$ and the distance of $L_{a}$ and $L_{b}$.
IV.2.1 Lower bound
Given that the energy barriers of $\delta_{a}$ and $\delta_{b}$ are $\Delta(\delta_{a})=E_{a}$ and $\Delta(\delta_{b})=E_{b}$. Here, we show that the energy barrier of the tensor product code $\delta_{c}$ can be lower bounded by the minimum of the distances of the two constituent codes, as formally described in the following lemma.
Lemma 11.
Given two classical codes $\delta_{a}$ and $\delta_{b}$ with parameters $[n_{a},k_{a},d_{a},E_{a}]$ and $[n_{b},k_{b},d_{b},E_{b}]$, where $E_{a}$ and $E_{b}$ are energy barriers of codes. Then the energy barrier of the tensor product code $\delta_{c}$ is lower bounded by
$\Delta(\delta_{c})\geqslant\min(d_{a},d_{b}).$ (46)
Proof.
Let us first establish a clear representation of the tensor product code structure. Given the parity check matrices $\delta_{a}\in\{0,1\}^{m_{a}\times n_{a}}$ and $\delta_{b}\in\{0,1\}^{m_{b}\times n_{b}}$, we can view the tensor product code $\delta_{c}$ as operating on a two-dimensional grid of size $n_{a}\times n_{b}$. In this representation, each row corresponds to a copy of code $\delta_{b}$, each column corresponds to a copy of code $\delta_{a}$.
The logical operator $L_{c}=L_{a}\otimes L_{b}$ can be visualized as a pattern on this grid, where $L_{a}$ specifies which columns are involved and $L_{b}$ specifies which rows are involved. Specifically, the support of $L_{c}$ consists of positions $(i,j)$ where $i\in\text{supp}(L_{a})$ and $j\in\text{supp}(L_{b})$, forming a subgrid of size $|L_{a}|\times|L_{b}|=d_{a}\times d_{b}$.
Now, consider any valid path $r=\{P_{r}^{0},P_{r}^{1},P_{r}^{2},\ldots,P_{r}^{L}\}$ from the identity operator $I$ to the logical operator $L_{c}$. By definition, this path must flip all the bits at the positions in the support of $L_{c}$.
For any intermediate state $P_{r}^{i}$ along this path, the syndrome weight $\text{wt}(\delta_{c}P_{r}^{i})$ measures the energy cost at that step. We will show that for any valid path, there must exist some intermediate state with syndrome weight at least $\min(d_{a},d_{b})$.
Consider the following key observation: To implement the logical operator $L_{c}$, the path must flip bits in at least one complete row or one complete column of the support of $L_{c}$. This is because:
  1. 1.
    If no complete row contains a logical operator of $\delta_{a}$, then the final state cannot be $L_{c}=L_{a}\otimes L_{b}$
  2. 2.
    Similarly, if no complete column contains a logical operator of $\delta_{b}$, then the final state cannot be $L_{c}$
Therefore, at some point in the path, there must exist an intermediate state $P_{r}^{i^{*}}$ where either:
  1. 1.
    Exactly one row $j\in\text{supp}(L_{b})$ contains a non-trivial logical operator $L^{\prime}_{a}$ of code $\delta_{a}$ (not necessarily to be $L_{a}$), or
  2. 2.
    Exactly one column $i\in\text{supp}(L_{a})$ contains a non-trivial logical operator $L^{\prime}_{b}$ of code $\delta_{b}$ (not necessarily to be $L_{b}$)
In the first case, consider the columns that contain the logical operator $L^{\prime}_{a}$. By definition, each of these columns contains at least one bit flip but does not form a complete logical operator of $\delta_{b}$. Therefore, each such column contributes at least 1 to the syndrome weight. Since any non-trivial logical operator of code $\delta_{a}$ has weight at least $d_{a}$, there are at least $d_{a}$ such columns. Thus, the total syndrome weight contribution is at least $d_{a}\times 1=d_{a}$.
In the second case, by same reasoning, the row containing the logical operator $L^{\prime}_{b}$ has at least $d_{b}$ bit flips, each contributing at least 1 to the syndrome weight, for a total contribution of at least $d_{b}$.
Therefore, for any valid path $r$, there exists some intermediate state with syndrome weight at least $\min(d_{a},d_{b})$. Since the energy barrier $\Delta(L_{c})$ is defined as the minimum over all paths of the maximum syndrome weight along the path, we have:
$\Delta(\delta_{c})\geqslant\min(d_{a},d_{b}).$ (47)
The intuition behind Lemma 11 can be illustrated through the 2D repetition code, which can be interpreted as the tensor product of two 1D repetition code. The 1D repetition code has parameters $[L,1,L,O(1)]$, where the energy barrier scales as $O(1)$. In contrast, the 2D repetition code exhibits parameters $\left[L^{2},1,L^{2},O(L)\right]$, with an energy barrier scaling as $O(L)$. These observations align with the lemma’s predictions, where $L$ denotes the linear dimension of both systems.
The energy barrier of the 2D repetition code can be rigorously analyzed through its tensor product structure. Consider a 2D repetition code on an $L\times L$ square lattice, where the logical operator corresponds to flipping all bits on the lattice. For any valid path implementing this logical operator, there necessarily exists an intermediate configuration where precisely a single row or a single column of bits has been flipped. At this configuration, the number of violated checks (i.e., the energy cost) is at least $L$, corresponding to the boundary of the flipped region. Consequently, the energy barrier of the 2D repetition code is lower-bounded by $L$, demonstrating the linear scaling with system size predicted by the lemma.
IV.2.2 Upper bound - “strip” argument
To establish an upper bound, we consider specific paths from the identity to a logical operator. We analyze two distinct implementation strategies: row-wise and column-wise implementation of logical operators. Same to the “strip” argument in Ref. Rakovszky and Khemani (2024).
We still view the tensor product code $\delta_{c}$ as operating on a 2D grid of size $n_{a}\times n_{b}$. In this representation, each row corresponds to a copy of code $\delta_{b}$, each column corresponds to a copy of code $\delta_{a}$.
For the row-wise implementation, let $r_{a}=\{P_{a}^{0},P_{a}^{1},\ldots,P_{a}^{F}\}$ represent an optimal path for implementing logical operator $L_{a}$, where $P_{a}^{0}=I$ and $P_{a}^{F}=L_{a}$. By definition, the maximum syndrome weight along the path $r_{a}$ is $E_{a}$. Because $P_{i}P_{i+1}=E_{i}$ is a single Pauli for all $i$, we denote $q_{a}=\{j_{1},j_{2},\ldots,j_{L_{a}}\}$ as the position of single Pauli in each step of $r_{a}$. These positions correspond to rows on the 2D grid of the tensor product code.
We consider a path for implementing $L_{c}$ by applying $L_{b}$ sequentially according to the positions in $q_{a}$. Specifically, Our path $r_{c}$ for implementing $L_{c}$ begins with $P_{c}^{0}=I$, for each position $j_{k}\in q_{a}$, we implement the logical operator $L_{b}$ on the corresponding row. After finishing all rows in $q_{a}$, we have $P_{c}^{F}=L_{c}$.
Now consider the energy barrier given by this specific $r_{c}$. Given any $P_{c}^{i}\in r_{c}$, the energy of $P_{c}^{i}$ is
$\epsilon(P_{c}^{i})=\text{wt}((\delta_{a}\otimes I_{n_{b}})P_{c}^{i})+\text{wt}((I_{n_{a}}\otimes\delta_{b})P_{c}^{i}).$ (48)
The maximum syndrome weight along this path is less than $d_{b}E_{a}+E_{b}$. This follows because for the part $\text{wt}((\delta_{a}\otimes I_{n_{b}})P_{c}^{i})$, each column contributes a syndrome weight of at most $E_{a}$, and there are $d_{b}$ columns where $L_{b}$ has support, which gives $d_{b}E_{a}$. While for the part $\text{wt}((I_{n_{a}}\otimes\delta_{b})P_{c}^{i})$, the maximum syndrome weight is $E_{b}$ because we implement $L_{b}$ sequentially.
Analogously, for column-wise implementation, let $r_{b}=\{P_{b}^{0},P_{b}^{1},\ldots,P_{b}^{F}\}$ represent an optimal path for implementing $L_{b}$, with maximum syndrome weight $E_{b}$. One can construct another path for $L_{c}$ by implementing $L_{a}$ on columns according to the position provided by path $r_{b}$. The maximum syndrome weight along this path is less than $d_{a}E_{b}+E_{a}$.
By selecting the implementation strategy with the lower maximum syndrome weight, we establish the upper bound:
$\Delta(L_{c})\leqslant\min(d_{b}E_{a}+E_{b},d_{a}E_{b}+E_{a}).$ (49)
In addition, we assert that (although we cannot provide a formal proof) there is no path with an energy barrier lower than $\min(d_{b}E_{a},d_{a}E_{b})$, which we formally state in the following conjecture.
Conjecture 12.
Given two classical codes $\delta_{a}$ and $\delta_{b}$ with parameters $[n_{a},k_{a},d_{a},E_{a}]$ and $[n_{b},k_{b},d_{b},E_{b}]$, the energy barrier of the tensor product code $\delta_{c}$ is
$\Delta(\delta_{c})\geqslant\min(d_{a}E_{b},E_{a}d_{b}).$ (50)
We conjecture this result and anticipate that an elegant, rigorous proof will be found in the future (See more discussion in the Appendix). Meanwhile, the lower bound provided in Lemma 11 is optimal with respect to the order of the distance, although the value $E_{a}$ and $E_{b}$ may yield a non-negligible difference in the absolute value of the energy barrier. Thus this lower bound is tight if underlying codes exhibit system size-dependent distances but constant energy barriers.
V Energy barrier of 3D hypergraph product codes
With all the necessary preparations are in place, we now turn our attention to the energy barrier of HHGP codes. In this section, we consider the simplest instance, 3D hypergraph product codes that constructed from three classical codes.
We begin by reviewing the structure of 3D hypergraph product codes and analyzing their logical operators, defining a complete set as elementary canonical logical operators. We first address Z logical operators. Using the strip argument from Ref. Rakovszky and Khemani (2024), we establish an upper bound for the energy barrier. We then derive a lower bound by examining the structure of those elementary canonical Z logical operators. Based on energy barrier of tensor product codes (Lemma 11), we demonstrate that this lower bound relates to the distances of the underlying classical codes. Finally, we discuss the energy barrier of the logical X operators.
V.1 Code construction of 3D hypergraph product codes
Consider three classical codes $\delta_{a}$, $\delta_{b}$, and $\delta_{c}$ with respective parameters $[n_{\ell},k_{\ell},d_{\ell}]$, where $\ell\in\{a,b,c\}$, and the parameters of their transposes $\delta_{\ell}^{T}$ are denoted by $[n_{\ell}^{T},k_{\ell}^{T},d_{\ell}^{T}]$. For convenience, we denote $r_{\ell}\equiv n_{\ell}^{T}$. Applying the 3D hypergraph product to these classical codes generates a length-4 chain complex Quintavalle et al. (2021)
$C_{0}\xrightarrow{\partial_{0}}C_{1}\xrightarrow{\partial_{1}}C_{2}\xrightarrow{\partial_{2}}C_{3},$ (51)
where $\partial_{0},\partial_{1}$ and $\partial_{2}$ are
$\partial_{0}=\left(\begin{array}[]{l}\delta_{a}\otimes I_{n_{b}}\otimes I_{n_{c}}\\ I_{n_{a}}\otimes\delta_{b}\otimes I_{n_{c}}\\ I_{n_{a}}\otimes I_{n_{b}}\otimes\delta_{c}\end{array}\right),$ (52)
$\partial_{1}=\left(\begin{array}[]{ccc}I_{r_{a}}\otimes\delta_{b}\otimes I_{n_{c}}&\delta_{a}\otimes I_{r_{b}}\otimes I_{n_{c}}&0_{r_{a}r_{b}n_{c}\times n_{a}n_{b}r_{c}}\\ I_{r_{a}}\otimes I_{n_{b}}\otimes\delta_{c}&0_{r_{a}n_{b}r_{c}\times n_{a}r_{b}n_{c}}&\delta_{a}\otimes I_{n_{b}}\otimes I_{r_{c}}\\ 0_{n_{a}r_{b}r_{c}\times r_{a}n_{b}n_{c}}&I_{n_{a}}\otimes I_{r_{b}}\otimes\delta_{c}&I_{n_{a}}\otimes\delta_{b}\otimes I_{r_{c}}\end{array}\right),$ (53)
$\partial_{2}=\left(I_{r_{a}}\otimes I_{r_{b}}\otimes\delta_{c}\quad I_{r_{a}}\otimes\delta_{b}\otimes I_{r_{c}}\quad\delta_{a}\otimes I_{r_{b}}\otimes I_{r_{c}}\right).$ (54)
Here $I_{n}$ denotes an identity matrix of size $n\times n$.
The relation $\partial_{1}\partial_{0}=0$ allows us to construct a quantum CSS code $\mathcal{C}\left(\delta_{a},\delta_{b},\delta_{c}\right)$ with parity check matrix $H_{Z}=\partial_{0}^{T}$ and $H_{X}=\partial_{1}$. The commutation condition follows as $H_{X}H_{Z}^{T}=\partial_{1}\partial_{0}=0$.
Additionally, since $\partial_{2}\partial_{1}=0$, the matrix $M=\partial_{2}$ serves as the meta-check for $H_{X}$. This means any valid X-syndrome must satisfy the constraints imposed by $M$.
The parameters of the code $\mathcal{C}\left(\delta_{a},\delta_{b},\delta_{c}\right)$ can be determined from the classical code $\delta_{a}$, $\delta_{b}$, and $\delta_{c}$. As demonstrated in Ref. Zeng and Pryadko (2019), the resulting quantum code $\mathcal{C}\left(\delta_{a},\delta_{b},\delta_{c}\right)$ is an $\left[\left[n,k,d_{x},d_{z}\right]\right]$ code with
[…]
V.2 Logical operators of 3D hypergraph product codes
The logical operators of stabilizer codes commute with all elements in the stabilizer group while remaining distinct from it. We first examine the Z-logical operators, which are operators in $\ker(H_{X})\backslash\text{im}(H_{Z}^{T})$.
Denote the binary form of logical operators of classical codes $\delta_{a}$, $\delta_{b}$, and $\delta_{c}$ as $L_{A}=\{L_{A}^{1},L_{A}^{2},\ldots,L_{A}^{k_{a}}\}$, $L_{B}=\{L_{B}^{1},L_{B}^{2},\ldots,L_{B}^{k_{b}}\}$, and $L_{C}=\{L_{C}^{1},L_{C}^{2},\ldots,L_{C}^{k_{c}}\}$, respectively. We first show the following three sets of binary vectors gives a complete set of Z logical operators.
$L_{Z}^{1}=\sum_{i,j,k}\alpha_{ijk}\left(\begin{array}[]{c}v_{i}\otimes L_{B}^{j}\otimes L_{C}^{k}\\ 0_{n_{a}r_{b}n_{c}}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right),$ (56)
$L_{Z}^{2}=\sum_{i,j,k}\beta_{ijk}\left(\begin{array}[]{c}0_{r_{a}n_{b}n_{c}}\\ L_{A}^{i}\otimes w_{j}\otimes L_{C}^{k}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right),$
$L_{Z}^{3}=\sum_{i,j,k}\gamma_{ijk}\left(\begin{array}[]{c}0_{r_{a}n_{b}n_{c}}\\ 0_{n_{a}r_{b}n_{c}}\\ L_{A}^{i}\otimes L_{B}^{j}\otimes u_{k}\end{array}\right),$
where $v_{i},w_{j}$, and $u_{k}$ are unit vectors with length $r_{a},r_{b}$, and $r_{c}$, respectively. $0_{n}$ is a zero vector with length $n$, and $\alpha_{ijk},\beta_{ijk},\gamma_{ijk}\in\{0,1\}$. Given that $\delta_{a}L_{A}^{i}=\delta_{b}L_{B}^{j}=\delta_{c}L_{C}^{k}=0$, one can confirm that $H_{X}L_{Z}^{l}=0$ for $l\in\{1,2,3\}$.
For $L_{Z}^{1}$, the condition $H_{X}L_{Z}^{1}=0$ ensures that any X-type stabilizer from $H_{X}$ commutes with $L_{Z}^{1}$. Similarly, any Z-type stabilizer from $H_{Z}$ naturally commutes with $L_{Z}^{1}$ as they both consist solely of Pauli Z operators. To prove that $L_{Z}^{1}$ are Z logical operators, we must prove that $L_{Z}^{1}$ are not in the image of $H_{Z}^{T}$. To proceed with this analysis, we first denote:
$L_{Z}^{1,i,j,k}=\left(\begin{array}[]{c}v_{i}\otimes L_{B}^{j}\otimes L_{C}^{k}\\ 0_{n_{a}r_{b}n_{c}}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right).$ (60)
For $a_{i}\in\{0,1\}$, suppose that the linear combination of these operators
$\sum_{i}a_{i}L_{Z}^{1,i,j,k}=\left(\begin{array}[]{c}(\sum_{i}a_{i}v_{i})\otimes L_{B}^{j}\otimes L_{C}^{k}\\ 0_{n_{a}r_{b}n_{c}}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right)$ (64)
is in the image of matrix $H_{Z}^{T}$. Then, there exists a vector $x$ such that
$\sum_{i}a_{i}L_{Z}^{1,i,j,k}$ $=$ $\left(\begin{array}[]{c}(\sum_{i}a_{i}v_{i})\otimes L_{B}^{j}\otimes L_{C}^{k}\\ 0_{n_{a}r_{b}n_{c}}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right)$
$=$ $H_{Z}^{T}(x\otimes L_{B}^{j}\otimes L_{C}^{k})$
$=$ $\left(\begin{array}[]{c}(\delta_{a}\otimes I_{n_{b}}\otimes I_{n_{c}})(x\otimes L_{B}^{j}\otimes L_{C}^{k})\\ (I_{n_{a}}\otimes\delta_{b}\otimes I_{n_{c}})(x\otimes L_{B}^{j}\otimes L_{C}^{k})\\ (I_{n_{a}}\otimes I_{n_{b}}\otimes\delta_{c})(x\otimes L_{B}^{j}\otimes L_{C}^{k})\end{array}\right)$ (72)
$=$ $\left(\begin{array}[]{c}\delta_{a}x\otimes L_{B}^{j}\otimes L_{C}^{k}\\ 0_{n_{a}r_{b}n_{c}}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right).$ (76)
Recall the number of encoded bits for $\delta_{a}^{T}$ is $k_{a}^{T}$, then $\dim(\text{im}(\delta_{a}))=n_{a}-k_{a}^{T}$, it is feasible to select a minimum of $k_{a}^{T}$ unit vectors $v_{i}$, ensuring that any combination of these $v_{i}$ remains external to the image of $\delta_{a}$. Specifically, for these $k_{a}^{T}$ unit vectors $v_{i}$, there does not exist any vector $x$ such that
$\sum_{i=1}^{k_{a}}a_{i}v_{i}=\delta_{a}x.$ (77)
For fixed $L_{B}^{j}$ and $L_{C}^{k}$, these $k_{a}^{T}$ unit vectors generate $k_{a}^{T}$ operators $L_{Z}^{1,i,j,k}$ that belong to $\ker(H_{X})\backslash\text{im}(H_{Z}^{T})$. Since $\delta_{b}$ has $k_{b}$ logical bits and $\delta_{c}$ has $k_{c}$ logical bits, there are $k_{b}$ distinct $L_{B}$ and $k_{c}$ distinct $L_{C}$ operators. This yields $k_{a}^{T}k_{b}k_{c}$ unique instances of $L_{Z}^{1,i,j,k}$, which function as logical Z operators of the quantum code.
Applying the same logic to operators $L_{Z}^{2}$ and $L_{Z}^{3}$ defined in Eq. (LABEL:3dzlogicaloperators), we can demonstrate that there exist $k_{a}k_{b}^{T}k_{c}$ distinct $L_{Z}^{2,i,j,k}$ and $k_{a}k_{b}k_{c}^{T}$ distinct $L_{Z}^{3,i,j,k}$ that serve as logical Z operators of the quantum code.
This completely characterizes all $k_{a}^{T}k_{b}k_{c}+k_{a}k_{b}^{T}k_{c}+k_{a}k_{b}k_{c}^{T}$ logical Z operators of the 3D hypergraph product codes. We refer to these as elementary canonical logical Z operators.
We now examine the logical X operators, which are in the set $\ker(H_{Z})\backslash\text{im}(H_{X}^{T})$. Consider the following three types of Pauli X operators,
$L_{X}^{1}=\sum_{i,j,k}\alpha_{ijk}^{\prime}\left(\begin{array}[]{c}L_{A}^{i^{\prime}}\otimes w_{j}\otimes u_{k}\\ 0_{n_{a}r_{b}n_{c}}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right),$ (81)
$L_{X}^{2}=\sum_{i,j,k}\beta_{ijk}^{\prime}\left(\begin{array}[]{c}0_{r_{a}n_{b}n_{c}}\\ v_{i}\otimes L_{B}^{j^{\prime}}\otimes u_{k}\\ 0_{n_{a}n_{b}n_{c}}\end{array}\right),$ (85)
$L_{X}^{3}=\sum_{i,j,k}\gamma_{ijk}^{\prime}\left(\begin{array}[]{c}0_{r_{a}n_{b}n_{c}}\\ 0_{n_{a}r_{b}n_{c}}\\ v_{i}\otimes w_{j}\otimes L_{C}^{k^{\prime}}\end{array}\right).$ (89)
Here $v_{i},w_{j}$, and $u_{k}$ are unit vectors of length $n_{a},n_{b}$ and $n_{c}$ respectively. Moreover, $L_{A}^{i^{\prime}},L_{B}^{j^{\prime}}$ and $L_{C}^{k^{\prime}}$ are logical operators of the code defined by the parity check matrix $\delta_{a}^{T},\delta_{b}^{T}$, and $\delta_{c}^{T}$ respectively.
For $L_{X}^{1}$, $H_{Z}L_{X}^{1}=0$ implies that any Z-type stabilizer from $H_{Z}$ commutes with $L_{X}^{1}$. Meanwhile, any X-type stabilizer from $H_{Z}$ also commutes with $L_{X}^{1}$ as their only composition of Pauli X operators. To identify logical X operators, we need to find Pauli X operators in the set $L_{X}^{1}$ that are not present in the image of $H_{X}^{T}$. We denote:
$L_{X}^{1,i,j,k}=\left(\begin{array}[]{c}L_{A}^{i^{\prime}}\otimes w_{j}\otimes u_{k}\\ 0_{n_{a}r_{b}n_{c}}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right).$ (94)
For $b_{j},c_{k}\in\{0,1\}$, suppose that the linear combination of these operators
$\sum_{jk}b_{j}c_{k}L_{X}^{1,i,j,k}=\left(\begin{array}[]{c}L_{A}^{i^{\prime}}\otimes\sum_{j}b_{j}w_{j}\otimes\sum_{k}c_{k}u_{k}\\ 0_{n_{a}r_{b}n_{c}}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right)$ (98)
is in the image of matrix $H_{X}^{T}$. Then, there exists vectors $y,z$ such that
$\sum_{jk}b_{j}c_{k}L_{X}^{1,i,j,k}$ $=$ $\left(\begin{array}[]{c}L_{A}^{i^{\prime}}\otimes\sum_{j}b_{j}w_{j}\otimes\sum_{k}c_{k}u_{k}\\ 0_{n_{a}r_{b}n_{c}}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right)$ (102)
$=$ $H_{X}^{T}(L_{A}^{i^{\prime}}\otimes y\otimes z)$ (103)
Recall that $k_{b}^{T}$ and $k_{c}^{T}$ represent the number of encoded bits for $\delta_{b}^{T}$ and $\delta_{c}^{T}$ respectively. Then with $\dim(\text{im}(\delta_{b}^{T}))=n_{b}-k_{b}$ and $\dim(\text{im}(\delta_{c}^{T}))=n_{c}-k_{c}$, one can select at least $k_{b}$ unit vectors $w_{j}$ and $k_{c}$ unit vectors $u_{k}$ such that any combination of these vectors remains outside the image of $\delta_{b}^{T}$ and $\delta_{c}^{T}$ respectively.
Given that there are $k_{a}^{T}$ different $L_{A}^{i^{\prime}}$, and paralleling our analysis of logical Z operators, we can verify that there exist $k_{a}^{T}k_{b}k_{c}$ distinct $L_{X}^{1,i,j,k}$, $k_{a}k_{b}^{T}k_{c}$ distinct $L_{X}^{2,i,j,k}$, and $k_{a}k_{b}k_{c}^{T}$ distinct $L_{X}^{3,i,j,k}$ that serve as logical X operators in $\ker(H_{Z})\backslash\text{im}(H_{X}^{T})$, we call those logical opertors as elementary canonical logical X operators. This accounts for all $k_{a}^{T}k_{b}k_{c}+k_{a}k_{b}^{T}k_{c}+k_{a}k_{b}k_{c}^{T}$ logical X operators of the code.
V.3 Energy barrier of Z logical operators
We now analyze the energy barrier for logical Z operators in 3D hypergraph product codes. First, we employ a strip argument to establish an upper bound. Then, using Lemma 11, we derive a lower bound for the energy barrier for logical Z operators.
V.3.1 Upper bound: “strip” argument
The upper bound on the energy barrier can be obtained by considering a specific path from the identity to any logical operator. By choosing this logical operator to be an elementary canonical logical operator, and consider the “strip” argument mentioned in Sec. IV.2, one can obtain an upper bound.
For elementary canonical logical operator in the form of $L_{Z}^{1,i,j,k}$, with fixed $i,j$, and $k$, the logical operator takes the form as in the Eq. (60). Since $v_{i}$ is a unit vector, the effective logical operator simplifies to $L_{Z}^{1^{\prime},j,k}=L_{B}^{j}\otimes L_{C}^{k}$. This can be interpreted exactly as a logical operator of a tensor product code. One can construct the same kind of strip configuration when implementing these logical operators. Thus, the energy barrier of this logical operator is upper bounded by the strip argument as $\Delta(L_{Z}^{1^{\prime},j,k})\leqslant\min\left(E_{b}d_{c},E_{c}d_{b}\right)$, where $E_{b}$ and $E_{c}$ are the energy barriers of the logical operators $L_{B}^{j}$ and $L_{C}^{k}$, respectively, and $d_{b}$ and $d_{c}$ are their corresponding distances.
With the same argument, the energy barrier of the canonical logical Z operator $L_{Z}^{2,i,j,k}$ is upper bounded by $\min\left(E_{a}d_{c},E_{c}d_{b}\right)$, and the energy barrier of $L_{Z}^{3,i,j,k}$ is upper bounded by $\min\left(E_{a}d_{b},E_{b}d_{a}\right)$.
In summary, consider a 3D hypergraph product code $\mathcal{C}$ constructed from classical codes $\delta_{a}$, $\delta_{b}$, and $\delta_{c}$ with parameters $\left[n_{\ell},k_{\ell},d_{\ell},E_{\ell}\right]$, where $\ell\in\{a,b,c\}$. Based on the structure of the canonical logical operators we have derived, the energy barrier of the logical Z operators can be shown to have the following upper bound.
$\Delta(L_{Z})\leqslant\min_{l,m\in\{a,b,c\}}(d_{l}E_{m}).$ (104)
V.3.2 Lower bound
We now establish a lower bound for the energy barrier of Z-type logical operators by analyzing the structure of these canonical logical operators. Consider a 3D hypergraph product code $\mathcal{C}$ constructed from classical codes $\delta_{a}$, $\delta_{b}$, and $\delta_{c}$ with parameters $\left[n_{\ell},k_{\ell},d_{\ell},E_{\ell}\right]$, where $\ell\in{a,b,c}$. The parity check matrices of this quantum code are $H_{Z}=\partial_{0}^{T}$ and $H_{X}=\partial_{1}$, as defined in Eq. (52) and Eq. (53). Recall that the Z-type canonical logical operators $L_{Z}$ fall into three categories.
$L_{Z}^{1}=\sum_{i,j,k}\alpha_{ijk}((v_{i}\otimes L_{B}^{j}\otimes L_{C}^{k})^{T},0^{T},0^{T})^{T},$ (105)
$L_{Z}^{2}=\sum_{i,j,k}\beta_{ijk}(0^{T},(L_{A}^{i}\otimes w_{j}\otimes L_{C}^{k})^{T},0^{T})^{T},$
$L_{Z}^{3}=\sum_{i,j,k}\gamma_{ijk}(0^{T},0^{T},(L_{A}^{i}\otimes L_{B}^{j}\otimes u_{k})^{T})^{T}.$
where (i) $\delta_{a}L_{A}^{i}=\delta_{b}L_{B}^{j}=\delta_{c}L_{C}^{k}=0$ and (ii) $v_{i}\notin\operatorname{Im}\left(\delta_{a}\right)$, $w_{j}\notin\operatorname{Im}\left(\delta_{b}\right)$ and $u_{k}\notin\operatorname{Im}\left(\delta_{c}\right)$ are unit vectors. We note that $i\in\{1,\ldots,k_{a}^{T}\}$, $j\in\{1,\ldots,k_{b}^{T}\}$ and $k\in\{1,\ldots,k_{c}^{T}\}$ and that the set of logical operators expressible in this form is complete. A canonical logical operator is elementary if only one of the coefficients, $\alpha_{ijk}$ (also $\beta_{ijk}$ or $\gamma_{ijk}$), is equal to one.
The Tanner graph representation provides an intuitive framework that is well-suited for our analysis. Let $\mathcal{G}_{1}(V_{1},C_{1})$, $\mathcal{G}_{2}(V_{2},C_{2})$ and $\mathcal{G}_{3}(V_{3},C_{3})$ be the Tanner graphs of codes defined by $\delta_{a}$, $\delta_{b}$ and $\delta_{c}$, respectively. Here, $V_{1}$, $V_{2}$ and $V_{3}$ represent the set of bits, and $C_{1}$, $C_{2}$ and $C_{3}$ denote the set of checks. We use $\{v_{1}^{i}:i\in\{1,2,\cdots n_{a}\}\}$ (resp. $\{v_{2}^{j}:j\in\{1,2,\cdots n_{b}\}\}$, $\{v_{3}^{k}:j\in\{1,2,\cdots n_{c}\}\}$) to refer to bit vertices in $V_{1}$ (resp. $V_{2}$, $V_{3}$), and also as length $n_{a}$ (resp. $n_{b}$, $n_{c}$) unit vectors with the $i$th (resp. $j$th, $k$th) entry as $1$.
The 3D hypergraph product $\mathcal{G}_{1}\times\mathcal{G}_{2}\times\mathcal{G}_{3}$ is a bipartite graph with vertex set $V\cup C$, where
$V=V_{1}\otimes V_{2}\otimes C_{3}\cup V_{1}\otimes C_{2}\otimes V_{3}\cup C_{1}\otimes V_{2}\otimes V_{3}$ (106)
is the qubit set, and
$C$ $=$ $V_{1}\otimes V_{2}\otimes V_{3}\cup C_{1}\otimes C_{2}\otimes V_{3}\cup C_{1}\otimes V_{2}\otimes C_{3}$ (107)
$\cup V_{1}\otimes C_{2}\otimes C_{3}\cup C_{1}\otimes C_{2}\otimes C_{3}$
is the check and meta-check set. More specifically, $V_{1}\otimes V_{2}\otimes V_{3}$ is the set the Z type checks, $C_{1}\otimes C_{2}\otimes V_{3}\cup C_{1}\otimes V_{2}\otimes C_{3}\cup V_{1}\otimes C_{2}\otimes C_{3}$ represents the set of X checks, and $C_{1}\otimes C_{2}\otimes C_{3}$ is meta-checks on X checks.
The set of qubits can be partitioned into three subsets, $V_{1}\otimes V_{2}\otimes C_{3}$, $V_{1}\otimes C_{2}\otimes V_{3}$, and $C_{1}\otimes V_{2}\otimes V_{3}$. For $H_{X}=\partial_{1}$, there are three column blocks (see Eq. (53)). The first column block of $H_{X}$ acts on $C_{1}\otimes V_{2}\otimes V_{3}$, the second column block of $H_{X}$ acts on $V_{1}\otimes C_{2}\otimes V_{3}$, and the third column block acts on $V_{1}\otimes V_{2}\otimes C_{3}$. Moreover, the subset $C_{1}\otimes V_{2}\otimes V_{3}$ can be further partitioned into $n_{a}^{T}\equiv r_{a}$ subsets $\{c_{1}^{1}\otimes V_{2}\otimes V_{3},c_{1}^{2}\otimes V_{2}\otimes V_{3},\cdots,c_{1}^{r_{a}}\otimes V_{2}\otimes V_{3}\}$, where $c_{1}^{k}\otimes V_{2}\otimes V_{3}:=\{c_{1}^{k}\otimes y\otimes z:y\in V_{2},z\in V_{3}\}$ and $V_{2}=\{v_{2}^{1},\ldots,v_{2}^{n_{b}}\}$, $V_{3}=\{v_{3}^{1},\ldots,v_{3}^{n_{c}}\}$.
Thus, a Z-type Pauli operator can be expressed as a bit-string $z=\left({z^{(1)}}^{T},{z^{(2)}}^{T},{z^{(3)}}^{T}\right)^{T}$, where $z^{(1)}$ is supported on the qubit subset $C_{1}\otimes V_{2}\otimes V_{3}$ with vector space $\mathbb{F}_{2}^{r_{a}}\otimes\mathbb{F}_{2}^{n_{b}}\otimes\mathbb{F}_{2}^{n_{c}}$, $z^{(2)}$ is supported on the qubit set $V_{1}\otimes C_{2}\otimes V_{3}$ with vector space $\mathbb{F}_{2}^{n_{a}}\otimes\mathbb{F}_{2}^{r_{b}}\otimes\mathbb{F}_{2}^{n_{c}}$ and $z^{(3)}$ is supported on the qubit set $V_{1}\otimes V_{2}\otimes C_{3}$ with vector space $\mathbb{F}_{2}^{n_{a}}\otimes\mathbb{F}_{2}^{n_{b}}\otimes\mathbb{F}_{2}^{r_{c}}$. Because $z^{(1)}$, $z^{(2)}$, and $z^{(3)}$ act as a tensor product of three vector spaces, one can view them also as 3D matrices. (For instance, $v_{i}\otimes u_{j}\otimes w_{k}$ for unit vectors $v_{i}$, $u_{j}$ and $w_{k}$ can be viewed as a 3D matrix whose entry is $1$ on the $i$’th a-slice, the $j$’th b-slice and the $k$’th c-slice and zero elsewhere. We call this procedure vector reshaping. After reshaping, $z^{(1)}$, $z^{(2)}$ and $z^{(3)}$ become $Z^{(1)}$, $Z^{(2)}$ and $Z^{(3)}$ respectively. Here, $Z^{(1)}$ is an $r_{a}\times n_{b}\times n_{c}$ matrix, and the entry $Z^{(1)}_{i,j,k}$ is supported on the qubit $r_{1}^{i}\otimes v_{2}^{j}\otimes v_{3}^{k}$. Moreover, the $i$th a-slice $Z^{(1)}_{i}$ is supported on qubits $\{r_{1}^{i}\otimes v_{2}^{1}\otimes v_{3}^{1},r_{1}^{i}\otimes v_{2}^{2}\otimes v_{3}^{1},\cdots,r_{1}^{i}\otimes v_{2}^{n_{b}}\otimes v_{3}^{n_{c}}\}$, which we refer to as $r_{1}^{i}\otimes V_{2}\otimes V_{3}$. Similarly, the subset $V_{1}\otimes C_{2}\otimes V_{3}$ can be partitioned into $r_{b}$ b-slices, and $j$th b-slice is supported on the qubit subset $V_{1}\otimes r_{2}^{j}\otimes V_{3}$. Also, the subset $V_{1}\otimes V_{2}\otimes C_{3}$ can be partitioned into $r_{c}$ c-slices, and $k$th c-slice is supported on the qubit subset $V_{1}\otimes V_{2}\otimes r_{3}^{k}$ (defined similarly).
An elementary canonical logical-Z operator, for example, in the form $(v_{i}\otimes L_{B}^{j}\otimes L_{C}^{k},0,0)^{T}$, is supported on the subset $r_{1}^{i}\otimes V_{2}\otimes V_{3}$ if $v_{i}=r_{1}^{i}$. In matrix form, it is supported on the $i$th a-slice of $Z^{(1)}$. Fig. 2 provides an illustrative example of the 3D Toric code structure.
Figure 2: The 3D toric code can be structured as a cubic lattice (with periodic boundary condition). From left to right, the figure illustrates its decomposition into a-slices, b-slices, and c-slices. The 3D Toric code encodes 3 logical qubits, with corresponding 3 logical Z operators. Each Z elementary canonical logical operator is supported on a single slice, with slices of the same color representing logical operators equivalent up to stabilizers. Generally, given any 3D hypergraph product code, one can create such a decomposition, although the structure within each slice will be more complicated and depends on the underlying classical codes.
We now prove a lower bound for the energy barrier of the canonical logical-Z operators, as shown in the following proposition.
Proposition 13.
For any nontrivial canonical logical-Z operator $L$,
$\Delta(L)\geq\min(d_{a},d_{b},d_{c})$ (108)
We introduce two lemmas to prove this proposition. We use the following convention in the proof. A path $r=\left\{P_{0},P_{1},\ldots,P_{F}\right\}$ is said to be supported on the $U\subseteq V$ if the support of all $P_{i}$ in $r$ is included in $U$.
Lemma 14.
For any elementary canonical logical-Z operator $L$ supported on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$ (resp. $V_{1}\otimes r_{2}^{\beta}\otimes V_{3}$, $V_{1}\otimes V_{2}\otimes r_{3}^{\gamma}$), its energy barrier is attained by a path supported on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$ (resp. $V_{1}\otimes r_{2}^{\beta}\otimes V_{3}$, $V_{1}\otimes V_{2}\otimes r_{3}^{\gamma}$).
Lemma 15.
For any nontrivial canonical logical-Z operator $L$, the energy barrier $\Delta(L)$ is greater than or equal to the minimum energy barrier of the elementary canonical logical-Z operators.
Consider an elementary canonical logical-Z operator $L$ supported on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$. Suppose $\Delta(L)$ is given by a path $r$. The main idea behind the proof of Lemma 14 is to deform a general path $r$ into a new path $r^{\prime}$, supported on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$. Importantly, we show that the energy barrier of $r^{\prime}$ is not greater than that of the original path $r$. A similar argument can be applied to prove the Lemma 15. The proofs closely extended from the arguments in the hypergraph product case Zhao et al. (2025). The details are provided in the appendix.
For the elementary canonical logical operator $L_{Z}^{1,\alpha}=((r_{1}^{\alpha}\otimes L_{B}^{j}\otimes L_{C}^{k})^{T},0^{T},0^{T})^{T}$, with fixed $\alpha$, $j$, and $k$. $L_{Z}^{1,\alpha}$ is supported on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$. To compute the energy barrier of $L_{Z}^{1,\alpha}$, consider a path $r=\{P_{0},P_{1},\cdots,P_{F}\}$. Then
$\Delta(L_{Z}^{1,\alpha})=\min_{r\in w(0,L_{Z}^{1,\alpha})}\epsilon_{\max}(r).$ (109)
Here, $P_{0}=(0,0,0)^{T}$ is a zero vector, and $P_{F}=((r_{1}^{\alpha}\otimes L_{B}^{j}\otimes L_{C}^{k})^{T},0^{T},0^{T})^{T}$. Note that $r$ is a path that walks throughout the set of qubits. Therefore, $P_{\ell}\in r$ could be represented as $P_{\ell}=((P_{\ell}^{1})^{T},(P_{\ell}^{2})^{T},(P_{\ell}^{3})^{T})^{T}$, where each component, $P_{\ell}^{1},P_{\ell}^{2}$, and $P_{\ell}^{3}$, could be nontrivial.
Lemma 14 implies that there is a path $r$ supported on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$ that attains the energy barrier of $\Delta(L)$. A path restricted to the first component means that for any $P_{\ell}\in r$, one can set $P_{\ell}^{2}=P_{\ell}^{3}=0$, the only nontrivial Paulis (nonzero binaries) are in $P_{\ell}^{1}$ with fixed $r_{1}^{\alpha}$. Furthermore, $r_{1}^{\alpha}$ is fixed means every $P_{\ell}^{1}$ has the form $r_{1}^{\alpha}\otimes y\otimes z$, where $y\in V_{2}$ and $z\in V_{3}$ are vectors of length $n_{b}$ and $n_{c}$ respectively.
Given a $P_{\ell}\in r$, the energy (the number of violated checks) of $P_{\ell}$ is given by $\epsilon(P_{\ell})=\mathrm{wt}\left(H_{X}P_{\ell}\right)$, has the form
[…]
By setting $P_{\ell}^{2}=P_{\ell}^{3}=0$, then
$\epsilon(P_{\ell})$ $=\mathrm{wt}\left(\left(\begin{array}[]{c}I_{n_{a}}\otimes\delta_{b}\otimes I_{n_{c}}\\ I_{n_{a}}\otimes I_{n_{b}}\otimes\delta_{c}\end{array}\right)P_{\ell}^{1}\right)$ (118)
$=\mathrm{wt}\left(I_{n_{a}}\otimes\delta_{b}\otimes I_{n_{c}}P_{\ell}^{1}\right)+\mathrm{wt}\left(I_{n_{a}}\otimes I_{n_{b}}\otimes\delta_{c}P_{\ell}^{1}\right).$
Furthermore, for fixed $r_{1}^{\alpha}$, the part $r_{1}^{\alpha}$ can be discarded by applying the map $(r_{1}^{\alpha})^{T}\otimes I_{n_{b}}\otimes I_{n_{c}}$ to $P_{\ell}^{1}$. Thus
$\epsilon(P_{\ell})=\mathrm{wt}\left(\delta_{b}\otimes I_{n_{c}}P_{\ell}^{1^{\prime}}\right)+\mathrm{wt}\left(I_{n_{b}}\otimes\delta_{c}P_{\ell}^{1^{\prime}}\right),$ (119)
with $P_{\ell}^{1^{\prime}}=((r_{1}^{\alpha})^{T}\otimes I_{n_{b}}\otimes I_{n_{c}})P_{\ell}^{1}$.
In summary, to calculate the energy barrier of the Z-type logical operator $L_{Z}^{1,\alpha,j,k}=((r_{1}^{\alpha}\otimes L_{B}^{j}\otimes L_{C}^{k})^{T},0^{T},0^{T})^{T}$, one can consider a path $r\in\{Q_{0},Q_{1},\cdots,Q_{L}\}$ supported on qubit subset $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$, then
$\Delta(L_{Z}^{1,\alpha,j,k})=\min_{r\in w(0,L_{B}^{j}\otimes L_{C}^{k})}\epsilon_{\max}(r).$ (120)
where for each $r$,
$\epsilon_{\max}(r)=\max\{\epsilon(Q_{0}),\epsilon(Q_{1}),\cdots,\epsilon(Q_{L})\},$ (121)
and the energy cost of the step $Q_{\ell}$ is
$\epsilon(Q_{\ell})=\mathrm{wt}\left(\delta_{b}\otimes I_{n_{c}}Q_{\ell}\right)+\mathrm{wt}\left(I_{n_{b}}\otimes\delta_{c}Q_{\ell}\right).$ (122)
This scenario corresponds exactly to computing the energy barrier of a classical tensor product code (see Eq. 42), where the underlying classical codes are $\delta_{b}$ and $\delta_{c}$. The subset $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$ is the bit set of this classical tensor product code. The path for applying logical operator $L_{Z}^{1,\alpha,j,k}$ effectively forms a path for the logical operator $L_{B}\otimes L_{C}$ in the tensor product code constructed by $\delta_{b}$ and $\delta_{c}$.
As stated in Lemma 11, the energy barrier of $L_{B}\otimes L_{C}$ is lower bounded by $\min\left(d_{b},d_{c}\right)$. Thus the energy barrier of the Z-type logical operator $L_{Z}^{1,\alpha,j,k}$ satisfies $\Delta\left(L_{Z}^{1,\alpha,j,k}\right)\geq$ $\min\left(d_{b},d_{c}\right)$.
Following the same reasoning, the energy barrier for logical Z-type operators of the form $L_{Z}^{2,i,j,k}=$ $\left(0^{T},(L_{A}^{i}\otimes w_{j}\otimes L_{C}^{k})^{T},0^{T}\right)^{T}$ is bounded below by $\min\left(d_{a},d_{c}\right)$. Similarly, logical operators of the form $L_{Z}^{3,i,j,k}=\left(0^{T},0^{T},(L_{A}^{i}\otimes L_{B}^{j}\otimes u_{k})^{T}\right)^{T}$ have an energy barrier lower bounded by $\min\left(d_{a},d_{b}\right)$.
Thus, for any elementary canonical Z logical operator $L$ in the 3D hypergraph product code, the energy barrier is lower bounded by
$\Delta(L)\geqslant\min\left(d_{a},d_{b},d_{c}\right).$ (123)
Using Lemma 15, we can establish that for any canonical Z-type logical operator, the energy barrier is lower bounded by $\min\left(d_{a},d_{b},d_{c}\right)$. Moreover, results from Ref. Zhao et al. (2025) demonstrate that for any logical operator $L$ and stabilizer $S$ with $\Delta(L)\geqslant\Delta(S)$, we have $\Delta(L)=\Delta(LS)$. Therefore, with the LDPC property yielding $\Delta(S)=O(1)$ for any $S$, and in cases where $d_{a},d_{b},d_{c}\geqslant O(1)$, this lower bound applies to all logical Z operators, not just the canonical ones.
Formally, we have the following theorem.
Theorem 16.
Let $\Delta(L_{Z})$ be the energy barrier of the Z logical operators of the 3D LDPC hypergraph product code constructed from classical codes $\delta_{a}$, $\delta_{b}$ and $\delta_{c}$ with parameters $[n_{\ell},k_{\ell},d_{\ell},E_{\ell}]$ for $\ell\in{a,b,c}$. If the resulting quantum code is LDPC with sparsity parameters $w_{c},w_{q}$, then in cases where $d_{a},d_{b},d_{c}\geqslant w_{c}w_{q}$, we have
$\Delta(L_{Z})\geqslant\min(d_{a},d_{b},d_{c}).$ (124)
Similarly, with Lemma 10, we can conclude that the energy barrier of 3D LDPC hypergraph product codes is lower bounded by the energy barriers of the underlying classical codes:
$\Delta(L_{Z})\geqslant\min(E_{a},E_{b},E_{c}).$ (125)
Moreover, if Conjecture 12 is true, we have
$\Delta(L_{Z})\geqslant\min_{l,m\in\{a,b,c\}}(d_{l}E_{m}).$ (126)
This result matches the upper bound obtained using the strip argument, effectively providing a tight bound.
V.4 Energy barrier of X logical operators
The canonical X logical operators have the form
$L_{X}^{1}=\sum_{i,j,k}\alpha_{ijk}^{\prime}((L_{A}^{i^{\prime}}\otimes w_{j}\otimes u_{k})^{T},0_{n_{a}r_{b}n_{c}}^{T},0_{n_{a}n_{b}r_{c}}^{T})^{T},$
$L_{X}^{2}=\sum_{i,j,k}\beta_{ijk}^{\prime}(0_{r_{a}n_{b}n_{c}}^{T},(v_{i}\otimes L_{B}^{j^{\prime}}\otimes u_{k})^{T},0_{n_{a}n_{b}n_{c}}^{T})^{T},$
$L_{X}^{3}=\sum_{i,j,k}\gamma_{ijk}^{\prime}(0_{r_{a}n_{b}n_{c}}^{T},0_{n_{a}r_{b}n_{c}}^{T},(v_{i}\otimes w_{j}\otimes L_{C}^{k^{\prime}})^{T})^{T},$
where (i) $\delta_{a}^{T}L_{A}^{i^{\prime}}=\delta_{b}^{T}L_{B}^{j^{\prime}}=\delta_{c}^{T}L_{C}^{k^{\prime}}=0$ and (ii) $v_{i}\notin\operatorname{Im}\left(\delta_{a}^{T}\right)$, $w_{j}\notin\operatorname{Im}\left(\delta_{b}^{T}\right)$ and $u_{k}\notin\operatorname{Im}\left(\delta_{c}^{T}\right)$ are unit vectors. We note that $i\in\{1,\ldots,k_{a}^{T}\}$, $j\in\{1,\ldots,k_{b}\}$ and $k\in\{1,\ldots,k_{c}\}$ and that the set of logical operators expressible in this form is complete. A canonical logical operator is elementary if only one of the coefficients, $\alpha_{ijk}^{\prime}$, $\beta_{ijk}^{\prime}$ or $\gamma_{ijk}^{\prime}$, is equal to one.
These elementary canonical logical X operators take the form of tensor products between one classical logical operator and two unit vectors, which resembles that of logical operators in hypergraph product codes, though in a higher dimension. By extending the reasoning previously applied to hypergraph product codes Zhao et al. (2025) to this higher-dimensional case, in the LDPC regime, one can establish a lower bound for the energy barrier of logical X operators as
$\Delta(L_{X})\geqslant\min\left(E_{a}^{T},E_{b}^{T},E_{c}^{T}\right).$ (128)
The upper bound can be also derived from the structure of these elementary canonical logical X operators Zhao et al. (2025), which is equal to $\left\{E_{a}^{T},E_{b}^{T},E_{c}^{T}\right\}$. In the case where $\min\left(E_{a}^{T},E_{b}^{T},E_{c}^{T}\right)\geqslant w_{c}w_{q}$, with $w_{c},w_{q}$ are the sparsity parameters of the resulting quantum code, then
$\Delta(L_{X})=\min\left(E_{a}^{T},E_{b}^{T},E_{c}^{T}\right).$ (129)
VI Enery barrier of 4D hypergraph product codes
In 3D hypergraph product codes, the energy barrier of Z-type logical operators is lower bounded by the distance of the underlying classical codes. Meanwhile, X logical operators retain the energy barrier of these classical codes. When classical codes possess large distances but small energy barriers, only Z-type logical operators have improved energy barrier, potentially enabling single-shot. The X-type logical operators, lacking this enhanced protection, remain vulnerable. In this section, we argue that 4D hypergraph product codes offer improved energy barrier for both X and Z logical operators.
This distinction between the 3D and 4D cases stems from structural differences in their elementary canonical logical operators. In 3D codes, Z logical operators take the form $L_{Z}=L_{c_{1}}\otimes L_{c_{2}}\otimes u_{k}$, while X logical operators are simply $L_{X}=v_{i}\otimes w_{j}\otimes L_{c_{3}}^{T}$, where $v_{i},w_{j}$, and $u_{k}$ are unit vectors, and $L_{c_{1}},L_{c_{2}},L_{c_{3}}$ are classical logical operators. In contrast, 4D codes exhibit a more balanced structure where both X and Z logical operators are tensor products of two classical logical operators (and two unit vectors, see Eq. 155 and Eq. 156), resulting in enhanced energy barriers for both types of logical operators.
Given four classical codes represented by parity check matrices $\delta_{a},\delta_{b},\delta_{c},\delta_{d}$, let $\left[n_{\ell},k_{\ell},d_{\ell},E_{\ell}\right]$ and $\left[n_{\ell}^{T},k_{\ell}^{T},d_{\ell}^{T},E_{\ell}^{T}\right]$ represent the parameters of the classical linear code defined by the parity check matrix $\delta_{\ell}$ and its transpose $\delta_{\ell}^{T}$. For convenience, we denote $r_{\ell}\equiv n_{\ell}^{T}$, where =$\ell=\{a,b,c,d\}$. The 4D hypergraph product is defined by a length-$5$ chain complex
$C_{0}\xrightarrow{\partial_{0}}C_{1}\xrightarrow{\partial_{1}}C_{2}\xrightarrow{\partial_{2}}C_{3}\xrightarrow{\partial_{3}}C_{4},$ (130)
Then, $\partial_{0},\partial_{1}$, $\partial_{2}$ and $\partial_{3}$ are
$\partial_{0}=\left(\begin{array}[]{c}\delta_{a}\otimes I_{n_{b}}\otimes I_{n_{c}}\otimes I_{n_{d}}\\ I_{n_{a}}\otimes\delta_{b}\otimes I_{n_{c}}\otimes I_{n_{d}}\\ I_{n_{a}}\otimes I_{n_{b}}\otimes\delta_{c}\otimes I_{n_{d}}\\ I_{n_{a}}\otimes I_{n_{b}}\otimes I_{n_{c}}\otimes\delta_{d}\end{array}\right),$ (135)
$\partial_{1}=\left(\begin{array}[]{cccc}I_{r_{a}}\otimes\delta_{b}\otimes I_{n_{c}}\otimes I_{n_{d}}&\delta_{a}\otimes I_{r_{b}}\otimes I_{n_{c}}\otimes I_{n_{d}}&0&0\\ I_{r_{a}}\otimes I_{n_{b}}\otimes\delta_{c}\otimes I_{n_{d}}&0&\delta_{a}\otimes I_{n_{b}}\otimes I_{r_{c}}\otimes I_{n_{d}}&0\\ I_{r_{a}}\otimes I_{n_{b}}\otimes I_{n_{c}}\otimes\delta_{d}&0&0&\delta_{a}\otimes I_{n_{b}}\otimes I_{n_{c}}\otimes I_{r_{d}}\\ 0&I_{n_{a}}\otimes I_{r_{b}}\otimes\delta_{c}\otimes I_{n_{d}}&I_{n_{a}}\otimes\delta_{b}\otimes I_{r_{c}}\otimes I_{n_{d}}&0\\ 0&I_{n_{a}}\otimes I_{r_{b}}\otimes I_{n_{c}}\otimes\delta_{d}&0&I_{n_{a}}\otimes\delta_{b}\otimes I_{n_{c}}\otimes I_{r_{d}}\\ 0&0&I_{n_{a}}\otimes I_{n_{b}}\otimes I_{r_{c}}\otimes\delta_{d}&I_{n_{a}}\otimes I_{n_{b}}\otimes\delta_{c}\otimes I_{r_{d}}\end{array}\right),$ (142)
$\partial_{2}=$
$\left({\begin{array}[]{cccccc}I_{r_{a}}\otimes I_{r_{b}}\otimes\delta_{c}\otimes I_{n_{d}}&I_{r_{a}}\otimes\delta_{b}\otimes I_{r_{c}}\otimes I_{n_{d}}&0&\delta_{a}\otimes I_{r_{b}}\otimes I_{r_{c}}\otimes I_{n_{d}}&0&0\\ I_{r_{a}}\otimes I_{r_{b}}\otimes I_{n_{c}}\otimes\delta_{d}&0&I_{r_{a}}\otimes\delta_{b}\otimes I_{n_{c}}\otimes I_{r_{d}}&0&\delta_{a}\otimes I_{r_{b}}\otimes I_{n_{c}}\otimes I_{r_{d}}&0\\ 0&I_{r_{a}}\otimes I_{n_{b}}\otimes I_{r_{c}}\otimes\delta_{d}&I_{r_{a}}\otimes I_{n_{b}}\otimes\delta_{c}\otimes I_{r_{d}}&0&0&\delta_{a}\otimes I_{n_{b}}\otimes I_{r_{c}}\otimes I_{r_{d}}\\ 0&0&0&I_{n_{a}}\otimes I_{r_{b}}\otimes I_{r_{c}}\otimes\delta_{d}&I_{n-a}\otimes I_{r_{b}}\otimes\delta_{c}\otimes I_{r_{d}}&I_{n_{a}}\otimes\delta_{b}\otimes I_{r_{c}}\otimes I_{r_{d}}\end{array}}\right),$ (148)
$\partial_{3}=\left(\begin{array}[]{cccc}I_{r_{a}}\otimes I_{r_{b}}\otimes I_{r_{c}}\otimes\delta_{d}&I_{r_{a}}\otimes I_{r_{b}}\otimes\delta_{c}\otimes I_{r_{d}}&I_{r_{a}}\otimes\delta_{b}\otimes I_{r_{c}}\otimes I_{r_{d}}&\delta_{a}\otimes I_{r_{b}}\otimes I_{r_{c}}\otimes I_{r_{d}}\end{array}\right).$ (151)
One can confirm that $\partial_{i}\partial_{i-1}=0\mod 2$ for $i\in{1,2,3}$, thus $H_{Z}=\partial_{1}^{T},H_{X}=\partial_{2}$ can be used to construct a quantum CSS code $H_{(\delta_{a},\delta_{b},\delta_{c},\delta_{d})}$ with parity check matrix
$H_{\left(\delta_{a},\delta_{b},\delta_{c},\delta_{c}\right)}=\left(\begin{array}[]{cc}H_{X}&0\\ 0&H_{Z}\end{array}\right).$ (153)
Moreover, as $\partial_{3}\partial_{1}=\partial_{1}\partial_{0}=0$, the remaining matrices $M_{1}=\partial_{3}$, $M_{2}=\partial_{0}^{T}$ work as a meta-check for $H_{X}$ and $H_{Z}$, respectively. This means that any valid X-syndrome adheres to the constraints defined by $M_{1}$, and any valid Z-syndrome satisfies the constraints defined by $M_{2}$.
Similarly to the 3D case, the parameters of the code $\mathcal{C}\left(\delta_{a},\delta_{b},\delta_{c},\delta_{d}\right)$ depend on the classical codes $\delta_{a},\delta_{b}$, $\delta_{c}$, and $\delta_{d}$. More specific, the above length-$5$ chain complex can be regarded as an $\left[\left[n,k,d_{x},d_{z}\right]\right]$ code, with
$n$ $=$ $n_{a}^{T}n_{b}^{T}n_{c}n_{d}+n_{a}^{T}n_{b}n_{c}^{T}n_{d}+n_{a}^{T}n_{b}n_{c}n_{d}^{T}+n_{a}n_{b}^{T}n_{c}^{T}n_{d}+n_{a}n_{b}^{T}n_{c}n_{d}^{T}+n_{a}n_{b}n_{c}^{T}n_{d}^{T},$
$k$ $=$ $k_{a}^{T}k_{b}^{T}k_{c}k_{d}+k_{a}^{T}k_{b}k_{c}^{T}k_{d}+k_{a}^{T}k_{b}k_{c}k_{d}^{T}+k_{a}k_{b}^{T}k_{c}^{T}k_{d}+k_{a}k_{b}^{T}k_{c}k_{d}^{T}+k_{a}k_{d}k_{c}^{T}k_{d}^{T},$
$d_{x}$ $=$ $\min\left(d_{a}^{T}d_{b}^{T},d_{c}^{T}d_{a}^{T},d_{a}^{T}d_{d}^{T},d_{c}^{T}d_{b}^{T},d_{b}^{T}b_{d}^{T},d_{c}^{T}d_{d}^{T}\right),$
$d_{z}$ $=$ $\min\left(d_{b}d_{c},d_{b}d_{d},d_{a}d_{d},d_{a}d_{c},d_{a}d_{b}\right).$ (154)
The Z-type canonical logical operators $L_{Z}$ can be categorized into the following six sets.
$L_{Z}^{1}=\sum_{i,j,k,l}a^{1}_{ijkl}((v_{i}\otimes w_{j}\otimes L_{C}^{k}\otimes L_{D}^{l})^{T},0^{T},0^{T},0^{T},0^{T},0^{T})^{T},$ (155)
$L_{Z}^{2}=\sum_{i,j,k,l}a^{2}_{ijkl}(0^{T},(v_{i}\otimes L_{B}^{j}\otimes u_{k}\otimes L_{D}^{l})^{T},0^{T},0^{T},0^{T},0^{T})^{T},$
$L_{Z}^{3}=\sum_{i,j,k,l}a^{3}_{ijkl}(0^{T},0^{T},(v_{i}\otimes L_{B}^{j}\otimes L_{C}^{k}\otimes t_{l})^{T},0^{T},0^{T},0^{T})^{T},$
$L_{Z}^{4}=\sum_{i,j,k,l}a^{4}_{ijkl}(0^{T},0^{T},0^{T},(L_{A}^{i}\otimes w_{j}\otimes u_{k}\otimes L_{D}^{l})^{T},0^{T},0^{T})^{T},$
$L_{Z}^{5}=\sum_{i,j,k,l}a^{5}_{ijkl}(0^{T},0^{T},0^{T},0^{T},(L_{A}^{i}\otimes w^{j}\otimes L_{C}^{k}\otimes t_{l})^{T},0^{T})^{T},$
$L_{Z}^{6}=\sum_{i,j,k,l}a^{6}_{ijkl}(0^{T},0^{T},0^{T},0^{T},0^{T},(L_{A}^{i}\otimes L_{B}^{j}\otimes u_{k}\otimes t_{l})^{T})^{T},$
where $v_{i},w_{j}$, $u_{k}$, and $t_{l}$ are unit vectors with length $r_{a},r_{b}$, $r_{c}$ and $r_{c}$, respectively. ${L_{A}^{i}},{L_{B}^{j}}$, ${L_{C}^{k}}$ and ${L_{D}^{l}}$ are logical operators of the code defined by the parity check matrix $\delta_{a},\delta_{b}$, $\delta_{c}$ and $\delta_{d}$, respectively. Given that $\delta_{a}L_{A}^{i}=\delta_{b}L_{B}^{j}=\delta_{c}L_{C}^{k}=\delta_{d}L_{D}^{l}=0$, one can confirmed that $H_{X}L_{Z}^{l}=0$ for $l\in\{1,2,3,4,5,6\}$.
Similarly, the X-type canonical logical operators $L_{X}$ can be categorized into six sets.
$L_{X}^{1}=\sum_{i,j,k,l}b^{1}_{ijkl}(({L_{A}^{i}}^{\prime}\otimes{L_{B}^{j}}^{\prime}\otimes u_{k}\otimes t_{l})^{T},0^{T},0^{T},0^{T},0^{T},0^{T})^{T},$ (156)
$L_{X}^{2}=\sum_{i,j,k,l}b^{2}_{ijkl}(0^{T},({L_{A}^{i}}^{\prime}\otimes w^{j}\otimes{L_{C}^{k}}^{\prime}\otimes t_{l})^{T},0^{T},0^{T},0^{T},0^{T})^{T},$
$L_{X}^{3}=\sum_{i,j,k,l}b^{3}_{ijkl}(0^{T},0^{T},({L_{A}^{i}}^{\prime}\otimes w_{j}\otimes u_{k}\otimes{L_{D}^{l}}^{\prime})^{T},0^{T},0^{T},0^{T})^{T},$
$L_{X}^{4}=\sum_{i,j,k,l}b^{4}_{ijkl}(0^{T},0^{T},0^{T},(v_{i}\otimes{L_{B}^{j}}^{\prime}\otimes{L_{C}^{k}}^{\prime}\otimes t_{l})^{T},0^{T},0^{T})^{T},$
$L_{X}^{5}=\sum_{i,j,k,l}b^{5}_{ijkl}(0^{T},0^{T},0^{T},0^{T},(v_{i}\otimes{L_{B}^{j}}^{\prime}\otimes u_{k}\otimes{L_{D}^{l}}^{\prime})^{T},0^{T})^{T},$
$L_{X}^{6}=\sum_{i,j,k,l}b^{6}_{ijkl}(0^{T},0^{T},0^{T},0^{T},0^{T},(v_{i}\otimes w_{j}\otimes{L_{C}^{k}}^{\prime}\otimes{L_{D}^{l}}^{\prime})^{T})^{T}.$
where $v_{i},w_{j}$, $u_{k}$ and $t_{l}$ are unit vectors of length $n_{a},n_{b}$, $n_{c}$ and $n_{d}$, respectively. ${L_{A}^{i}}^{\prime},{L_{B}^{j}}^{\prime}$, ${L_{C}^{k}}^{\prime}$ and ${L_{D}^{l}}^{\prime}$ are logical operators of the code defined by the parity check matrix $\delta_{a}^{T},\delta_{b}^{T}$, $\delta_{c}^{T}$ and $\delta_{d}^{T}$, respectively.
Both X and Z elementary canonical logical operators have the structure of tensor product of two classical logical operators (and two unit vectors). Thus, in the LDPC regime, with Lemma 11, by applying the same reasoning used for logical Z operators in 3D hypergraph product codes but extended to one more dimension, we can establish the following lower bounds on the energy barriers for logical Z and X operators in the 4D hypergraph product code.
$\Delta(L_{Z})$$\geqslant$ $\min(d_{a},d_{b},d_{c},d_{d}),$
$\Delta(L_{X})$$\geqslant$ $\min(d_{a}^{T},d_{b}^{T},d_{c}^{T},d_{d}^{T}).$ (157)
We state it formally in the following Theorem
Theorem 17.
Let $\Delta(L_{Z})$, $\Delta(L_{X})$ be the energy barrier of the Z, X logical operators of the 4D hypergraph product code constructed from classical codes $\delta_{a}$, $\delta_{b}$, $\delta_{c}$ and $\delta_{d}$, with parameters $[n_{\ell},k_{\ell},d_{\ell},E_{\ell}]$ for $\ell\in\{a,b,c,d\}$. If the resulting quantum code is LDPC with sparsity parameters $w_{c},w_{q}$. When $d_{\ell},d_{\ell}^{T}\geqslant w_{c}w_{q}$ for $\ell\in\{a,b,c,d\}$, we have
$\Delta(L_{Z})$$\geqslant$ $\min(d_{a},d_{b},d_{c},d_{d}),$ (158)
$\Delta(L_{X})$$\geqslant$ $\min(d_{a}^{T},d_{b}^{T},d_{c}^{T},d_{d}^{T}).$ (159)
Similarly, with Lemma 10, we have
$\Delta(L_{Z})$$\geqslant$ $\min(E_{a},E_{b},E_{c},E_{d}),$ (160)
$\Delta(L_{X})$$\geqslant$ $\min(E_{a}^{T},E_{b}^{T},E_{c}^{T},E_{d}^{T}).$ (161)
Moreover, if Conjecture 12 is true, we have
$\Delta(L_{Z})$$\geqslant$ $\min_{l,m\in\{a,b,c,d\}}(d_{l}E_{m}),$ (162)
$\Delta(L_{X})$$\geqslant$ $\min_{l,m\in\{a,b,c,d\}}(d_{l}^{T}E_{m}^{T}),$ (163)
which is tight.
VII Conclusion and outlook
In this work, we gave a comprehensive introduction to the confinement, soundness, and expansion properties of codes. We also discussed the relationships among these properties. Specifically, we showed the equivalence between left-expansion and linear confinement, and demonstrated formally that confinement property inherently induces its energy barrier. We then established a lower bound for the energy barrier of tensor product codes in Lemma 11, showing its relation to the distance of the underlying codes.
Using this lemma, we proved that for 3D LDPC hypergraph product codes constructed from classical codes $\delta_{a}$, $\delta_{b}$, and $\delta_{c}$ with parameters $\left[n_{\ell},k_{\ell},d_{\ell},E_{\ell}\right]$, where $\ell\in{a,b,c}$, the energy barrier of logical Z operators is lower-bounded by $d_{\ell}$, which improves upon the $O(d_{\ell}^{\frac{1}{3}})$ bound derived from the confinement property alone. For logical X operators, the energy barrier corresponds to that of the transpose codes of these classical codes, paralleling the behavior in standard hypergraph products.
In the 4D case, we established that both X and Z logical operators have energy barriers lower-bounded by $d_{\ell}$ and $d_{\ell}^{T}$, respectively. This represents a significant improvement over previous results derived from Ref. Campbell (2019); Quintavalle et al. (2021), where under the same conditions, the energy barriers derived from soundness and confinement conditions are only $O(d_{\ell}^{\frac{1}{3}})$ and $O((d_{\ell}^{T})^{\frac{1}{3}})$, respectively.
Several questions remain for further investigation. An immediate inquiry concerns the mechanism behind energy barrier enhancement in HHGP codes. A possible explanation lies in the increased check redundancy introduced by these higher-dimensional structures. This raises follow-up questions, what is the precise quantitative relation between redundancy and energy barrier? What amount or structure of redundancy is sufficient to achieve a specified energy barrier?
Researchers have shown that the redundancy provided by higher-dimensional hypergraph products is enough for soundness and confinement Campbell (2019); Quintavalle et al. (2021). Several works also demonstrate the effectiveness of a redundant set of parity checks for simultaneously handling measurement and qubit errors Fujiwara (2014); Ashikhmin et al. (2020); Delfosse et al. (2022). However, it remains an open question how to establish a general framework for introducing redundancy that is sufficient for confinement across different code families.
Additionally, quantum expander codes present an intriguing example, as they exhibit confinement despite lacking check redundancy. This inspires the thought to explore alternative methods of providing confinement beyond redundancy or expansion, such as symmetries Roberts and Bartlett (2020); Stahl (2023). Moreover, quantum expander codes exhibit unique statistical properties, as demonstrated in Ref. Hong et al. (2025); Yin and Lucas (2024).
Moreover, quantum codes with large energy barriers are conjectured to support local decoders Leverrier et al. (2015); Fawzi et al. (2018b); Kubica and Preskill (2019); Breuckmann et al. (2017); Hastings (2014); Dennis et al. (2002); Herold et al. (2015); Breuckmann and Ni (2018); Dauphinais and Poulin (2017); Pastawski et al. (2011); Vasmer et al. (2021), based on intuition from statistical mechanics. In such systems, local cooling processes can reduce system energy, suggesting that for codes with macroscopic energy barriers, a simple iterative local process could function effectively as a decoder. However, the design of local decoders for higher-dimensional hypergraph product codes remains an open research question.
Acknowledgments- I thank Qi Ye of Tsinghua University for providing the argument for the energy barrier of two-dimensional repetition code. I thank Isaac Kim for helpful discussions and feedback on an earlier version of this manuscript. I also thank Andrew Doherty, Xiao-Tong Ni for valuable discussions. The author acknowledges the financial support from Sydney Quantum Academy. This work was supported by the Australian Research Council Centre of Excellence for Engineered Quantum Systems (EQUS, CE170100009).