### v3

# Improved energy barrier in higher-dimensional hypergraph product codes

Abstract
Single-shot error correction outperforms conventional approaches by requiring only one round of stabilizer measurements for decoding, even in the presence of measurement errors. This capability relates to the soundness and confinement property of codes, which provides an energy barrier lower bound. Earlier research established a confinement property for 3D hypergraph product codes [Quintavalle et al. 2021 PRX Quantum], yielding an energy barrier lower bound for these codes. In this work, by analyzing the structure of logical operators, we show an improved energy barrier lower bound for higher-dimensional hypergraph product (HHGP) codes with low-density parity check (LDPC) property. Our bound exceeds results derived from soundness and confinement, and unlike standard hypergraph product codes, these higher dimensional variants can possess macroscopic energy barriers even when the underlying classical codes lack this property. Specifically, our results show that the energy barrier of LDPC HHGP codes is lower bounded by the distance of the underlying classical codes. This bound is tight if the underlying classical codes exhibit system size-dependent distances but constant energy barriers, like 3D and 4D toric codes.
I Introduction
Quantum error correction is crucial for fault tolerant quantum computing, yet it introduces significant overhead. Within the stabilizer codes gottesman1997stabilizer implemented through the quantum circuit model nielsen2001quantum, a significant portion of overhead stems from the stabilizer measurements required to detect and identify errors. Specifically, along the quantum circuit, stabilizer measurements and decoding must be performed at regular intervals to maintain the integrity of the code state. Moreover, to overcome measurement errors, conventional protocols require multiple rounds of stabilizer measurements for each decoding process, with the number of rounds typically scaling with the code distance shor1996fault; divincenzo1996fault.
In response to this overhead, single-shot error correction emerges as a promising protocol by requiring only one round of stabilizer measurements per decoding process bombin2015single, even in the presence of measurement errors. Although this approach still leaves some residual error after each decoding, the residual error remains bounded. Such controlled error propagation is sufficient for fault-tolerant quantum computation.
Single-shot error correction requires quantum codes with specialized structures. Bombín first demonstrated that codes with confinement property enable single-shot correction bombin2015single. Campbell later proposed that good soundness is also sufficient for single-shot error correction campbell2019theory. He also showed that single-shot error correction can be realized for any code by measuring a carefully chosen set of stabilizers. Further advancing the theory, Quintavalle et al. provided a more general definition of confinement, and proved that confinement is sufficient for single-shot decoding in the case of adversarial errors Quintavalle2021single.
So far, codes shown to support single-shot error correction include the 3D gauge color code bombin2015single; Ben2016nc, 3D hypergraph product codes (only for either X or Z noise, depending on code structure)kubica2019cellular; vasmer2021cellular; Quintavalle2021single; Higgott2023improved, 4D hypergraph product codesniko2017local; campbell2019theory; zeng2019higher, 4D hyperbolic codes Breuckmann_2022, quantum expander codes fawzi2020constant, and quantum Tanner codes gu2024single.
Among these codes, higher-dimensional hypergraph product (HHGP) codes zeng2019higher are an important family of quantum codes that extend the original hypergraph product construction 2014-Tillich into higher dimensions. While the standard hypergraph product combines only two classical codes to form a quantum code, this higher-dimensional approach incorporates additional classical codes. These additional structure provide meta-checks for X and Z stabilizers, effectively validating the stabilizers themselves, thereby providing an intuitive foundation for the single-shot error correction capability these codes exhibit.
Single-shot error correction shares fundamental connections with self-correcting memory  ben2016rmp; terhal2015quantum, which requires the system to preferentially stay within the code space at finite temperature. A thorough analysis of self-correction requires evaluation of free energy dynamics. While entropic effects play a crucial role in determining the system’s stability plischke1994equilibrium; bravyi2013; Bravyi.2011, a macroscopic energy barrier constitutes a necessary condition for self-correction. The term “macroscopic” means that energy barrier scales with system size.
Previously, the author established a lower bound for the energy barrier of hypergraph product codes with low-density parity check (LDPC) property zhao2024on. In that case, the quantum energy barrier is lower bounded by the energy barriers of the underlying classical codes. Thus, constructing quantum codes with macroscopic energy barriers necessitates classical codes with the same property. However, finding the classical codes with macroscopic energy barriers remains challenging, raising the question of how to construct quantum codes that exhibit macroscopic energy barriers when the classical codes do not.
The HHGP codes indeed meet this criterion. In this work, we will show that the energy barrier of LDPC HHGP codes is lower bounded not only by the energy barriers of the underlying classical codes (as in standard hypergraph products), but also by the distances of these underlying classical codes. Thus to construct quantum codes with macroscopic energy barrier, one can just use classical codes that have large distances.
Notably, previous research has demonstrated that HHGP codes exhibit good soundness and confinement, making them single-shot correctable against adversarial errors campbell2019theory; Quintavalle2021single. These soundness and confinement properties induce a lower bound on the energy barrier. However, we demonstrate that this bound is not optimal. Our approach, based on analyzing logical operator structures, provides a tighter characterization of the energy landscape for HHGP codes.
To guide the reader, our paper is structured as follows. In Sec. II, we review key concepts including soundness, confinement, energy barrier and HHGP codes. In Sec. III, we formally demonstrate how confinement induces a lower bound on energy barriers and use the confinement results from Ref. Quintavalle2021single to establish an energy barrier lower bound for 3D hypergraph product codes. Then, in Sec. IV, we provide an energy barrier lower bound for tensor product codes, as formally stated in Lemma 11. This lower bound serves as a key insight for proving the energy barrier lower bound of LDPC HHGP codes. Finally, by analyzing the structure of logical operators of HHGP codes and applying Lemma 11, we derive energy barrier lower bounds for 3D and 4D hypergraph product codes in Sec. V and Sec. VI, respectively. These bounds, presented in Theorem 16 and Theorem 17, relate directly to the distances of the underlying classical codes.
II Preliminaries
II.1 Quantum LDPC codes
Stabilizer codes, conceptualized by Daniel Gottesman in 1996 gottesman1997stabilizer, encode quantum information in a protected subspace of a larger Hilbert space. Specifically, these codes store quantum states within the +1+1 co-eigenspace of a set of $m$ commuting Pauli operators, $\mathcal{S}=\{S_{1},S_{2},\ldots,S_{m}\}$, defined on an $n$-qubit Hilbert space $\mathcal{H}_{2}^{\otimes n}$. Each operator $S_{i}\in\{I,X,Y,Z\}^{\otimes n}$ consists of a tensor product of the Pauli operators, with the requirement that all these operators commute with each other. $\mathcal{S}$ excludes the operator I-I to ensure a non-trivial code space. Logical operators are defined as the elements of $\S\mathcal{C}(S)\backslash$ S.
Calderbank-Shor-Steane (CSS) codes calderbank1996good; steane1996error are an important family of stabilizer codes with special structural properties. Their stabilizer generators are exclusively either X-type or Z-type Pauli operators. This separation allows all X-type stabilizers to be represented by a parity check matrix $H_{X}$, while all Z-type stabilizers are represented by a parity check matrix $H_{Z}$. Consequently, the parity check matrix for the quantum CSS code takes the form
$H=\left(\begin{array}[]{cc}0&H_{Z}\\ H_{X}&0\end{array}\right),$ (3)
with the condition $H_{X}H_{Z}^{T}=0$ ensuring that all stabilizers commute with each other.
Quantum LDPC codes are stabilizer codes with sparse parity check matrices (see zunaira2015; Breuckmann2021ldpc for good reviews). The sparsity is defined by two parameters: $w_{c}$, the maximum weight amongst all stabilizer generators, and $w_{q}$, the maximum number of stabilizer generators associated with any single qubit. Formally, a code is classified as LDPC if both $w_{c}$ and $w_{q}$ remain constant as the code size increases, specifically $w_{c},w_{q}=O(1)$.
Quantum LDPC codes are promising because they often provide lower overhead and can be designed for fault-tolerance gottesman2014faulttolerant. Recent research has led to several families of quantum LDPC codes with enhanced parameters freedman2002z2; evra2022decodable; kaufman2021new; hastings2021fiber; panteleev2021quantum; Breuckmann_2021; bacon2017sparse, with some efforts even achieving good quantum LDPC codes Panteleev.2021; leverrier2022quantum; dinur2022good.
Various methods exist for constructing quantum LDPC codes. From a mathematical perspective, the hypergraph product construction 2014-Tillich; zeng2019higher represents a well-established approach, providing a systematic way to construct quantum codes from any classical codes. Alternative methods such as two-block group-algebra (2BGA) codes wang2023abelian; lin2024twoblock offer different advantages, like suitability for two-dimensional layouts bravyi2024bbcode, making them useful for practical quantum computing architectures.
II.2 Energy barrier of codes
Following Ref. Bravyi2009nogo, we present a formal definition of the energy barrier for stabilizer codes. Let $\mathcal{C}$ be the code subspace of a stabilizer code, defined in terms of the stabilizer group $S$. The code subspace can be viewed as the ground state subspace of a Hamiltonian of the form $\hat{H}=\sum_{i=1}^{m}(I-s_{i})/2$, where $\{s_{1},\ldots,s_{m}\}\subset S$ is a set of stabilizer generators.
[…]
$\epsilon(P)=\text{wt}(Hv(P)).$ (4)
A sequence $P_{0},P_{1},\ldots,P_{t}$ from the Pauli group $\mathcal{P}$ forms a path from $P_{0}$ to $P_{t}$ if for each index $i$, the operators $P_{i}$ and $P_{i+1}$ differ at no more than one qubit, we call it local error condition. The notation $w(P_{0},P_{t})$ represents the collection of all such paths from $P_{0}$ to $P_{t}$. For $r\in w(P_{0},P_{t})$, $\epsilon_{\max}(r)$ denotes the highest energy along path $r$, i.e., $\epsilon_{\max}(r)=\max_{P_{i}\in r}\epsilon(P_{i})$, as the energy barrier of $E$ along the path $r$.
The minimum energy associated with a Pauli $P$ is the smallest value of $\epsilon_{\text{max}}$ across all possible paths from $I$ to $P$. This is the energy barrier of $P$, denoted as $\Delta(P)$,
$\Delta(P)=\min_{r\in w(I,P)}\epsilon_{\max}(r).$ (5)
The energy barrier of the quantum code is the minimum energy barrier over the set of nontrivial logical operators.
Definition 1.
Let $S$ be a stabilizer group and $L(S)$ be the set of nontrivial logical operators. The energy barrier is
$\Delta(H):=\min\{\Delta(\ell):\ell\in L(S)\}.$ (6)
This is the smallest energy the environment has to overcome to enact a logical operation on the encoded qubit. We can similarly define the energy barrier of classical codes by only considering the path formed by Pauli-Xs. We shall denote this energy barrier also as $\Delta(H)$, where $H$ in this case is the parity check matrix of the classical code.
There are two main obstacles for computing the energy barrier of quantum codes. First, numerous paths exist for implementing any logical operator, creating a vast solution space to explore. Second, in quantum codes, any stabilizer can be multiplied with a logical operator to produce an equivalent logical operator, while for those equivalent logical operators, the energy barriers may differ. That is to say, for a logical operator $L$ and stabilizer $S$, usually $\Delta(L)\neq\Delta(LS)$. This complicates the analysis as one must consider the full equivalence class of logical operators.
Quantum LDPC codes offer significant advantages for energy barrier computation by reducing analytical complexity. As shown in Ref. zhao2024on, in LDPC codes, two logical operators that are equivalent under stabilizers have identical energy barriers, provided at least one of their energy barriers equals or exceeds $w_{c}w_{q}$, the product of sparsity parameters. With $w_{c}w_{q}=O(1)$, this property yields a powerful simplification: if any non-trivial logical operator has an energy barrier of $\Omega(1)$, all equivalent logical operators share the same energy barrier.
Therefore, to determine the energy barriers of quantum LDPC codes, one needs only analyze a fixed complete set of logical operators. Once the energy barrier is established for this complete set, the energy barrier of the code follows directly from these results.
II.3 Soundness
Soundness originated in the study of locally testable codes (LTCs) blum1990self; arora1994probabilistic; rubinfeld1996robust; friedl1995some. LTCs are a class of error-correcting codes that allow efficient verification of whether a string is a valid codeword or is far from the code space. The soundness of a code, denoted by $R(\delta)$, measures the likelihood that a randomly selected local constraint is violated by a string that is at Hamming distance at least $\delta n$ from the code space, where $n$ is the codeword length. The concept of locally testability has deeply influenced fields like PCP, combinatorial optimization, property testing, program verification, and cryptography.
Aharonov and Eldar introduced quantum locally testable codes (QLTCs) aharonov2015quantum, with interest surging after Eldar and Harrow showed eldar2017local that QLTCs with constant soundness, locality, and relative distance could construct Hamiltonians lacking low-energy trivial states—addressing the NLTS conjecture freedman2014quantum. QLTCs are closely related to the quantum PCP conjecture aharonov2013guest, a fundamental problem in quantum complexity theory related to the quantum analog of the classical PCP theorem.
Roughly speaking, the soundness property ensures that high-weight errors produce high-weight syndromes. For stabilizer codes, soundness can be characterized as follows. Consider a code $C$ defined on $n$ qubits, determined by a generating set $\mathcal{G}$ where each generator has support over $w$ qubits (representing the stabilizer size). For any error $E$ acting on the code with a syndrome weight of at least $\delta n$, a randomly chosen generator $g\in\mathcal{G}$ fails to commute with $E$ with probability at least $R(\delta)$. This property ensures that errors can be detected probabilistically by examining only a limited number of stabilizers. Additionally, this soundness property induces a macroscopic energy barrier rakovszky2023physics.
To analyze single-shot property of quantum codes, Campbell provided the following alternative definition of soundness campbell2019theory.
Definition 2 (Soundness, Definition 3 in campbell2019theory).
Let $t$ be an integer and $f:\mathbb{Z}\rightarrow\mathbb{R}$ be a function termed the soundness function, where $f(0)=$ 0. A set of Pauli stabilizers $\mathcal{M}$ is called $(t,f)$-sound if, for every Pauli error $E$ with syndrome weight $|\sigma(E)|=x<t$, there exists an $E^{\star}$ such that $\sigma\left(E^{\star}\right)=\sigma(E)$ and the weight wt $\left(E^{\star}\right)\leqslant f(x)$.
In essence, this definition requires that low-weight syndromes are caused by low-weight errors. This formulation represents the contrapositive of definition of locally testable codes, with the key distinction being its focus on scenarios where $|\sigma(E)|=x<t$. Under this definition, good soundness is characterized by function $f(x)$, which is a monotonically increasing polynomial function of $x$ that remains independent of the size of the check set.
Campbell demonstrated that codes possessing good soundness naturally exhibit single-shot error correction capabilities, with the specific performance determined by the soundness parameters $t,f$ and the code parameters $[n,k,d]$ campbell2019theory. Moreover, good soundness in LDPC codes implies the existence of a macroscopic energy barrier aharonov2015quantum; campbell2019theory, as formalized in the following lemma:
Lemma 3 (Lemma 3 in campbell2019theory).
Consider a $\left[\left[n,k,d\right]\right]$ quantum code with checks $\mathcal{M}$ that is $(t,f)$-sound and where all qubits are involved in no more than $w_{c}$ checks. It follows that the energy barrier is at least $f^{-1}(c)$ where $c=\min\left[(t-1)/w_{c},\left(d-1\right)/2\right]$ and $f^{-1}$ is the inverse of the soundness function.
It is important to note that the converse statement, any LDPC check family with a macroscopic energy barrier necessarily possesses good soundness, does not hold. The expander code provides a clear counterexample: despite exhibiting bad soundness due to its lack of check redundancy ben2010locally, it has been proven to support single-shot error correction with the small-set flip decoder fawzi2020constant; fawzi2018efficient. In this context, check redundancy refers to the ratio between the check set size (encompassing all checks, including dependent ones) and the number of independent checks.
The soundness of a stabilizer code depends specifically on the selection of the stabilizer set, not merely on the stabilizer generators. Campbell also demonstrated that for any stabilizer code, it is possible to strategically choose a stabilizer set that endows the code with good soundness properties campbell2019theory. The price is that the size of some checks could be excessively large. LDPC property may be lost due to the existence of high-weight checks.
To maintain the LDPC characteristic, Campbell discovered an elegant solution for achieving single-shot properties: applying the hypergraph product construction twice campbell2019theory. The resulting structures naturally possess meta-checks, a direct consequence of the iterative hypergraph product process. Through this approach, Campbell effectively combined soundness and LDPC properties by strategically leveraging check redundancy. The following lemma shows this formally.
Lemma 4 (Lemma 5 and Lemma 6 in campbell2019theory).
Let $C_{0}\xleftarrow{\partial_{0}}C_{1}$ be a chain complex. Applying the hypergraph product, we obtain a new chain complex $\tilde{C}_{-1}\xleftarrow{\tilde{\delta}_{-1}}\tilde{C}_{0}\xleftarrow{\tilde{\delta}_{0}}\tilde{C}_{1}$, where the maps $\tilde{\delta}_{0}^{T}$ and $\tilde{\delta}_{-1}$ are $(t,f)$-sound with $f(x)=x^{2}/4$ and $t=\min\left[d_{0},d_{0}^{T}\right]$, where $d_{0},d_{0}^{T}$ are the distances of codes defined by $\delta_{0},\delta_{0}^{T}$, respectively. Applying the hypergraph products again, we obtain a new chain complex $\breve{C}_{-2}\xleftarrow{\breve{\delta}_{-2}}\breve{C}_{-1}\xleftarrow{\breve{\delta}_{-1}}\breve{C}_{0}\xleftarrow{\breve{\delta}_{0}}\breve{C}_{1}\xleftarrow{\breve{\delta}_{1}}\breve{C}_{2}$, where the maps $\breve{\delta}_{0}$ and $\breve{\delta}_{-1}^{T}$ are $(t,g)$-sound with soundness function $g(x)=x^{3}/4$, and $t=\min\left[d_{0},d_{0}^{T}\right]$.
For a 4D hypergraph product code that constructed with four identical classical codes defined by the same parity check matrix $\delta$ with distance $d$ ($\delta^{T}$ with distance $d^{T}$), according to Lemma 3, the energy barrier for logical operators of the resulting quantum code then has a lower bound of $\Omega\left({(\min[d,d^{T}])}^{\frac{1}{3}}\right)$. Later, we will demonstrate that the optimal lower bound for this energy barrier scales as $\Omega(\min[d,d^{T}])$ in the given scenario.
II.4 Confinement
Confinement is another robustness measure of quantum codes. Bombín originally introduced the concept of confinement as a mechanism that allows quantum codes to achieve single-shot error correction bombin2015single. However, in this discussion, we use a more developed definition provided in Quintavalle2021single.
Definition 5 (Confinement, Definition 1 in Quintavalle2021single).
Let $t$ be an integer and $f:\mathbb{Z}\rightarrow$ $\mathbb{Z}$ some increasing function with $f(0)=0$. We say that a stabilizer code has $(t,f)$-confinement if, for all errors $e$ with $\mathrm{wt}_{\mathrm{red}}(e)\leq t$, it holds
$f(|\sigma(e)|)\geq\mathrm{wt}_{\mathrm{red}}(e).$ (7)
In this definition, $\text{wt}_{\mathrm{red}}(e)$ denotes the reduced weight of the error $e$, which represents the smallest weight of any error $e^{\prime}$ that produces the same error syndrome as $e$.
Unlike Bombín’s more restrictive formulation (Definition 16 in bombin2015single), the above definition permits nonlinear functions $f(x)$. 2D repetition code and 3D hypergraph product codes both exhibit superlinear confinement functions while supporting single-shot correction. These important cases would be excluded under Bombín’s original framework.
A code family is said to have good confinement if each code within the family satisfies $(t,f)$-confinement with the following criteria: (1) $t$ increases with the length of the code $n$, specifically $t\in\Omega\left(n^{b}\right)$ for some $b>0$, and (2) the confinement function $f(\cdot)$ increases monotonically and does not depend on $n$. Additionally, a code family is said to exhibit good X-confinement if this property applies only to Pauli Z errors Quintavalle2021single.
Confinement is a weaker requirement than soundness. While good soundness requires low-weight syndromes to correspond to low-weight errors, good confinement demands that low-weight errors produce low-weight syndromes. For an LDPC code that is $(t,f)$-sound, if the maximum qubit degree is $\omega_{q}$ (the sparsity parameter), then the code has $({t}/{\omega_{q}},f)$ confinement (see Lemma 2 in  Quintavalle2021single). This follows from a direct analysis: if $e$ is an error set with $\text{wt}_{\mathrm{red}}(e)\leq{t}/{\omega_{q}}$, then its syndrome satisfies $|\sigma(e)|\leq{t}/{\omega_{q}}\cdot\omega_{q}=t$. By the soundness of the code, we have $f(|\sigma(e)|)\geq\text{wt}_{\mathrm{red}}(e)$.
Importantly, one can show that codes with good confinement possess single-shot error correction capabilities Quintavalle2021single.
Lemma 6 (Theorem 1 in Quintavalle2021single).
Consider a family of $[[n,k,d]]$ quantum-LDPC codes with good confinement such that $d\geq an^{b}$ with $a,b>0$. This code family is single-shot for the adversarial noise model. If the code family only has good X-confinement then it is single-shot with respect to Pauli Z noise.
Confinement is also related to the energy barrier of a code bombin2015single; hong2024quantum; placke2024topological. In Sec. III, we demonstrate formally that confinement naturally induces an energy barrier lower bound.
II.5 Expander code
An expander graph embodies a bipartite graph characterized by significant connectivity. Let $G=(V\cup C,E)$ be a bipartite graph. For a vertex subset $S\subseteq V$, a vertex $g\in C$ is a neighbor of $S$ if it is connected to at least one vertex in $S$, $N(S)$ is the set of neighbors of $S$. We say $g\in C$ is a unique neighbor of $S$ if it is only connected to a single vertex in $S$. We use $U(S)$ to denote the collection of unique neighbors of $S$. The expander graph can be defined as follows:
Definition 7 (Expander graph).
Let $G=(V\cup C,E)$ be a bipartite graph with left and right degrees bounded by $w$. Let $|V|=n$ and $|C|=m$. We say $G$ is $(\gamma_{1},\partial_{1})$-left-expanding with constants $\gamma_{1},\partial_{1}>0$, if for any subset $S\subseteq V$ with $|S|\leq\gamma_{1}n$, the neighbor set $N(S)$ of $S$ in $C$ satisfies
$|N(S)|\geqslant w(1-\partial_{1})|S|.$ (8)
Similarly, we say $G$ is $(\gamma_{2},\partial_{2})$-right-expanding with constants $\gamma_{2},\partial_{2}>0$, if for any subset $T\subseteq C$ with $|T|\leq\gamma_{2}n$, the neighbor $N(T)$ of $T$ in the subset $V$ satisfies $|N(T)|\geqslant w(1-\partial_{2})|T|$. We call $G$ an $\left(n,m,w,\gamma_{1},\partial_{1},\gamma_{2},\partial_{2}\right)$ expander graph if it is both left and right expanding.
Let $H_{G}$ be the parity check matrix of an expander code defined by this $\left(n,m,w,\gamma_{1},\partial_{1},\gamma_{2},\partial_{2}\right)$ expander graph (choosing $V$ and $C$ as the variable and check nodes of the Tanner graph) with $\partial_{1}<1/2$. Then for any $S\subset V$ such that $|S|\leq\gamma_{1}n,\quad U(S)\geq w(1-2\partial_{1})|S|$. Furthermore, the distance $d_{C}$ of the code $\mathcal{C}(G)$ is greater than $\gamma_{1}n$ Sipser1996.
The above analysis directly establishes the confinement property of expander codes. For any error $e$ with reduced weight $\text{wt}_{\mathrm{red}}(e)<d$, we have $|\sigma(e)|\geqslant U(e)\geqslant w(1-2\partial_{1})|e|$. Noting that in classical codes, $\text{wt}_{\mathrm{red}}(e)=|e|$. We can then conclude that $({1}/({w(1-2\partial_{1})}))|\sigma(e)|\geqslant\text{wt}_{\mathrm{red}}(e)$. This relationship demonstrates that expander codes exhibit linear confinement.
Conversely, linear confinement also induces left-expansion. For $(t,f)$-confinement with $t=an$ and $f(x)=cx$, where $a$ and $c$ are constants, any error $S$ with $|S|<an$ satisfies $c|\sigma(S)|\geqslant|S|$. Thus $N(S)\geqslant|\sigma(S)|\geqslant(1/c)|S|$. With degree $w$, we have $N(S)\geqslant w(1-(1-{1}/{(cw)}))|S|$. Therefore, this code is an expander code with (a,11/(cw))(a,1-{1}/{(cw)})-left-expanding.
Furthermore, the energy barrier of expander codes can be characterized by the following lemma.
Lemma 8.
Let $\Delta(H_{G})$ be the energy barrier of classical expander code defined by $H_{G}$ with parameters $(n,m,w,\gamma_{1},\partial_{1},\gamma_{2},\partial_{2})$ and $\partial_{1}<{1}/{2}$, then
$\Delta(H_{G})\geq cn,$ (9)
where $c>0$ is a constant.
Proof.
Without loss of generality, suppose that energy barrier $\Delta(H_{G})$ of code is given by the energy barrier of logical operator $L$, with path $r_{L}$. Since $|L|\geq\gamma_{1}n$, there is a point along the path (denoted as $P$) such that $\left|P\right|=\gamma_{1}n$. With the expansion property $U(P)\geq w(1-2\partial_{1})|P|$, we get $\Delta(H_{G})\geq U(P)\geq w(1-2\partial_{1})|P|=w(1-2\partial_{1})\gamma_{1}n$. Then, we have $\Delta(H_{G})\geq cn$ for $c=w(1-2\partial_{1})\gamma_{1}>0$. ∎
From Lemma 8, it follows that the classical expander code has an energy barrier scaling linearly with the number of bits. A similar conclusion holds for the transpose code in which the variable and the check nodes are exchanged. Therefore, according to the results in zhao2024on, the quantum expander code defined by parity check matrix $H_{(H_{G},H_{G})}$ has $\Omega(\sqrt{n})$ energy barrier, where $n$ here is the number of qubits.
Researches have shown that good soundness requires check redundancy in LDPC codes ben2010locally; ben2003stoc, the expander code exhibits poor soundness due to its lack of check redundancy. However, despite this limitation, the quantum expander code has been proven to support single-shot error correction when implemented with the small-set flip decoder fawzi2020constant; fawzi2018efficient. This makes confinement a more general and inclusive property compared to soundness, in relation to single-shot error correction.
II.6 Higher-dimensional hypergraph product
The hypergraph products provide a way to construct quantum codes with any two classical linear codes 2014-Tillich. It enables the application of extensive insights from classical coding theory to quantum coding. Those codes have good properties, such as circuit-level distance preservation manes2023distance; tan2025effective.
Given two classical codes $C_{1}$ and $C_{2}$, with respective parity check matrices $H_{1}$ and $H_{2}$, the hypergraph product defines a quantum code with parity check matrices
$H_{X}$ $=\left(H_{1}\otimes I\quad I\otimes H_{2}^{T}\right),$ (10)
$H_{Z}$ $=\left(I\otimes H_{2}\quad H_{1}^{T}\otimes I\right).$
One can verify that $H_{X}H_{Z}^{T}=0$ within the field $\mathbb{F}_{2}$. Thus these two parity check matrices define a CSS code, with a quantum parity check matrix of the form
$H_{\left(H_{1},H_{2}\right)}=\left(\begin{array}[]{cc}H_{X}&0\\ 0&H_{Z}\end{array}\right).$ (11)
Hypergraph product codes can be interpreted through chain complexes. Classical codes correspond to length-2 chain complexes, where boundary maps define parity check matrices. A hypergraph product code arises from the tensor product of two such chain complexes, yielding a length-3 chain complex. This structure provides two boundary maps that serve as parity checks. The fundamental property that “a boundary has no boundary” ensures these checks commute, allowing the construction of CSS quantum codes directly from the boundary maps of the resulting length-3 chain complex.
HHGP codes zeng2019higher extends this concept by generalizing the construction to create a length-$n$ chain complex. This generalization has important connections to topological codes: $n$-dimensional toric codes on hypercubic lattices constitute a length-$n$ chain complex bombin2007homological; dennis2002topological, and can be perfectly described by higher-dimensional hypergraph products.
Following Ref. zeng2019higher, we give a short introduction to the chain complex interpretation of hypergraph product. A chain complex is a sequence of finite-dimensional vector spaces $\{\ldots,\mathcal{A}_{j-1},\mathcal{A}_{j},\ldots\}$ connected by boundary operators $\partial_{j}:\mathcal{A}_{j-1}\leftarrow\mathcal{A}_{j}$, satisfying $\partial_{j}\partial_{j+1}=0$ for all $j\in\mathbb{Z}$.
For qubit systems, we consider vector spaces $\mathcal{A}_{j}=\mathbb{F}_{2}^{n_{j}}$, consisting of binary vectors of length $n_{j}$. $\mathcal{A}\equiv\mathcal{K}\left(A_{1},\ldots,A_{m}\right)$ can be defined as a length-$(m+1)$ chain complex with boundary map $\partial_{j}$ represented by $n_{j-1}\times n_{j}$ binary matrices $A_{j}$:
$\mathcal{A}:\{0\}\stackrel{{\scriptstyle\partial_{0}}}{{\leftarrow}}\mathcal{A}_{0}\stackrel{{\scriptstyle A_{1}}}{{\leftarrow}}\mathcal{A}_{1}\ldots\stackrel{{\scriptstyle A_{m}}}{{\leftarrow}}\mathcal{A}_{m}\stackrel{{\scriptstyle\partial_{m+1}}}{{\leftarrow}}\{0\},$ (12)
where adjacent matrices satisfy the orthogonality condition $A_{j-1}A_{j}=0$ for $j\in\{1,\ldots,m\}$. The trivial operators $\partial_{0}:\{0\}\leftarrow\mathcal{A}_{0}$ and $\partial_{m+1}:\mathcal{A}_{m}\leftarrow\{0\}$ are treated as zero matrices of dimensions $0\times n_{0}$ and $n_{m}\times 0$, respectively.
The subspace $\operatorname{Im}\left(A_{j+1}\right)\subseteq\mathcal{A}_{j}$ consists of boundaries, which can be represented as linear combinations of the columns of $A_{j+1}$. This subspace defines a binary linear code with generator matrix $A_{j+1}^{T}$, denoted as $\operatorname{Im}\left(A_{j+1}\right)=\mathcal{C}_{A_{j+1}^{T}}$. For the special case where $j=m$, we have $\operatorname{Im}\left(\partial_{m+1}\right)={0}$, which forms a trivial vector space.
The subspace $\operatorname{Ker}\left(A_{j}\right)\subseteq\mathcal{A}_{j}$ consists of cycles, which are vectors $x\in\mathcal{A}_{j}$ orthogonal to the rows of $A_{j}$, satisfying $A_{j}x^{T}=0$. This subspace defines a binary linear code with parity check matrix $A_{j}$, expressed as $\operatorname{Ker}\left(A_{j}\right)=\mathcal{C}_{A_{j}}^{\perp}$. In the special case where $j=0$, we have $\operatorname{Ker}\left(\partial_{0}\right)=\mathcal{A}_{0}$, encompassing the entire vector space.
Within the binary vector space $\mathcal{A}_{j}$, we have defined two key subspaces: $\operatorname{Im}\left(A_{j+1}\right)$, consisting of boundaries, and $\operatorname{Ker}\left(A_{j}\right)$, comprising cycles. A fundamental property of chain complexes is that all boundaries are cycles, though not all cycles are boundaries. This relationship is mathematically expressed through the orthogonality condition $A_{j}A_{j+1}=0$.
Based on these subspaces, we can define the $j$-th homology group as
$H_{j}(\mathcal{A})\equiv\operatorname{Ker}\left(A_{j}\right)/\operatorname{Im}\left(A_{j+1}\right).$ (13)
This quotient space contains all elements that are cycles but not boundaries, which we refer to as homologically nontrivial cycles.
The rank of the $j$-th homology group, representing the dimension of the associated vector space, is given by
$k_{j}\equiv\operatorname{rank}H_{j}(\mathcal{A})=n_{j}-\operatorname{rank}A_{j}-\operatorname{rank}A_{j+1}.$ (14)
Additionally, the homological distance $d_{j}$ is defined as the minimum Hamming weight of a nontrivial element in $H_{j}(\mathcal{A})$, expressed as:
$d_{j}=\min\{\operatorname{wt}(x):x\in H_{j}(\mathcal{A}),x\neq 0\}.$ (15)
A quantum CSS code can be derived from a length-3 chain complex $\mathcal{A}_{0}\stackrel{{\scriptstyle A_{1}}}{{\leftarrow}}\mathcal{A}_{1}\stackrel{{\scriptstyle A_{2}}}{{\leftarrow}}\mathcal{A}_{2}$ by setting $H_{Z}^{T}=A_{1}$ and $H_{X}=A_{2}$. In chain complex terminology, the code length equals $\operatorname{dim}\mathcal{A}_{1}$, while its dimension corresponds to the first homology group $\mathcal{H}_{1}=\operatorname{ker}A_{1}/\operatorname{Im}A_{2}$, or equivalently, the first cohomology group $\mathcal{H}_{1}^{*}=\operatorname{ker}A_{1}^{T}/\operatorname{Im}A_{2}^{T}$. The X and Z distances are determined by the minimum weights of non-zero vectors in $\mathcal{H}_{1}^{*}$ and $\mathcal{H}_{1}$ respectively.
Classical codes can be described by length-2 chain complexes. To construct quantum codes from classical codes, it is just a question of how to build a length-3 chain complex from two length-2 chain complexes. For this purpose, we employ the tensor product of chain complexes.
The tensor product $\mathcal{A}\times\mathcal{B}$ of two chain complexes $\mathcal{A}$ and $\mathcal{B}$ is defined as a new chain complex in the following way
$(\mathcal{A}\times\mathcal{B})_{l}=\bigoplus_{i+j=l}\mathcal{A}_{i}\otimes\mathcal{B}_{j}.$ (16)
The vector spaces in the new chain complex are direct sums of Kronecker products of vector spaces from the original chain complexes. Moreover, the boundary operators in this tensor product construction act as
$\partial_{i+j}(a\otimes b)=\partial_{i}^{\prime}a\otimes b+(-1)^{i}a\otimes\partial_{j}^{\prime\prime}b,$ (17)
where $a\in\mathcal{A}_{i},b\in\mathcal{B}_{j}$, and $\partial_{i}^{\prime}$ and $\partial_{j}^{\prime\prime}$ are the boundary operators in the corresponding space $\mathcal{A}$ and $\mathcal{B}$. If both $\mathcal{A}$ and $\mathcal{B}$ have finite dimension, then the dimension of $\mathcal{C}=\mathcal{A}\times\mathcal{B}$ is
$n_{j}(\mathcal{C})=\sum_{i}n_{i}(\mathcal{A})n_{j-i}(\mathcal{B}).$ (18)
By the Künneth theorem, the homology groups of the product complex $\mathcal{C}=\mathcal{A}\times\mathcal{B}$ are given by
$H_{j}(\mathcal{C})\cong\bigoplus_{i}H_{i}(\mathcal{A})\otimes H_{j-i}(\mathcal{B}).$ (19)
Consequently, the rank $k_{j}(\mathcal{C})$ of the $j$-th homology group $H_{j}(\mathcal{C})$ can be expressed as
$k_{j}(\mathcal{C})=\sum_{i}k_{i}(\mathcal{A})k_{j-i}(\mathcal{B}).$ (20)
The hypergraph product code is defined as the tensor product of two length-2 chain complexes. The higher-dimensional hypergraph product extends this concept to the tensor product of two chain complexes of arbitrary lengths. As demonstrated in Ref. zeng2019higher, for HHGP codes formed from a length-m chain complex $\mathcal{A}$ and a length-l chain complex $\mathcal{B}$ with $l=2$, there exists a lower bound on the distance of logical operators in the resulting quantum codes:
$d_{j}(\mathcal{A}\times\mathcal{B})=\min\left[d_{j-1}(\mathcal{A})d_{1}(\mathcal{B}),d_{j}(\mathcal{A})d_{0}(\mathcal{B})\right].$ (21)
In this work, we focus specifically on those 3D and 4D hypergraph product codes that can be regarded as tensor products of three or four classical codes, respectively. This allows us to apply the distance bounds given by above equation.
III Confinement and energy barrier
The relation between confinement and energy barrier is well-known bombin2015single; hong2024quantum; placke2024topological. In this section, we formalize this connection through the following lemma.
Lemma 9.
For a $[[n,k,d]]$ quantum code $\mathcal{C}$ with $(t,f)$ confinement where $t\leqslant d$ and $f(\cdot)$ is increasing monotonically, the energy barrier of the code is at least $f^{-1}(w)$, where $w=\min[t,{(d-1)}/{2}]$.
Proof.
Without loss of generality, suppose that the code energy barrier $\Delta(\mathcal{C})$ is given by logical operator $L$, with a path $r_{L}$. Since $|L|\geqslant d\geqslant t$, we first show that there always exist a step $P_{c}$ along the path $r_{L}$, such that $\text{wt}_{\mathrm{red}}(P_{c})\geqslant(d-1)/2$. This can be argued as follows campbell2019theory:
Suppose $r_{L}=\{P_{0},P_{1},\cdots,P_{L}\}$, and each $E_{j}=P_{j}P_{j-1}$ is single qubit Pauli operator. For each $P_{j}$, we define the reduced weight as:
$\mathrm{wt}_{\mathrm{red}}(P_{j}):=\min_{V}\{\mathrm{wt}(P_{j}V):V\in\mathcal{P},\sigma(V)=0\},$ (22)
where the minimization is over Pauli operators $V$ with trivial syndrome. Let $V_{j}$ denote the Pauli operators achieving this minimum for $P_{j}$. Since $\sigma\left(V_{j}\right)=0$, $V_{j}$ is either a stabilizer or a nontrivial logical operator.
Trivially, we find $V_{0}=I$ and $V_{L}=P_{L}$, indicating that the sequence starts with a stabilizer and ends with a logical operator. Therefore, there exists an index $j^{*}$ such that $V_{j^{*}}$ is a stabilizer, while $V_{j^{*}+1}$ is a nontrivial logical operator. Then $V_{j^{*}}V_{j^{*}+1}$ is also a nontrivial logical operator, leading to:
$d\leq\mathrm{wt}\left(V_{j^{*}}V_{j^{*}+1}\right).$ (23)
Notice that
$\operatorname{wt}\left(V_{j^{*}}V_{j^{*}+1}\right)$ $=$ $\operatorname{wt}\left(V_{j^{*}}V_{j^{*}+1}P_{j^{*}}P_{j^{*}}^{\dagger}P_{j^{*}+1}P_{j^{*}+1}^{\dagger}\right)$
$=$ $\operatorname{wt}\left(V_{j^{*}}P_{j^{*}}V_{j^{*}+1}P_{j^{*}+1}P_{j^{*}}^{\dagger}P_{j^{*}+1}^{\dagger}\right).$
Applying the triangle inequality twice gives:
$\mathrm{wt}\left(V_{j^{*}}V_{j^{*}+1}\right)\leq\mathrm{wt}_{\mathrm{red}}\left(P_{j^{*}}\right)+\mathrm{wt}_{\mathrm{red}}\left(P_{j^{*}+1}\right)+1.$ (25)
Thus, combining this result, we obtain:
$d\leq 2\cdot\max\left[\mathrm{wt}_{\mathrm{red}}\left(P_{j^{*}}\right),\mathrm{wt}_{\mathrm{red}}\left(P_{j^{*}+1}\right)\right]+1,$ (26)
which implies
$\frac{d-1}{2}\leq\max\left[\mathrm{wt}_{\mathrm{red}}\left(P_{j^{*}}\right),\mathrm{wt}_{\mathrm{red}}\left(P_{j^{*}+1}\right)\right].$ (27)
Thus, the sequence of reduced weights $\left\{\mathrm{wt}_{\mathrm{red}}\left(P_{0}\right),\mathrm{wt}_{\text{red}}\left(P_{1}\right),\cdots,\mathrm{wt}_{\text{red }}\left(P_{L}\right)\right\}$ starts and ends with zero, and it reach a value of at least $\left(d-1\right)/2$ in the middle. Moreover, the local error condition ensures that $\left|\mathrm{wt}_{\mathrm{red}}\left(P_{j+1}\right)-\mathrm{wt}_{\text{red }}\left(P_{j}\right)\right|$ is either 0 or 1, then the sequence must include every integer from 0 to $\left(d-1\right)/2$.
Recall that $(t,f)$ confinement means that for all errors $e$ with $\text{wt}_{\mathrm{red}}(e)\leq t$, $f(|\sigma(e)|)\geq\text{wt}_{\mathrm{red}}(e)$. Then for $w=\min[t,(d-1)/2]$, there must exist an $P_{w}$ with $\text{wt}_{\text{red}}(P_{w})=w$, such that $f(|\sigma(P_{w})|)\geqslant\text{wt}_{\mathrm{red}}(P_{w})=w$. Given $f(\cdot)$ is monotonically increasing, $|\sigma(P_{c})|\geqslant f^{-1}(w)$. This implies that the energy barrier of code $\mathcal{C}$ is at least $f^{-1}(w)$. ∎
The converse statement of Lemma 9—that any quantum code with a macroscopic energy barrier necessarily exhibits confinement—does not generally hold. This can be demonstrated with a straightforward counterexample. Consider a quantum LDPC code $\mathcal{C}$ constructed as a composite system comprising two regions: Region A, based on a 2D quantum repetition code $[[n,1,n]]$ with energy barrier $O(\sqrt{n})$, and Region B, based on a 1D quantum repetition code $[[n,1,n]]$ with constant energy barrier $O(1)$, see Fig. 1. The overall code is a $[[2n,1,2n]]$ quantum repetition code with energy barrier $O(\sqrt{n})$ determined by Region A. However, for an error $e$ contained entirely within Region B, we can have $\text{wt}_{\mathrm{red}}(e)=|e|$ that grows up to (d1)/2(d-1)/2 with $|\sigma(e)|\leq c$, where $c$ is constant. Consequently, the overall code fails to satisfy the confinement criterion.
Quintavalle et al. proved that all 3D hypergraph product codes exhibit $(t,f)$ X-confinement (according to their convention) where $t$ equals the minimum distance of classical codes and $f(x)=x^{3}/4$ (see Lemma 10 of Ref Quintavalle2021single). For a 3D hypergraph product code constructed with three identical classical codes defined by the same parity check matrix $\delta$ with distance $d$, the result in Ref Quintavalle2021single implies a lower bound for the energy barrier of Z-type logical operators as $\Omega\left(d^{\frac{1}{3}}\right)$. In section V, we will demonstrate that the optimal lower bound for this energy barrier actually scales as $\Omega(d)$ in this scenario.
Region ARegion B
Figure 1: A composite system consisting of a 2D quantum repetition code (Region A) and a 1D quantum repetition code arranged in a snake pattern (Region B). The overall code exhibits an energy barrier determined by the 2D quantum repetition code, which scales as the linear dimension of Region A. However, confinement properties are influenced by the 1D quantum repetition component, resulting in the absence of confinement for the composite system.
IV Tensor product code and its energy barrier
In this section, we analyze the energy barrier of tensor product codes — a specific classical code structure that builds new classical codes from two existing ones. We will demonstrate that the energy barrier of the resulting code is related to the distance of the underlying classical codes.
This result is essential for our subsequent analysis of HHGP codes. Specifically, we establish that the logical operators of HHGP codes admit a tensor product decomposition. This structural property enables us to derive a lower bound on the energy barrier of HHGP codes in terms of the distances of their constituent classical codes.
IV.1 Code construction
Given two classical codes $\mathcal{C}_{a}$ and $\mathcal{C}_{b}$ with parity check matrix $\delta_{a}$ and $\delta_{b}$, the tensor product code $\mathcal{C}_{c}$ is a classical code defined to have the parity check matrix
$\delta_{c}=\left(\begin{array}[]{c}\delta_{a}\otimes I_{n_{b}}\\ I_{n_{a}}\otimes\delta_{b}\end{array}\right).$ (30)
This tensor product structure can be also understood from the perspective of code space. Let $\mathcal{C}_{a}$ with parameters $\left[n_{a},k_{a},d_{a}\right]$ and $\mathcal{C}_{b}$ with parameters $\left[n_{b},k_{b},d_{b}\right]$ be two classical codes. The tensor product of these two codes, denoted as $\mathcal{C}_{c}=\mathcal{C}_{a}\otimes\mathcal{C}_{b}$, encodes messages as follows: The message of the tensor product code $\mathcal{C}_{a}\otimes\mathcal{C}_{b}$ can be represented as a $k_{a}\times k_{b}$ matrix $M$. The encoding proceeds in two stages: first, each column of $M$ is encoded using $\mathcal{C}_{a}$, yielding a $n_{a}\times k_{b}$ intermediate matrix; second, each row of this intermediate matrix is encoded using $\mathcal{C}_{b}$, producing the final $n_{a}\times n_{b}$ codeword.
Let $L_{a}$ be the logical operator of $\delta_{a}$, and $L_{b}$ be the logical operator of $\delta_{b}$, one can verify that $L_{c}=L_{a}\otimes L_{b}$ is the logical operator of $\delta_{c}$ because
$\delta_{c}L_{c}$ $=$ $\left(\begin{array}[]{c}\delta_{a}\otimes I_{n_{b}}\\ I_{n_{a}}\otimes\delta_{b}\end{array}\right)(L_{a}\otimes L_{b})$ (33)
$=$ $\left(\begin{array}[]{c}(\delta_{a}\otimes I_{n_{b}})(L_{a}\otimes L_{b})\\ (I_{n_{a}}\otimes\delta_{b})(L_{a}\otimes L_{b})\end{array}\right)$ (36)
$=$ $0.$ (37)
Here we used the relation $\delta_{a}L_{a}=\delta_{b}L_{b}=0$.
Given $k_{a}$ distinct logical operators $L_{a}$ and $k_{b}$ distinct logical operators $L_{b}$, one can construct $k_{a}k_{b}$ distinct logical operators $L_{c}$ through this tensor product construction.
Moreover, the distance of the tensor product code is at least $d_{a}d_{b}$. A simple argument demonstrates this: consider two different message matrices of code $\mathcal{C}_{c}$, denoted as $M_{a}$ and $M_{b}$. After the first encoding step, we obtain intermediate matrices $N_{a}$ and $N_{b}$, and the second encoding step produces the final codewords $C_{a}$ and $C_{b}$. If $M_{a}$ and $M_{b}$ differ in the $i$-th row, then $N_{a}$ and $N_{b}$ differ in at least $d_{a}$ positions in that row, corresponding to $d_{a}$ columns. Let the indices of these differing columns be $\{j_{1},\ldots,j_{d_{a}}\}$. During the second encoding step, each of these columns generates differences in at least $d_{b}$ positions. Consequently, $C_{a}$ and $C_{b}$ differ in at least $d_{a}d_{b}$ positions. In summary, the tensor product code $\delta_{c}$ has parameters $n_{a}n_{b},k_{a}k_{b},d_{a}d_{b}]$.
IV.2 Energy barrier of tensor product code
Let $L_{a}$ and $L_{b}$ be the logical operators of $\delta_{a}$ and $\delta_{b}$. Our goal is to determine the energy barrier of the operator $L_{c}=L_{a}\otimes L_{b}$, which serves as a logical operator for $\delta_{c}$.
To compute the energy barrier of $L_{c}$, we consider a path $r=\{P_{0},P_{1},\cdots,P_{F}\}$ with $P_{0}=I$ and $P_{F}=L_{c}$. Then
$\Delta(L_{c})=\min\{\epsilon_{\max}(r):r\in w(0,L_{c})\}.$ (38)
Note that $r$ is a path on the bit of code $\delta_{c}$. Given a $P_{\ell}\in r$, the energy (the number of violated checks) of $P_{\ell}$ is
$\epsilon(P_{\ell})$ $=$ $\text{wt}(\delta_{c}P_{\ell})$ (41)
$=$ $\text{wt}\left(\left(\begin{array}[]{c}\delta_{a}\otimes I_{n_{b}}\\ I_{n_{a}}\otimes\delta_{b}\end{array}\right)P_{\ell}\right)$
$=$ $\text{wt}((\delta_{a}\otimes I_{n_{b}})P_{\ell})+\text{wt}((I_{n_{a}}\otimes\delta_{b})P_{\ell}).$ (42)
Because $\text{wt}(\cdot)\geqslant 0$, then we have
$\epsilon(P_{\ell})\geqslant\text{wt}((\delta_{a}\otimes I_{n_{b}})P_{\ell}),$ (43)
or
$\epsilon(P_{\ell})\geqslant\text{wt}((I_{n_{a}}\otimes\delta_{b})P_{\ell}).$ (44)
Because the path $r$ contains paths for $L_{a}$ and $L_{b}$, the above inequalities indicate that the energy barrier of $L_{c}$ is lower bounded by the energy barrier of $L_{a}$ or $L_{b}$. Furthermore, the energy barrier of code $\delta_{c}$ is then lower bounded by energy barrier of codes $\delta_{a}$ and $\delta_{b}$. We conclude this as the following Lemma
Lemma 10.
Given two classical codes $\delta_{a}$ and $\delta_{b}$ with parameters $n_{a},k_{a},d_{a},E_{a}]$ and $n_{b},k_{b},d_{b},E_{b}]$, where $E_{a}$ and $E_{b}$ are energy barriers of codes. Then the energy barrier of the tensor product code $\delta_{c}$ is lower bounded by
$\Delta(\delta_{c})\geqslant\min[E_{a},E_{b}].$ (45)
However, this may not be the tightest bound, particularly when codes $\delta_{a}$ and $\delta_{b}$ have constant energy barriers. In the following, we establish a relationship between the energy barrier of $L_{c}$ and the distance of $L_{a}$ and $L_{b}$.
IV.2.1 Lower bound
Given two classical codes $\delta_{a}$ and $\delta_{b}$ with energy barriers $\Delta(\delta_{a})=E_{a}$ and $\Delta(\delta_{b})=E_{b}$. Here, we show that the energy barrier of the tensor product code $\delta_{c}$ can be lower bounded by the distances of the two constituent codes. Formally, we state the following lemma.
Lemma 11.
Given two classical codes $\delta_{a}$ and $\delta_{b}$ with parameters $n_{a},k_{a},d_{a},E_{a}]$ and $n_{b},k_{b},d_{b},E_{b}]$. The energy barrier of the tensor product code $\delta_{c}$ is lower bounded by
$\Delta(\delta_{c})\geqslant\min[d_{a},d_{b}].$ (46)
Proof.
Let us first establish a clear representation of the tensor product code structure. Given the parity check matrices $\delta_{a}\in\{0,1\}^{m_{a}\times n_{a}}$ and $\delta_{b}\in\{0,1\}^{m_{b}\times n_{b}}$, we can arrange the code words of $\delta_{c}$ on a two-dimensional grid of size $n_{a}\times n_{b}$. Then, each row of the 2D grid corresponds to a copy of code $\delta_{b}$, and each column corresponds to a copy of code $\delta_{a}$.
The logical operator $L_{c}=L_{a}\otimes L_{b}$ can be visualized as a pattern on this grid, where $L_{a}$ specifies which rows are involved and $L_{b}$ specifies which columns are involved. Specifically, the support of $L_{c}$, denoted as $\mathrm{supp}(L_{c})$, consists of positions $(i,j)$ where $i\in\mathrm{supp}(L_{a})$ and $j\in\mathrm{supp}(L_{b})$, forming a subgrid of size $|L_{a}|\cdot|L_{b}|=d_{a}d_{b}$.
Any valid path $r=\{P_{r}^{0},P_{r}^{1},P_{r}^{2},\ldots,P_{r}^{L}\}$ from the identity operator $I$ to the logical operator $L_{c}$, by definition, must flip all the bits in the subset $\mathrm{supp}(L_{c})$. For any intermediate state $P_{r}^{i}$ along this path, the syndrome weight $\text{wt}(\delta_{c}P_{r}^{i})$ measures the energy cost at that step. To finish the proof, We show that for any valid path, there must exist a intermediate state with syndrome weight at least $\min[d_{a},d_{b}]$.
To flip all bits in $\mathrm{supp}(L_{c})$, the path must eventually form a logical operator in each of the $d_{a}$ rows indexed by $\mathrm{supp}(L_{a})$, as well as in each of the $d_{b}$ columns indexed by $\mathrm{supp}(L_{b})$. More precisely, in each row $i\in\mathrm{supp}(L_{a})$, the final configuration must contain a codeword of $\delta_{b}$. Similarly, each column $j\in\mathrm{supp}(L_{b})$ must contain a codeword of $\delta_{a}$.
Since the path proceeds by single bit flips, there must exist a critical intermediate state $P^{i^{*}}$, which represents the first moment when a logical operator of either $\delta_{a}$ or $\delta_{b}$ appears. More precisely, $P^{i^{*}}$ could be one of the following configurations:
  1. 1.
    Exactly one row $i\in\text{supp}(L_{a})$ contains a non-trivial logical operator $L^{\prime}_{b}$ of code $\delta_{b}$ (not necessarily to be $L_{b}$), while other rows and columns contain none, or
  2. 2.
    Exactly one column $j\in\text{supp}(L_{b})$ contains a non-trivial logical operator $L^{\prime}_{a}$ of code $\delta_{a}$ (not necessarily to be $L_{a}$), while other rows and columns contain none, or
  3. 3.
    In rare cases, exactly one row $i\in\text{supp}(L_{a})$ contains a non-trivial logical operator $L^{\prime}_{b}$ of code $\delta_{b}$ (not necessarily to be $L_{b}$), while simultaneously, exactly one column $j\in\text{supp}(L_{b})$ contains a non-trivial logical operator $L^{\prime}_{a}$ of code $\delta_{a}$ (not necessarily to be $L_{a}$), with all other rows and columns containing none.
Case 1: The row containing the logical operator has $d_{b}$ flipped bits at positions $(i,j_{1}),\ldots,(i,j_{d_{b}^{\prime}})$ where $\{j_{1},\ldots,j_{d_{b}^{\prime}}\}\subseteq\mathrm{supp}(L_{b}^{\prime})$. Each of these bits lies in a distinct column. Since these columns do not contain complete codewords of $\delta_{a}$, each such incomplete column contributes at least 1 to the syndrome weight. With $d_{b}^{\prime}\geqslant d_{b}$, the total syndrome weight is at least $d_{b}$.
Case 2: Similarly, the column containing the logical operator $L_{a}^{\prime}$ has $d_{a}^{\prime}$ flipped bits, each in a distinct row that does not form a complete codeword of $\delta_{b}$. Thus each incomplete row contributes at least 1 to the syndrome weight, yielding a total contribution of at least $d_{a}$.
Case 3: The row containing a logical operator of $\delta_{b}$ contributes at least $d_{b}$ flipped bits, and the column containing a logical operator of $\delta_{a}$ contributes at least $d_{a}$ flipped bits. With out loss of generality, suppose these two logical operators intersect at position $(i,j)$. The row induces syndrome violations in $d_{b}-1$ columns (excluding the column that contains a logical operator), and the column induces violations in $d_{a}-1$ rows (excluding the row that contains a logical operator). Thus, the syndrome weight is at least $(d_{a}-1)+(d_{b}-1)=d_{a}+d_{b}-2\geq\min[d_{a},d_{b}]$, where the final inequality holds when $d_{a},d_{b}\geq 2$.
In all cases, provided $d_{a},d_{b}\geq 2$ (which is standard for codes with non-trivial distance), the syndrome weight at the critical state $P^{i^{*}}$ is at least $\min[d_{a},d_{b}]$. Since the energy barrier $\Delta(L_{c})$ is defined as the minimum over all paths of the maximum syndrome weight along each path, we conclude $\Delta(L_{c})\geq\min[d_{a},d_{b}].$
The above result holds for any logical operator $L_{c}$ of $\delta_{c}$. Let us denote the set of logical operators of $\delta_{c}$ as $L(\delta_{c})$, then we have
$\Delta(\delta)=\min\{\Delta(L_{c}):L_{c}\in L(\delta_{c})\}\geq\min[d_{a},d_{b}].$ (47)
The intuition of Lemma 11 is from the 2D repetition code, which can be interpreted as the tensor product of two 1D repetition code. The 1D repetition code has parameters $L,1,L,\Omega(1)]$, where the energy barrier scales as $O(1)$. In contrast, the 2D repetition code exhibits parameters $\left[L^{2},1,L^{2},\Omega(L)\right]$, with an energy barrier scaling as $\Omega(L)$.
The energy barrier of the 2D repetition code can be rigorously analyzed through its tensor product structure. Consider a 2D repetition code on an $L\times L$ square lattice, where the logical operator corresponds to flipping all bits on the lattice. For any valid path (meets local error condition) implementing this logical operator, there necessarily exists an intermediate configuration where precisely a single row or a single column of bits has been flipped (or, a single row and a single column simultaneously). At this configuration, the number of violated checks (i.e., the energy cost) is at least $L$, as verified by analyzing the violated checks on each column and row. Consequently, the energy barrier of the 2D repetition code is lower-bounded by $L$.
IV.2.2 Upper bound - “strip” argument
To establish an upper bound, one can just consider any specific path from the identity to a logical operator. We analyze two distinct implementation strategies: row-wise and column-wise implementation of logical operators. Same to the “strip” argument in Ref. rakovszky2024physicsgoodldpccodes.
We still view the tensor product code $\delta_{c}$ as operating on a 2D grid of size $n_{a}\times n_{b}$. In this representation, each row corresponds to a copy of code $\delta_{b}$, each column corresponds to a copy of code $\delta_{a}$.
For the row-wise implementation, let $r_{a}=\{P_{a}^{0},P_{a}^{1},\ldots,P_{a}^{F}\}$ represent an optimal path for implementing logical operator $L_{a}$, where $P_{a}^{0}=I$ and $P_{a}^{F}=L_{a}$. By definition, the maximum syndrome weight along the path $r_{a}$ is $E_{a}$. Because $P_{i}P_{i+1}=E_{i}$ is a single Pauli for all $i$, we denote $q_{a}=\{j_{1},j_{2},\ldots,j_{L_{a}}\}$ as the position of single Pauli in each step of $r_{a}$. These positions correspond to rows on the 2D grid of the tensor product code.
We consider a path for implementing $L_{c}$ by applying $L_{b}$ sequentially according to the positions in $q_{a}$. Specifically, Our path $r_{c}$ for implementing $L_{c}$ begins with $P_{c}^{0}=I$, for each position $j_{k}\in q_{a}$, we implement the logical operator $L_{b}$ on the corresponding row. After finishing all rows in $q_{a}$, we have $P_{c}^{F}=L_{c}$.
Now consider the energy barrier given by this specific $r_{c}$. Given any $P_{c}^{i}\in r_{c}$, the energy of $P_{c}^{i}$ is
$\epsilon(P_{c}^{i})=\text{wt}((\delta_{a}\otimes I_{n_{b}})P_{c}^{i})+\text{wt}((I_{n_{a}}\otimes\delta_{b})P_{c}^{i}).$ (48)
The maximum syndrome weight along this path is less than $d_{b}E_{a}+E_{b}$. This follows because for the part $\text{wt}((\delta_{a}\otimes I_{n_{b}})P_{c}^{i})$, each column contributes a syndrome weight of at most $E_{a}$, and there are $d_{b}$ columns where $L_{b}$ has support, which gives $d_{b}E_{a}$. While for the part $\text{wt}((I_{n_{a}}\otimes\delta_{b})P_{c}^{i})$, the maximum syndrome weight is $E_{b}$ because we implement $L_{b}$ sequentially.
Analogously, for column-wise implementation, let $r_{b}=\{P_{b}^{0},P_{b}^{1},\ldots,P_{b}^{F}\}$ represent an optimal path for implementing $L_{b}$, with maximum syndrome weight $E_{b}$. One can construct another path for $L_{c}$ by implementing $L_{a}$ on columns according to the positions provided by path $r_{b}$. The maximum syndrome weight along this path is less than $d_{a}E_{b}+E_{a}$.
By selecting the implementation strategy with the lower maximum syndrome weight, we establish the upper bound:
$\Delta(L_{c})\leqslant\min[d_{b}E_{a}+E_{b},d_{a}E_{b}+E_{a}].$ (49)
In addition, we assert that (although we cannot provide a formal proof) there is no path with an energy barrier lower than $\min[d_{b}E_{a},d_{a}E_{b}]$, which we formally state in the following conjecture.
Conjecture 12.
Given two classical codes $\delta_{a}$ and $\delta_{b}$ with parameters $n_{a},k_{a},d_{a},E_{a}]$ and $n_{b},k_{b},d_{b},E_{b}]$, the energy barrier of the tensor product code $\delta_{c}$ is
$\Delta(\delta_{c})\geqslant\min[d_{a}E_{b},E_{a}d_{b}].$ (50)
We conjecture this result and anticipate that an elegant, rigorous proof will be found in the future (See more discussion in the Appendix). Meanwhile, the lower bound provided in Lemma 11 is optimal with respect to the order of the distance, although the value $E_{a}$ and $E_{b}$ may yield a non-negligible difference in the absolute value of the energy barrier. Thus the lower bound in Lemma 11 is tight if underlying codes exhibit system size-dependent distances but constant energy barriers.
V Energy barrier of 3D hypergraph product codes
With all the necessary preparations in place, we now turn our attention to the energy barrier of HHGP codes. In this section, we consider the simplest instance, 3D hypergraph product codes that are constructed from three classical codes.
We begin by reviewing the structure of 3D hypergraph product codes and analyzing their logical operators, defining a complete set as elementary canonical logical operators. We first address Z logical operators. Using the strip argument from Ref. rakovszky2024physicsgoodldpccodes, we establish an upper bound for the energy barrier. We then derive a lower bound by examining the structure of those elementary canonical Z logical operators. Based on energy barrier of tensor product codes (Lemma 11), we demonstrate that this lower bound relates to the distances of the underlying classical codes. Finally, we discuss the energy barrier of the logical X operators.
V.1 Code construction of 3D hypergraph product codes
Consider three classical codes $\delta_{a}$, $\delta_{b}$, and $\delta_{c}$ with respective parameters $n_{\ell},k_{\ell},d_{\ell}]$, where $\ell\in\{a,b,c\}$, and the parameters of their transposes $\delta_{\ell}^{T}$ are denoted by $n_{\ell}^{T},k_{\ell}^{T},d_{\ell}^{T}]$. For convenience, we denote $r_{\ell}\equiv n_{\ell}^{T}$. Applying the 3D hypergraph product to these classical codes generates a length-4 chain complex Quintavalle2021single
$C_{0}\xleftarrow{\partial_{0}}C_{1}\xleftarrow{\partial_{1}}C_{2}\xleftarrow{\partial_{2}}C_{3},$ (51)
where $\partial_{0},\partial_{1}$ and $\partial_{2}$ are
$\partial_{0}=\left(\begin{array}[]{l}\delta_{a}\otimes I_{n_{b}}\otimes I_{n_{c}}\\ I_{n_{a}}\otimes\delta_{b}\otimes I_{n_{c}}\\ I_{n_{a}}\otimes I_{n_{b}}\otimes\delta_{c}\end{array}\right),$ (52)
$\partial_{1}=\left(\begin{array}[]{ccc}I_{r_{a}}\otimes\delta_{b}\otimes I_{n_{c}}&\delta_{a}\otimes I_{r_{b}}\otimes I_{n_{c}}&0_{r_{a}r_{b}n_{c}\times n_{a}n_{b}r_{c}}\\ I_{r_{a}}\otimes I_{n_{b}}\otimes\delta_{c}&0_{r_{a}n_{b}r_{c}\times n_{a}r_{b}n_{c}}&\delta_{a}\otimes I_{n_{b}}\otimes I_{r_{c}}\\ 0_{n_{a}r_{b}r_{c}\times r_{a}n_{b}n_{c}}&I_{n_{a}}\otimes I_{r_{b}}\otimes\delta_{c}&I_{n_{a}}\otimes\delta_{b}\otimes I_{r_{c}}\end{array}\right),$ (53)
$\partial_{2}=\left(I_{r_{a}}\otimes I_{r_{b}}\otimes\delta_{c}\quad I_{r_{a}}\otimes\delta_{b}\otimes I_{r_{c}}\quad\delta_{a}\otimes I_{r_{b}}\otimes I_{r_{c}}\right).$ (54)
Here $I_{n}$ denotes an identity matrix of size $n\times n$.
The relation $\partial_{1}\partial_{0}=0$ allows us to construct a quantum CSS code $\mathcal{C}\left(\delta_{a},\delta_{b},\delta_{c}\right)$ with parity check matrix $H_{Z}=\partial_{0}^{T}$ and $H_{X}=\partial_{1}$. The commutation relation follows as $H_{X}H_{Z}^{T}=\partial_{1}\partial_{0}=0$.
Additionally, since $\partial_{2}\partial_{1}=0$, the matrix $M=\partial_{2}$ serves as the meta-check for $H_{X}$. This means any valid X-syndrome must satisfy the constraints imposed by $M$.
The parameters of the code $\mathcal{C}\left(\delta_{a},\delta_{b},\delta_{c}\right)$ can be determined from the classical code $\delta_{a}$, $\delta_{b}$, and $\delta_{c}$. As demonstrated in Ref. zeng2019higher, the resulting quantum code $\mathcal{C}\left(\delta_{a},\delta_{b},\delta_{c}\right)$ is an $\left[\left[n,k,d_{x},d_{z}\right]\right]$ code with
$\begin{aligned} n&=n_{a}^{T}n_{b}n_{c}+n_{a}n_{b}^{T}n_{c}+n_{a}n_{b}n_{c}^{T},\\ k&=k_{a}^{T}k_{b}k_{c}+k_{a}k_{b}^{T}k_{c}+k_{a}k_{b}k_{c}^{T},\\ d_{x}&=\min\left[d_{a}^{T},d_{b}^{T},d_{c}^{T}\right],\\ d_{z}&=\min\left[d_{b}d_{c},d_{a}d_{c},d_{a}d_{b}\right].\end{aligned}$ (55)
V.2 Logical operators of 3D hypergraph product codes
[…]
Denote the binary form of logical operators of classical codes $\delta_{a}$, $\delta_{b}$, and $\delta_{c}$ as $L_{A}=\{L_{A}^{1},L_{A}^{2},\ldots,L_{A}^{k_{a}}\}$, $L_{B}=\{L_{B}^{1},L_{B}^{2},\ldots,L_{B}^{k_{b}}\}$, and $L_{C}=\{L_{C}^{1},L_{C}^{2},\ldots,L_{C}^{k_{c}}\}$, respectively. We first show that the following three sets of binary vectors give a complete set of Z logical operators.
$L_{Z}^{1}=\sum_{i,j,k}\alpha_{ijk}\left(\begin{array}[]{c}v_{i}\otimes L_{B}^{j}\otimes L_{C}^{k}\\ 0_{n_{a}r_{b}n_{c}}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right),$ (56)
$L_{Z}^{2}=\sum_{i,j,k}\beta_{ijk}\left(\begin{array}[]{c}0_{r_{a}n_{b}n_{c}}\\ L_{A}^{i}\otimes w_{j}\otimes L_{C}^{k}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right),$
$L_{Z}^{3}=\sum_{i,j,k}\gamma_{ijk}\left(\begin{array}[]{c}0_{r_{a}n_{b}n_{c}}\\ 0_{n_{a}r_{b}n_{c}}\\ L_{A}^{i}\otimes L_{B}^{j}\otimes u_{k}\end{array}\right),$
where $v_{i},w_{j}$, and $u_{k}$ are unit vectors with length $r_{a},r_{b}$, and $r_{c}$, respectively. $0_{n}$ is a zero vector with length $n$, and $\alpha_{ijk},\beta_{ijk},\gamma_{ijk}\in\{0,1\}$. Given that $\delta_{a}L_{A}^{i}=\delta_{b}L_{B}^{j}=\delta_{c}L_{C}^{k}=0$, one can confirm that $H_{X}L_{Z}^{l}=0$ for $l\in\{1,2,3\}$.
For $L_{Z}^{1}$, the condition $H_{X}L_{Z}^{1}=0$ ensures that any X-type stabilizer from $H_{X}$ commutes with $L_{Z}^{1}$. Similarly, any Z-type stabilizer from $H_{Z}$ naturally commutes with $L_{Z}^{1}$ as they both consist solely of Pauli Z operators. To prove that $L_{Z}^{1}$ are Z logical operators, we must prove that $L_{Z}^{1}$ are not in the image of $H_{Z}^{T}$. To proceed with this analysis, we first denote:
$L_{Z}^{1,i,j,k}=\left(\begin{array}[]{c}v_{i}\otimes L_{B}^{j}\otimes L_{C}^{k}\\ 0_{n_{a}r_{b}n_{c}}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right).$ (60)
For $a_{i}\in\{0,1\}$, suppose that the linear combination of these operators
$\sum_{i}a_{i}L_{Z}^{1,i,j,k}=\left(\begin{array}[]{c}(\sum_{i}a_{i}v_{i})\otimes L_{B}^{j}\otimes L_{C}^{k}\\ 0_{n_{a}r_{b}n_{c}}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right)$ (64)
is in the image of matrix $H_{Z}^{T}$. Then, there exists a vector $x$ such that
$\sum_{i}a_{i}L_{Z}^{1,i,j,k}$ $=$ $\left(\begin{array}[]{c}(\sum_{i}a_{i}v_{i})\otimes L_{B}^{j}\otimes L_{C}^{k}\\ 0_{n_{a}r_{b}n_{c}}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right)$
$=$ $H_{Z}^{T}(x\otimes L_{B}^{j}\otimes L_{C}^{k})$
$=$ $\left(\begin{array}[]{c}(\delta_{a}\otimes I_{n_{b}}\otimes I_{n_{c}})(x\otimes L_{B}^{j}\otimes L_{C}^{k})\\ (I_{n_{a}}\otimes\delta_{b}\otimes I_{n_{c}})(x\otimes L_{B}^{j}\otimes L_{C}^{k})\\ (I_{n_{a}}\otimes I_{n_{b}}\otimes\delta_{c})(x\otimes L_{B}^{j}\otimes L_{C}^{k})\end{array}\right)$ (72)
$=$ $\left(\begin{array}[]{c}\delta_{a}x\otimes L_{B}^{j}\otimes L_{C}^{k}\\ 0_{n_{a}r_{b}n_{c}}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right).$ (76)
Recall the number of encoded bits for $\delta_{a}^{T}$ is $k_{a}^{T}$, then $\dim(\text{im}(\delta_{a}))=n_{a}-k_{a}^{T}$, it is feasible to select a minimum of $k_{a}^{T}$ unit vectors $v_{i}$, ensuring that any combination of these $v_{i}$ remains external to the image of $\delta_{a}$. Specifically, for these $k_{a}^{T}$ unit vectors $v_{i}$, there does not exist any vector $x$ such that
$\sum_{i=1}^{k_{a}^{T}}a_{i}v_{i}=\delta_{a}x.$ (77)
[…]
Applying the same logic to operators $L_{Z}^{2}$ and $L_{Z}^{3}$ defined in Eq. (LABEL:3dzlogicaloperators), we can demonstrate that there exist $k_{a}k_{b}^{T}k_{c}$ distinct $L_{Z}^{2,i,j,k}$ and $k_{a}k_{b}k_{c}^{T}$ distinct $L_{Z}^{3,i,j,k}$ that serve as logical Z operators of the quantum code.
This completely characterizes all $k_{a}^{T}k_{b}k_{c}+k_{a}k_{b}^{T}k_{c}+k_{a}k_{b}k_{c}^{T}$ logical Z operators of the 3D hypergraph product codes. We refer to these as elementary canonical logical Z operators.
[…]
$L_{X}^{1}=\sum_{i,j,k}\alpha_{ijk}^{\prime}\left(\begin{array}[]{c}L_{A}^{i^{\prime}}\otimes w_{j}\otimes u_{k}\\ 0_{n_{a}r_{b}n_{c}}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right),$ (81)
$L_{X}^{2}=\sum_{i,j,k}\beta_{ijk}^{\prime}\left(\begin{array}[]{c}0_{r_{a}n_{b}n_{c}}\\ v_{i}\otimes L_{B}^{j^{\prime}}\otimes u_{k}\\ 0_{n_{a}n_{b}n_{c}}\end{array}\right),$ (85)
$L_{X}^{3}=\sum_{i,j,k}\gamma_{ijk}^{\prime}\left(\begin{array}[]{c}0_{r_{a}n_{b}n_{c}}\\ 0_{n_{a}r_{b}n_{c}}\\ v_{i}\otimes w_{j}\otimes L_{C}^{k^{\prime}}\end{array}\right).$ (89)
Here $v_{i},w_{j}$, and $u_{k}$ are unit vectors of length $n_{a},n_{b}$ and $n_{c}$ respectively. Moreover, $L_{A}^{i^{\prime}},L_{B}^{j^{\prime}}$ and $L_{C}^{k^{\prime}}$ are logical operators of the code defined by the parity check matrix $\delta_{a}^{T},\delta_{b}^{T}$, and $\delta_{c}^{T}$ respectively.
For $L_{X}^{1}$, $H_{Z}L_{X}^{1}=0$ implies that any Z-type stabilizer from $H_{Z}$ commutes with $L_{X}^{1}$. Meanwhile, any X-type stabilizer in $H_{X}$ also commutes with $L_{X}^{1}$ as they consist only of Pauli X operators. To identify logical X operators, we need to find Pauli X operators in the set $L_{X}^{1}$ that are not present in the image of $H_{X}^{T}$. We denote:
$L_{X}^{1,i,j,k}=\left(\begin{array}[]{c}L_{A}^{i^{\prime}}\otimes w_{j}\otimes u_{k}\\ 0_{n_{a}r_{b}n_{c}}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right).$ (94)
For $b_{j},c_{k}\in\{0,1\}$, suppose that the linear combination of these operators
$\sum_{jk}b_{j}c_{k}L_{X}^{1,i,j,k}=\left(\begin{array}[]{c}L_{A}^{i^{\prime}}\otimes\sum_{j}b_{j}w_{j}\otimes\sum_{k}c_{k}u_{k}\\ 0_{n_{a}r_{b}n_{c}}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right)$ (98)
is in the image of matrix $H_{X}^{T}$. Then, there exists vectors $y,z$ such that
$\sum_{jk}b_{j}c_{k}L_{X}^{1,i,j,k}$ $=$ $\left(\begin{array}[]{c}L_{A}^{i^{\prime}}\otimes\sum_{j}b_{j}w_{j}\otimes\sum_{k}c_{k}u_{k}\\ 0_{n_{a}r_{b}n_{c}}\\ 0_{n_{a}n_{b}r_{c}}\end{array}\right)$ (102)
$=$ $H_{X}^{T}(L_{A}^{i^{\prime}}\otimes y\otimes z)$ (103)
Recall that $k_{b}^{T}$ and $k_{c}^{T}$ represent the number of encoded bits for $\delta_{b}^{T}$ and $\delta_{c}^{T}$ respectively. Then with $\dim(\text{im}(\delta_{b}^{T}))=n_{b}-k_{b}$ and $\dim(\text{im}(\delta_{c}^{T}))=n_{c}-k_{c}$, one can select at least $k_{b}$ unit vectors $w_{j}$ and $k_{c}$ unit vectors $u_{k}$ such that any combination of these vectors remains outside the image of $\delta_{b}^{T}$ and $\delta_{c}^{T}$ respectively.
[…]
V.3 Energy barrier of Z logical operators
We now analyze the energy barrier for logical Z operators in 3D hypergraph product codes. First, we employ a strip argument to establish an upper bound. Then, using Lemma 11, we derive a lower bound for the energy barrier for logical Z operators.
V.3.1 Upper bound: “strip” argument
The upper bound on the energy barrier can be obtained by considering a specific path from the identity to any logical operator. By choosing this logical operator to be an elementary canonical logical operator, and considering the “strip” argument mentioned in Sec. IV.2, one can obtain an upper bound.
For elementary canonical logical operator in the form of $L_{Z}^{1,i,j,k}$, with fixed $i,j$, and $k$, the logical operator takes the form as in the Eq. (60). Since $v_{i}$ is a unit vector, the effective logical operator simplifies to $L_{Z}^{1^{\prime},j,k}=L_{B}^{j}\otimes L_{C}^{k}$. This can be interpreted exactly as a logical operator of a tensor product code. One can construct the same kind of strip configuration when implementing these logical operators. Thus, the energy barrier of this logical operator is upper bounded by the strip argument as $\Delta(L_{Z}^{1^{\prime},j,k})\leqslant\min\left[E_{b}d_{c},E_{c}d_{b}\right]$, where $E_{b}$ and $E_{c}$ are the energy barriers of the logical operators $L_{B}^{j}$ and $L_{C}^{k}$, respectively, and $d_{b}$ and $d_{c}$ are their corresponding distances.
With the same argument, the energy barrier of the canonical logical Z operator $L_{Z}^{2,i,j,k}$ is upper bounded by $\min\left[E_{a}d_{c},E_{c}d_{a}\right]$, and the energy barrier of $L_{Z}^{3,i,j,k}$ is upper bounded by $\min\left[E_{a}d_{b},E_{b}d_{a}\right]$.
In summary, consider a 3D hypergraph product code $\mathcal{C}$ constructed from classical codes $\delta_{a}$, $\delta_{b}$, and $\delta_{c}$ with parameters $\left[n_{\ell},k_{\ell},d_{\ell},E_{\ell}\right]$, where $\ell\in\{a,b,c\}$. Based on the structure of the canonical logical operators we have derived, the energy barrier of the logical Z operators can be shown to have the following upper bound.
$\Delta(L_{Z})\leqslant\min\{d_{\ell}E_{m}:\ell,m\in\{a,b,c\},\ell\neq m\}.$ (104)
V.3.2 Lower bound
We now establish a lower bound for the energy barrier of Z-type logical operators by analyzing the structure of these canonical logical operators. Consider a 3D hypergraph product code $\mathcal{C}$ constructed from classical codes $\delta_{a}$, $\delta_{b}$, and $\delta_{c}$ with parameters $\left[n_{\ell},k_{\ell},d_{\ell},E_{\ell}\right]$, where $\ell\in\{a,b,c\}$. The parity check matrices of this quantum code are $H_{Z}=\partial_{0}^{T}$ and $H_{X}=\partial_{1}$, as defined in Eq. (52) and Eq. (53). Recall that the Z-type canonical logical operators $L_{Z}$ fall into three categories.
$L_{Z}^{1}=\sum_{i,j,k}\alpha_{ijk}((v_{i}\otimes L_{B}^{j}\otimes L_{C}^{k})^{T},0^{T},0^{T})^{T},$ (105)
$L_{Z}^{2}=\sum_{i,j,k}\beta_{ijk}(0^{T},(L_{A}^{i}\otimes w_{j}\otimes L_{C}^{k})^{T},0^{T})^{T},$
$L_{Z}^{3}=\sum_{i,j,k}\gamma_{ijk}(0^{T},0^{T},(L_{A}^{i}\otimes L_{B}^{j}\otimes u_{k})^{T})^{T}.$
where (i) $\delta_{a}L_{A}^{i}=\delta_{b}L_{B}^{j}=\delta_{c}L_{C}^{k}=0$ and (ii) $v_{i}\notin\operatorname{Im}\left(\delta_{a}\right)$, $w_{j}\notin\operatorname{Im}\left(\delta_{b}\right)$ and $u_{k}\notin\operatorname{Im}\left(\delta_{c}\right)$ are unit vectors. We note that $i\in\{1,\ldots,k_{a}^{T}\}$, $j\in\{1,\ldots,k_{b}^{T}\}$ and $k\in\{1,\ldots,k_{c}^{T}\}$ and that the set of logical operators expressible in this form is complete. A canonical logical operator is elementary if only one of the coefficients, $\alpha_{ijk}$ (also $\beta_{ijk}$ or $\gamma_{ijk}$), is equal to one.
The Tanner graph representation provides an intuitive framework that is well-suited for our analysis. Let $\mathcal{G}_{1}(V_{1},C_{1})$, $\mathcal{G}_{2}(V_{2},C_{2})$ and $\mathcal{G}_{3}(V_{3},C_{3})$ be the Tanner graphs of codes defined by $\delta_{a}$, $\delta_{b}$ and $\delta_{c}$, respectively. Here, $V_{1}$, $V_{2}$ and $V_{3}$ represent the set of bits, and $C_{1}$, $C_{2}$ and $C_{3}$ denote the set of checks. We use $\{v_{1}^{i}:i\in\{1,2,\cdots n_{a}\}\}$ (resp. $\{v_{2}^{j}:j\in\{1,2,\cdots n_{b}\}\}$, $\{v_{3}^{k}:j\in\{1,2,\cdots n_{c}\}\}$) to refer to bit vertices in $V_{1}$ (resp. $V_{2}$, $V_{3}$), and also as length $n_{a}$ (resp. $n_{b}$, $n_{c}$) unit vectors with the $i$th (resp. $j$th, $k$th) entry as $1$.
The 3D hypergraph product $\mathcal{G}_{1}\times\mathcal{G}_{2}\times\mathcal{G}_{3}$ is a bipartite graph with vertex set $V\cup C$, where
$V=V_{1}\otimes V_{2}\otimes C_{3}\cup V_{1}\otimes C_{2}\otimes V_{3}\cup C_{1}\otimes V_{2}\otimes V_{3}$ (106)
is the qubit set, and
$C$ $=$ $V_{1}\otimes V_{2}\otimes V_{3}\cup C_{1}\otimes C_{2}\otimes V_{3}\cup C_{1}\otimes V_{2}\otimes C_{3}$ (107)
$\cup V_{1}\otimes C_{2}\otimes C_{3}\cup C_{1}\otimes C_{2}\otimes C_{3}$
is the check and meta-check set. More specifically, $V_{1}\otimes V_{2}\otimes V_{3}$ is the set the Z type checks, $C_{1}\otimes C_{2}\otimes V_{3}\cup C_{1}\otimes V_{2}\otimes C_{3}\cup V_{1}\otimes C_{2}\otimes C_{3}$ represents the set of X checks, and $C_{1}\otimes C_{2}\otimes C_{3}$ is meta-checks on X checks.
The set of qubits can be partitioned into three subsets, $V_{1}\otimes V_{2}\otimes C_{3}$, $V_{1}\otimes C_{2}\otimes V_{3}$, and $C_{1}\otimes V_{2}\otimes V_{3}$. For $H_{X}=\partial_{1}$, there are three column blocks (see Eq. (53)). The first column block of $H_{X}$ acts on $C_{1}\otimes V_{2}\otimes V_{3}$, the second column block of $H_{X}$ acts on $V_{1}\otimes C_{2}\otimes V_{3}$, and the third column block acts on $V_{1}\otimes V_{2}\otimes C_{3}$. Moreover, the subset $C_{1}\otimes V_{2}\otimes V_{3}$ can be further partitioned into $n_{a}^{T}\equiv r_{a}$ subsets $\{c_{1}^{1}\otimes V_{2}\otimes V_{3},c_{1}^{2}\otimes V_{2}\otimes V_{3},\cdots,c_{1}^{r_{a}}\otimes V_{2}\otimes V_{3}\}$, where $c_{1}^{k}\otimes V_{2}\otimes V_{3}:=\{c_{1}^{k}\otimes y\otimes z:y\in V_{2},z\in V_{3}\}$ and $V_{2}=\{v_{2}^{1},\ldots,v_{2}^{n_{b}}\}$, $V_{3}=\{v_{3}^{1},\ldots,v_{3}^{n_{c}}\}$.
Thus, a Z-type Pauli operator can be expressed as a bit-string $z=\left({z^{(1)}}^{T},{z^{(2)}}^{T},{z^{(3)}}^{T}\right)^{T}$, where $z^{(1)}$ is supported on the qubit subset $C_{1}\otimes V_{2}\otimes V_{3}$ with vector space $\mathbb{F}_{2}^{r_{a}}\otimes\mathbb{F}_{2}^{n_{b}}\otimes\mathbb{F}_{2}^{n_{c}}$, $z^{(2)}$ is supported on the qubit set $V_{1}\otimes C_{2}\otimes V_{3}$ with vector space $\mathbb{F}_{2}^{n_{a}}\otimes\mathbb{F}_{2}^{r_{b}}\otimes\mathbb{F}_{2}^{n_{c}}$ and $z^{(3)}$ is supported on the qubit set $V_{1}\otimes V_{2}\otimes C_{3}$ with vector space $\mathbb{F}_{2}^{n_{a}}\otimes\mathbb{F}_{2}^{n_{b}}\otimes\mathbb{F}_{2}^{r_{c}}$. Because $z^{(1)}$, $z^{(2)}$, and $z^{(3)}$ act as a tensor product of three vector spaces, one can view them also as 3D matrices. (For instance, $v_{i}\otimes u_{j}\otimes w_{k}$ for unit vectors $v_{i}$, $u_{j}$ and $w_{k}$ can be viewed as a 3D matrix whose entry is $1$ on the $i$’th a-slice, the $j$’th b-slice and the $k$’th c-slice and zero elsewhere. We call this procedure vector reshaping. After reshaping, $z^{(1)}$, $z^{(2)}$ and $z^{(3)}$ become $Z^{(1)}$, $Z^{(2)}$ and $Z^{(3)}$ respectively. Here, $Z^{(1)}$ is an $r_{a}\times n_{b}\times n_{c}$ matrix, and the entry $Z^{(1)}_{i,j,k}$ is supported on the qubit $r_{1}^{i}\otimes v_{2}^{j}\otimes v_{3}^{k}$. Moreover, the $i$th a-slice $Z^{(1)}_{i}$ is supported on qubits $\{r_{1}^{i}\otimes v_{2}^{1}\otimes v_{3}^{1},r_{1}^{i}\otimes v_{2}^{2}\otimes v_{3}^{1},\cdots,r_{1}^{i}\otimes v_{2}^{n_{b}}\otimes v_{3}^{n_{c}}\}$, which we refer to as $r_{1}^{i}\otimes V_{2}\otimes V_{3}$. Similarly, the subset $V_{1}\otimes C_{2}\otimes V_{3}$ can be partitioned into $r_{b}$ b-slices, and $j$th b-slice is supported on the qubit subset $V_{1}\otimes r_{2}^{j}\otimes V_{3}$. Also, the subset $V_{1}\otimes V_{2}\otimes C_{3}$ can be partitioned into $r_{c}$ c-slices, and $k$th c-slice is supported on the qubit subset $V_{1}\otimes V_{2}\otimes r_{3}^{k}$ (defined similarly).
An elementary canonical logical-Z operator, for example, in the form $(v_{i}\otimes L_{B}^{j}\otimes L_{C}^{k},0,0)^{T}$, is supported on the subset $r_{1}^{i}\otimes V_{2}\otimes V_{3}$ if $v_{i}=r_{1}^{i}$. In matrix form, it is supported on the $i$th a-slice of $Z^{(1)}$. Fig. 2 provides an illustrative example of the 3D Toric code structure.
Figure 2: The 3D toric code can be structured as a cubic lattice (with periodic boundary conditions). From left to right, the figure illustrates its decomposition into a-slices, b-slices, and c-slices. The 3D Toric code encodes 3 logical qubits, with corresponding 3 logical Z operators. Each Z elementary canonical logical operator is supported on a single slice, with slices of the same color representing logical operators equivalent up to stabilizers. Generally, given any 3D hypergraph product code, one can create such a decomposition, although the structure within each slice will be more complicated and depends on the underlying classical codes.
We now prove a lower bound for the energy barrier of the canonical logical-Z operators, as shown in the following proposition.
Proposition 13.
For any nontrivial canonical logical-Z operator $L$,
$\Delta(L)\geq\min[d_{a},d_{b},d_{c}]$ (108)
We introduce two lemmas to prove this proposition. We use the following convention in the proof. For the subsequent proof, we adopt the following notation: a path $r=\{P_{0},P_{1},\ldots,P_{\ell}\}$ is said to be supported on a subset $U\subseteq V$ if $\mathrm{supp}(P_{i})\subseteq U$ for all $0\leq i\leq\ell$.
Lemma 14.
For any elementary canonical logical-Z operator $L$ supported on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$ (resp. $V_{1}\otimes r_{2}^{\beta}\otimes V_{3}$, $V_{1}\otimes V_{2}\otimes r_{3}^{\gamma}$), its energy barrier is attained by a path supported on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$ (resp. $V_{1}\otimes r_{2}^{\beta}\otimes V_{3}$, $V_{1}\otimes V_{2}\otimes r_{3}^{\gamma}$).
Lemma 15.
For any nontrivial canonical logical-Z operator $L$, the energy barrier $\Delta(L)$ is greater than or equal to the minimum energy barrier of the elementary canonical logical-Z operators.
Consider an elementary canonical logical-Z operator $L$ supported on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$. Suppose $\Delta(L)$ is given by a path $r$. The main idea behind the proof of Lemma 14 is to deform a general path $r$ into a new path $r^{\prime}$, supported only on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$, and the energy barrier of $r^{\prime}$ is not greater than that of the original path $r$. A similar argument can be applied to prove Lemma 15. The proofs closely extended from the arguments in the hypergraph product case zhao2024on. The details are provided in the appendix [see Sec IX.2 and  IX.3].
For the elementary canonical logical operator $L_{Z}^{1,\alpha}=((r_{1}^{\alpha}\otimes L_{B}^{j}\otimes L_{C}^{k})^{T},0^{T},0^{T})^{T}$, with fixed $\alpha$, $j$, and $k$. $L_{Z}^{1,\alpha}$ is supported on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$. To compute the energy barrier of $L_{Z}^{1,\alpha}$, consider a path $r=\{P_{0},P_{1},\cdots,P_{F}\}$. Then
$\Delta(L_{Z}^{1,\alpha})=\min\{\epsilon_{\max}(r):r\in w(0,L_{Z}^{1,\alpha})\}.$ (109)
Here, $P_{0}=(0,0,0)^{T}$ is a zero vector, and $P_{F}=((r_{1}^{\alpha}\otimes L_{B}^{j}\otimes L_{C}^{k})^{T},0^{T},0^{T})^{T}$. Note that $r$ is a path that walks throughout the set of qubits. Therefore, $P_{\ell}\in r$ could be represented as $P_{\ell}=((P_{\ell}^{1})^{T},(P_{\ell}^{2})^{T},(P_{\ell}^{3})^{T})^{T}$, where each component, $P_{\ell}^{1},P_{\ell}^{2}$, and $P_{\ell}^{3}$, could be nontrivial.
Lemma 14 implies that there is a path $r$ supported on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$ that attains the energy barrier of $\Delta(L)$. A path restricted to the first component means that for any $P_{\ell}\in r$, one can set $P_{\ell}^{2}=P_{\ell}^{3}=0$, the only nontrivial Paulis (nonzero binaries) are in $P_{\ell}^{1}$ with fixed $r_{1}^{\alpha}$. Furthermore, $r_{1}^{\alpha}$ is fixed means every $P_{\ell}^{1}$ has the form $r_{1}^{\alpha}\otimes y\otimes z$, where $y\in V_{2}$ and $z\in V_{3}$ are vectors of length $n_{b}$ and $n_{c}$ respectively.
Given a $P_{\ell}\in r$, the energy (the number of violated checks) of $P_{\ell}$ is given by $\epsilon(P_{\ell})=\mathrm{wt}\left(H_{X}P_{\ell}\right)$, has the form
[…]
By setting $P_{\ell}^{2}=P_{\ell}^{3}=0$, then
$\epsilon(P_{\ell})$ $=\mathrm{wt}\left(\left(\begin{array}[]{c}I_{n_{a}}\otimes\delta_{b}\otimes I_{n_{c}}\\ I_{n_{a}}\otimes I_{n_{b}}\otimes\delta_{c}\end{array}\right)P_{\ell}^{1}\right)$ (118)
$=\mathrm{wt}\left(I_{n_{a}}\otimes\delta_{b}\otimes I_{n_{c}}P_{\ell}^{1}\right)+\mathrm{wt}\left(I_{n_{a}}\otimes I_{n_{b}}\otimes\delta_{c}P_{\ell}^{1}\right).$
Furthermore, for fixed $r_{1}^{\alpha}$, because $r_{1}^{\alpha}$ is a unit vector, the part $r_{1}^{\alpha}$ can be discarded by applying the map $(r_{1}^{\alpha})^{T}\otimes I_{n_{b}}\otimes I_{n_{c}}$ to $P_{\ell}^{1}$. Thus
$\epsilon(P_{\ell})=\mathrm{wt}\left(\delta_{b}\otimes I_{n_{c}}P_{\ell}^{1^{\prime}}\right)+\mathrm{wt}\left(I_{n_{b}}\otimes\delta_{c}P_{\ell}^{1^{\prime}}\right),$ (119)
with $P_{\ell}^{1^{\prime}}=((r_{1}^{\alpha})^{T}\otimes I_{n_{b}}\otimes I_{n_{c}})P_{\ell}^{1}$.
In summary, to calculate the energy barrier of the Z-type logical operator $L_{Z}^{1,\alpha,j,k}=((r_{1}^{\alpha}\otimes L_{B}^{j}\otimes L_{C}^{k})^{T},0^{T},0^{T})^{T}$, one can consider a path $r\in\{Q_{0},Q_{1},\cdots,Q_{L}\}$ supported on qubit subset $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$, then
$\Delta(L_{Z}^{1,\alpha,j,k})=\min\{\epsilon_{\max}(r):r\in w(0,L_{B}^{j}\otimes L_{C}^{k})\}.$ (120)
where for each $r$,
$\epsilon_{\max}(r)=\max[\epsilon(Q_{0}),\epsilon(Q_{1}),\cdots,\epsilon(Q_{L})],$ (121)
and the energy cost of the step $Q_{\ell}$ is
$\epsilon(Q_{\ell})=\mathrm{wt}\left(\delta_{b}\otimes I_{n_{c}}Q_{\ell}\right)+\mathrm{wt}\left(I_{n_{b}}\otimes\delta_{c}Q_{\ell}\right).$ (122)
This scenario corresponds exactly to computing the energy barrier of a classical tensor product code (Eq. (42)), where the underlying classical codes are $\delta_{b}$ and $\delta_{c}$. The subset $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$ is the bit set of this classical tensor product code. The path for applying logical operator $L_{Z}^{1,\alpha,j,k}$ effectively forms a path for the logical operator $L_{B}\otimes L_{C}$ in the tensor product code constructed by $\delta_{b}$ and $\delta_{c}$.
According to Lemma 11, the energy barrier of $L_{B}\otimes L_{C}$ is lower bounded by $\min\left[d_{b},d_{c}\right]$. Thus the energy barrier of the Z-type logical operator $L_{Z}^{1,\alpha,j,k}$ satisfies $\Delta\left(L_{Z}^{1,\alpha,j,k}\right)\geq$ $\min\left[d_{b},d_{c}\right]$.
Following the same reasoning, the energy barrier for logical Z-type operators of the form $L_{Z}^{2,i,j,k}=$ $\left(0^{T},(L_{A}^{i}\otimes w_{j}\otimes L_{C}^{k})^{T},0^{T}\right)^{T}$ is lower bounded by $\min\left[d_{a},d_{c}\right]$, and the logical operators of the form $L_{Z}^{3,i,j,k}=\left(0^{T},0^{T},(L_{A}^{i}\otimes L_{B}^{j}\otimes u_{k})^{T}\right)^{T}$ have an energy barrier lower bounded by $\min\left[d_{a},d_{b}\right]$.
In summary, for any elementary canonical Z logical operator $L$ in the 3D hypergraph product code, the energy barrier is lower bounded by
$\Delta(L)\geqslant\min\left[d_{a},d_{b},d_{c}\right].$ (123)
With Lemma 15, we can establish that for any canonical Z-type logical operator (not just elementary ones), the energy barrier is lower bounded by $\min\left[d_{a},d_{b},d_{c}\right]$. Moreover, results from Ref. zhao2024on demonstrate that for any logical operator $L$ and stabilizer $S$ with $\Delta(L)\geqslant\Delta(S)$, we have $\Delta(L)=\Delta(LS)$. Therefore, with the LDPC property yielding $\Delta(S)=O(1)$ for any $S$, and in cases where $d_{a},d_{b},d_{c}\geqslant O(1)$, this lower bound applies to all logical Z operators, not just the canonical ones.
Formally, we have the following theorem.
Theorem 16.
Let $\Delta(L_{Z})$ be the energy barrier of the Z logical operators of the 3D hypergraph product code constructed from classical codes $\delta_{a}$, $\delta_{b}$ and $\delta_{c}$ with parameters $n_{\ell},k_{\ell},d_{\ell},E_{\ell}]$ for $\ell\in\{a,b,c\}$. If the resulting quantum code is LDPC with sparsity parameters $w_{c},w_{q}$, then in cases where $d_{a},d_{b},d_{c}\geqslant w_{c}w_{q}$, we have
$\Delta(L_{Z})\geqslant\min[d_{a},d_{b},d_{c}].$ (124)
Similarly, with Lemma 10, we can conclude that the energy barrier of 3D LDPC hypergraph product codes is lower bounded by the energy barriers of the underlying classical codes:
$\Delta(L_{Z})\geqslant\min[E_{a},E_{b},E_{c}].$ (125)
Moreover, if Conjecture 12 is true, we have
$\Delta(L_{Z})\geqslant\min\{d_{\ell}E_{m}:\ell,m\in\{a,b,c\},\ell\neq m\}.$ (126)
This lower bound matches the upper bound obtained using the strip argument, effectively providing a tight bound.
V.4 Energy barrier of X logical operators
The canonical X logical operators have the form
$L_{X}^{1}=\sum_{i,j,k}\alpha_{ijk}^{\prime}((L_{A}^{i^{\prime}}\otimes w_{j}\otimes u_{k})^{T},0_{n_{a}r_{b}n_{c}}^{T},0_{n_{a}n_{b}r_{c}}^{T})^{T},$
$L_{X}^{2}=\sum_{i,j,k}\beta_{ijk}^{\prime}(0_{r_{a}n_{b}n_{c}}^{T},(v_{i}\otimes L_{B}^{j^{\prime}}\otimes u_{k})^{T},0_{n_{a}n_{b}n_{c}}^{T})^{T},$
$L_{X}^{3}=\sum_{i,j,k}\gamma_{ijk}^{\prime}(0_{r_{a}n_{b}n_{c}}^{T},0_{n_{a}r_{b}n_{c}}^{T},(v_{i}\otimes w_{j}\otimes L_{C}^{k^{\prime}})^{T})^{T},$
where (i) $\delta_{a}^{T}L_{A}^{i^{\prime}}=\delta_{b}^{T}L_{B}^{j^{\prime}}=\delta_{c}^{T}L_{C}^{k^{\prime}}=0$ and (ii) $v_{i}\notin\operatorname{Im}\left(\delta_{a}^{T}\right)$, $w_{j}\notin\operatorname{Im}\left(\delta_{b}^{T}\right)$ and $u_{k}\notin\operatorname{Im}\left(\delta_{c}^{T}\right)$ are unit vectors. We note that $i\in\{1,\ldots,k_{a}^{T}\}$, $j\in\{1,\ldots,k_{b}\}$ and $k\in\{1,\ldots,k_{c}\}$ and that the set of logical operators expressible in this form is complete. A canonical logical operator is elementary if only one of the coefficients, $\alpha_{ijk}^{\prime}$, $\beta_{ijk}^{\prime}$ or $\gamma_{ijk}^{\prime}$, is equal to one.
These elementary canonical logical X operators take the form of tensor products between one classical logical operator and two unit vectors, which resembles that of logical operators in hypergraph product codes, though in a higher dimension. By extending the reasoning previously applied to hypergraph product codes zhao2024on to this higher-dimensional case, in the LDPC regime, one can establish a lower bound for the energy barrier of logical X operators as
$\Delta(L_{X})\geqslant\min\left[E_{a}^{T},E_{b}^{T},E_{c}^{T}\right].$ (128)
The upper bound can be also derived from the structure of these elementary canonical logical X operators zhao2024on, which is also $\min[E_{a}^{T},E_{b}^{T},E_{c}^{T}]$. In the case where $\min\left[E_{a}^{T},E_{b}^{T},E_{c}^{T}\right]\geqslant w_{c}w_{q}$, with $w_{c},w_{q}$ are the sparsity parameters, we have
$\Delta(L_{X})=\min\left[E_{a}^{T},E_{b}^{T},E_{c}^{T}\right].$ (129)
VI Energy barrier of 4D hypergraph product codes
In 3D hypergraph product codes, the energy barrier of Z-type logical operators is lower bounded by the distance of the underlying classical codes. While X logical operators retain the energy barrier of these classical codes. When classical codes possess large distances but small energy barriers, only Z-type logical operators have improved energy barrier, potentially enabling single-shot. The X-type logical operators, lacking this enhanced protection, remain vulnerable. In this section, we show that 4D hypergraph product codes offer improved energy barrier for both X and Z logical operators.
The distinction between the 3D and 4D cases stems from structural differences in their elementary canonical logical operators. In 3D codes, Z logical operators take the form $L_{Z}=L_{c_{1}}\otimes L_{c_{2}}\otimes u_{k}$, while X logical operators are simply $L_{X}=v_{i}\otimes w_{j}\otimes L_{c_{3}}^{T}$, where $v_{i},w_{j}$, and $u_{k}$ are unit vectors, and $L_{c_{1}},L_{c_{2}},L_{c_{3}}$ are classical logical operators. In contrast, 4D codes exhibit a more balanced structure where both X and Z logical operators are tensor products of two classical logical operators (and two unit vectors, see Eq. 155 and Eq. 156), resulting in enhanced energy barriers for both types of logical operators.
Given four classical codes represented by parity check matrices $\delta_{a},\delta_{b},\delta_{c},\delta_{d}$, let $\left[n_{\ell},k_{\ell},d_{\ell},E_{\ell}\right]$ and $\left[n_{\ell}^{T},k_{\ell}^{T},d_{\ell}^{T},E_{\ell}^{T}\right]$ represent the parameters of the classical linear code defined by the parity check matrix $\delta_{\ell}$ and its transpose $\delta_{\ell}^{T}$. For convenience, we denote $r_{\ell}\equiv n_{\ell}^{T}$, where $\ell=\{a,b,c,d\}$. The 4D hypergraph product is defined by a length-$5$ chain complex
$C_{0}\xleftarrow{\partial_{0}}C_{1}\xleftarrow{\partial_{1}}C_{2}\xleftarrow{\partial_{2}}C_{3}\xleftarrow{\partial_{3}}C_{4},$ (130)
Boundary maps $\partial_{0},\partial_{1}$, $\partial_{2}$ and $\partial_{3}$ are
$\partial_{0}=\left(\begin{array}[]{c}\delta_{a}\otimes I_{n_{b}}\otimes I_{n_{c}}\otimes I_{n_{d}}\\ I_{n_{a}}\otimes\delta_{b}\otimes I_{n_{c}}\otimes I_{n_{d}}\\ I_{n_{a}}\otimes I_{n_{b}}\otimes\delta_{c}\otimes I_{n_{d}}\\ I_{n_{a}}\otimes I_{n_{b}}\otimes I_{n_{c}}\otimes\delta_{d}\end{array}\right),$ (135)
$\partial_{1}=\left(\begin{array}[]{cccc}I_{r_{a}}\otimes\delta_{b}\otimes I_{n_{c}}\otimes I_{n_{d}}&\delta_{a}\otimes I_{r_{b}}\otimes I_{n_{c}}\otimes I_{n_{d}}&0&0\\ I_{r_{a}}\otimes I_{n_{b}}\otimes\delta_{c}\otimes I_{n_{d}}&0&\delta_{a}\otimes I_{n_{b}}\otimes I_{r_{c}}\otimes I_{n_{d}}&0\\ I_{r_{a}}\otimes I_{n_{b}}\otimes I_{n_{c}}\otimes\delta_{d}&0&0&\delta_{a}\otimes I_{n_{b}}\otimes I_{n_{c}}\otimes I_{r_{d}}\\ 0&I_{n_{a}}\otimes I_{r_{b}}\otimes\delta_{c}\otimes I_{n_{d}}&I_{n_{a}}\otimes\delta_{b}\otimes I_{r_{c}}\otimes I_{n_{d}}&0\\ 0&I_{n_{a}}\otimes I_{r_{b}}\otimes I_{n_{c}}\otimes\delta_{d}&0&I_{n_{a}}\otimes\delta_{b}\otimes I_{n_{c}}\otimes I_{r_{d}}\\ 0&0&I_{n_{a}}\otimes I_{n_{b}}\otimes I_{r_{c}}\otimes\delta_{d}&I_{n_{a}}\otimes I_{n_{b}}\otimes\delta_{c}\otimes I_{r_{d}}\end{array}\right),$ (142)
$\partial_{2}=$
$\left({\begin{array}[]{cccccc}I_{r_{a}}\otimes I_{r_{b}}\otimes\delta_{c}\otimes I_{n_{d}}&I_{r_{a}}\otimes\delta_{b}\otimes I_{r_{c}}\otimes I_{n_{d}}&0&\delta_{a}\otimes I_{r_{b}}\otimes I_{r_{c}}\otimes I_{n_{d}}&0&0\\ I_{r_{a}}\otimes I_{r_{b}}\otimes I_{n_{c}}\otimes\delta_{d}&0&I_{r_{a}}\otimes\delta_{b}\otimes I_{n_{c}}\otimes I_{r_{d}}&0&\delta_{a}\otimes I_{r_{b}}\otimes I_{n_{c}}\otimes I_{r_{d}}&0\\ 0&I_{r_{a}}\otimes I_{n_{b}}\otimes I_{r_{c}}\otimes\delta_{d}&I_{r_{a}}\otimes I_{n_{b}}\otimes\delta_{c}\otimes I_{r_{d}}&0&0&\delta_{a}\otimes I_{n_{b}}\otimes I_{r_{c}}\otimes I_{r_{d}}\\ 0&0&0&I_{n_{a}}\otimes I_{r_{b}}\otimes I_{r_{c}}\otimes\delta_{d}&I_{n_{a}}\otimes I_{r_{b}}\otimes\delta_{c}\otimes I_{r_{d}}&I_{n_{a}}\otimes\delta_{b}\otimes I_{r_{c}}\otimes I_{r_{d}}\end{array}}\right),$ (148)
$\partial_{3}=\left(\begin{array}[]{cccc}I_{r_{a}}\otimes I_{r_{b}}\otimes I_{r_{c}}\otimes\delta_{d}&I_{r_{a}}\otimes I_{r_{b}}\otimes\delta_{c}\otimes I_{r_{d}}&I_{r_{a}}\otimes\delta_{b}\otimes I_{r_{c}}\otimes I_{r_{d}}&\delta_{a}\otimes I_{r_{b}}\otimes I_{r_{c}}\otimes I_{r_{d}}\end{array}\right).$ (151)
One can confirm that $\partial_{i}\partial_{i-1}=0\mod 2$ for $i\in{1,2,3}$. In this case, $H_{Z}=\partial_{1}^{T},H_{X}=\partial_{2}$ is used to construct a quantum CSS code $H_{(\delta_{a},\delta_{b},\delta_{c},\delta_{d})}$ with parity check matrix
$H_{\left(\delta_{a},\delta_{b},\delta_{c},\delta_{d}\right)}=\left(\begin{array}[]{cc}H_{X}&0\\ 0&H_{Z}\end{array}\right).$ (153)
Moreover, as $\partial_{3}\partial_{2}=\partial_{1}\partial_{0}=0$, the remaining matrices $M_{1}=\partial_{3}$, $M_{2}=\partial_{0}^{T}$ work as a meta-check for $H_{X}$ and $H_{Z}$, respectively. This means that any valid X-syndrome adheres to the constraints defined by $M_{1}$, and any valid Z-syndrome satisfies the constraints defined by $M_{2}$.
The parameters of the resulting quantum code $\mathcal{C}\left(\delta_{a},\delta_{b},\delta_{c},\delta_{d}\right)$ depend on the parameters of classical codes $\delta_{a},\delta_{b}$, $\delta_{c}$, and $\delta_{d}$, where
$n$ $=$ $n_{a}^{T}n_{b}^{T}n_{c}n_{d}+n_{a}^{T}n_{b}n_{c}^{T}n_{d}+n_{a}^{T}n_{b}n_{c}n_{d}^{T}+n_{a}n_{b}^{T}n_{c}^{T}n_{d}+n_{a}n_{b}^{T}n_{c}n_{d}^{T}+n_{a}n_{b}n_{c}^{T}n_{d}^{T},$
$k$ $=$ $k_{a}^{T}k_{b}^{T}k_{c}k_{d}+k_{a}^{T}k_{b}k_{c}^{T}k_{d}+k_{a}^{T}k_{b}k_{c}k_{d}^{T}+k_{a}k_{b}^{T}k_{c}^{T}k_{d}+k_{a}k_{b}^{T}k_{c}k_{d}^{T}+k_{a}k_{b}k_{c}^{T}k_{d}^{T},$
$d_{x}$ $=$ $\min\left[d_{a}^{T}d_{b}^{T},d_{c}^{T}d_{a}^{T},d_{a}^{T}d_{d}^{T},d_{c}^{T}d_{b}^{T},d_{b}^{T}d_{d}^{T},d_{c}^{T}d_{d}^{T}\right],$
$d_{z}$ $=$ $\min\left[d_{b}d_{c},d_{b}d_{d},d_{a}d_{d},d_{a}d_{c},d_{a}d_{b}\right].$ (154)
The Z-type canonical logical operators $L_{Z}$ can be categorized into the following six sets.
$L_{Z}^{1}=\sum_{i,j,k,l}a^{1}_{ijkl}((v_{i}\otimes w_{j}\otimes L_{C}^{k}\otimes L_{D}^{l})^{T},0^{T},0^{T},0^{T},0^{T},0^{T})^{T},$ (155)
$L_{Z}^{2}=\sum_{i,j,k,l}a^{2}_{ijkl}(0^{T},(v_{i}\otimes L_{B}^{j}\otimes u_{k}\otimes L_{D}^{l})^{T},0^{T},0^{T},0^{T},0^{T})^{T},$
$L_{Z}^{3}=\sum_{i,j,k,l}a^{3}_{ijkl}(0^{T},0^{T},(v_{i}\otimes L_{B}^{j}\otimes L_{C}^{k}\otimes t_{l})^{T},0^{T},0^{T},0^{T})^{T},$
$L_{Z}^{4}=\sum_{i,j,k,l}a^{4}_{ijkl}(0^{T},0^{T},0^{T},(L_{A}^{i}\otimes w_{j}\otimes u_{k}\otimes L_{D}^{l})^{T},0^{T},0^{T})^{T},$
$L_{Z}^{5}=\sum_{i,j,k,l}a^{5}_{ijkl}(0^{T},0^{T},0^{T},0^{T},(L_{A}^{i}\otimes w_{j}\otimes L_{C}^{k}\otimes t_{l})^{T},0^{T})^{T},$
$L_{Z}^{6}=\sum_{i,j,k,l}a^{6}_{ijkl}(0^{T},0^{T},0^{T},0^{T},0^{T},(L_{A}^{i}\otimes L_{B}^{j}\otimes u_{k}\otimes t_{l})^{T})^{T},$
where $v_{i},w_{j}$, $u_{k}$, and $t_{l}$ are unit vectors with length $r_{a},r_{b}$, $r_{c}$ and $r_{d}$, respectively. ${L_{A}^{i}},{L_{B}^{j}}$, $L_{C}^{k}$} and $L_{D}^{l}$} are logical operators of the code defined by the parity check matrix $\delta_{a},\delta_{b}$, $\delta_{c}$ and $\delta_{d}$, respectively. Given that $\delta_{a}L_{A}^{i}=\delta_{b}L_{B}^{j}=\delta_{c}L_{C}^{k}=\delta_{d}L_{D}^{l}=0$, one can confirm that $H_{X}L_{Z}^{l}=0$ for $l\in\{1,2,3,4,5,6\}$.
Similarly, the X-type canonical logical operators $L_{X}$ can be categorized into six sets.
$L_{X}^{1}=\sum_{i,j,k,l}b^{1}_{ijkl}(({L_{A}^{i}}^{\prime}\otimes{L_{B}^{j}}^{\prime}\otimes u_{k}\otimes t_{l})^{T},0^{T},0^{T},0^{T},0^{T},0^{T})^{T},$ (156)
$L_{X}^{2}=\sum_{i,j,k,l}b^{2}_{ijkl}(0^{T},({L_{A}^{i}}^{\prime}\otimes w^{j}\otimes{L_{C}^{k}}^{\prime}\otimes t_{l})^{T},0^{T},0^{T},0^{T},0^{T})^{T},$
$L_{X}^{3}=\sum_{i,j,k,l}b^{3}_{ijkl}(0^{T},0^{T},({L_{A}^{i}}^{\prime}\otimes w_{j}\otimes u_{k}\otimes{L_{D}^{l}}^{\prime})^{T},0^{T},0^{T},0^{T})^{T},$
$L_{X}^{4}=\sum_{i,j,k,l}b^{4}_{ijkl}(0^{T},0^{T},0^{T},(v_{i}\otimes{L_{B}^{j}}^{\prime}\otimes{L_{C}^{k}}^{\prime}\otimes t_{l})^{T},0^{T},0^{T})^{T},$
$L_{X}^{5}=\sum_{i,j,k,l}b^{5}_{ijkl}(0^{T},0^{T},0^{T},0^{T},(v_{i}\otimes{L_{B}^{j}}^{\prime}\otimes u_{k}\otimes{L_{D}^{l}}^{\prime})^{T},0^{T})^{T},$
$L_{X}^{6}=\sum_{i,j,k,l}b^{6}_{ijkl}(0^{T},0^{T},0^{T},0^{T},0^{T},(v_{i}\otimes w_{j}\otimes{L_{C}^{k}}^{\prime}\otimes{L_{D}^{l}}^{\prime})^{T})^{T}.$
where $v_{i},w_{j}$, $u_{k}$ and $t_{l}$ are unit vectors of length $n_{a},n_{b}$, $n_{c}$ and $n_{d}$, respectively. ${L_{A}^{i}}^{\prime},{L_{B}^{j}}^{\prime}$, $L_{C}^{k}$$^{\prime}$ and $L_{D}^{l}$$^{\prime}$ are logical operators of the code defined by the parity check matrix $\delta_{a}^{T},\delta_{b}^{T}$, $\delta_{c}^{T}$ and $\delta_{d}^{T}$, respectively.
Both X and Z elementary canonical logical operators have the structure of tensor product of two classical logical operators (and two unit vectors). Thus, in the LDPC regime, with Lemma 11, by applying the same reasoning used for logical Z operators in 3D hypergraph product codes but extended to four dimension, we can establish the following lower bounds on the energy barriers for logical Z and X operators in the 4D hypergraph product code.
$\Delta(L_{Z})$ $\geqslant$ $\min[d_{a},d_{b},d_{c},d_{d}],$
$\Delta(L_{X})$ $\geqslant$ $\min[d_{a}^{T},d_{b}^{T},d_{c}^{T},d_{d}^{T}].$ (157)
We state it formally in the following Theorem
Theorem 17.
Let $\Delta(L_{Z})$, $\Delta(L_{X})$ be the energy barrier of the Z, X logical operators of the 4D hypergraph product code constructed from classical codes $\delta_{a}$, $\delta_{b}$, $\delta_{c}$ and $\delta_{d}$, with parameters $n_{\ell},k_{\ell},d_{\ell},E_{\ell}]$ for $\ell\in\{a,b,c,d\}$. If the resulting quantum code is LDPC with sparsity parameters $w_{c},w_{q}$. When $d_{\ell},d_{\ell}^{T}\geqslant w_{c}w_{q}$ for $\ell\in\{a,b,c,d\}$, then we have
$\Delta(L_{Z})$ $\geqslant$ $\min[d_{a},d_{b},d_{c},d_{d}],$ (158)
$\Delta(L_{X})$ $\geqslant$ $\min[d_{a}^{T},d_{b}^{T},d_{c}^{T},d_{d}^{T}].$ (159)
Similarly, with Lemma 10, we have
$\Delta(L_{Z})$ $\geqslant$ $\min[E_{a},E_{b},E_{c},E_{d}],$ (160)
$\Delta(L_{X})$ $\geqslant$ $\min[E_{a}^{T},E_{b}^{T},E_{c}^{T},E_{d}^{T}].$ (161)
Moreover, if Conjecture 12 is true, we have
$\Delta(L_{Z})$ $\geqslant$ $\min\{d_{\ell}E_{m}:\ell,m\in\{a,b,c,d\},\ell\neq m\},$ (162)
$\Delta(L_{X})$ $\geqslant$ $\min\{d_{\ell}^{T}E_{m}^{T}:\ell,m\in\{a,b,c,d\},\ell\neq m\},$ (163)
which is tight.
VII Conclusion and outlook
In this work, we first gave a introduction to the confinement, soundness, and expansion properties of codes. We also discussed the relationships among these properties. Specifically, we showed the equivalence between left-expansion and linear confinement, and demonstrated formally that confinement property inherently induces a lower bound on energy barrier.
We then established a lower bound for the energy barrier of tensor product codes in Lemma 11, showing its relation to the distance of the underlying codes. Using this lemma, we proved that for 3D LDPC hypergraph product codes constructed from classical codes $\delta_{a}$, $\delta_{b}$, and $\delta_{c}$ with parameters $\left[n_{\ell},k_{\ell},d_{\ell},E_{\ell}\right]$, where $\ell\in{a,b,c}$, the energy barrier of logical Z operators is lower-bounded by $\Omega(d_{\ell})$, which improves upon the $\Omega(d_{\ell}^{\frac{1}{3}})$ bound derived from the soundness and confinement property. For logical X operators, the energy barrier corresponds to that of the transpose codes of these classical codes, paralleling the behavior in standard hypergraph products. In the 4D case, we showed that X and Z logical operators have energy barriers lower-bounded by $\Omega(d_{\ell})$ and $\Omega(d_{\ell}^{T})$, respectively.
Several questions remain for further investigation. An immediate one concerns the mechanism behind energy barrier enhancement in HHGP codes. A possible explanation lies in the increased check redundancy introduced by these higher-dimensional structures. This raises follow-up questions: Is there a quantitative relationship between redundancy and energy barriers? What amount or structure of redundancy is sufficient to achieve a specified energy barrier?
The redundancy provided by higher-dimensional hypergraph products is sufficient for achieving soundness and confinement campbell2019theory; Quintavalle2021single. Several works have also demonstrated the effectiveness of redundant parity check sets for simultaneously handling measurement and qubit errors fujiwara2014ability; Ashikhmin2020quantum; Delfosse2022beyond. However, establishing a general framework for introducing redundancy that ensures both confinement and macroscopic energy barriers remains an open question.
In contrast, quantum expander codes exhibit confinement even without check redundancy. This inspires the thought to explore alternative methods of providing confinement beyond redundancy or expansion, such as symmetries sam2020symmetry; charles2023self. Furthermore, quantum expander codes display unique statistical properties that may intrigue researchers, as demonstrated in recent studies hong2024quantum; yin2024low; placke2024topological; deroeck2024ldpc.
Finally, quantum codes with macroscopic energy barriers are conjectured to support local decoders Leverrier_2015; Fawzi_2018; kubica2019cellular; niko2017local; Hastings2014decoding; dennis2002topological; herold2015cellular; Breuckmann2018Scalable; dauphinais2017fault; Pastawski2011quantum; vasmer2021cellular, based on intuition from statistical mechanics. In such systems, local cooling processes can reduce system energy, suggesting that for codes with macroscopic energy barriers, a simple iterative local process could function effectively as a decoder. However, the design of local decoders for higher-dimensional hypergraph product codes remains an open research question.
VIII Acknowledgments
I thank Qi Ye of Tsinghua University for providing the argument for the energy barrier of two-dimensional repetition code. I thank Isaac Kim for helpful discussions and feedback on an earlier version of this manuscript. I also thank Andrew Doherty, Xiaotong Ni for valuable discussions. The author acknowledges the financial support from Sydney Quantum Academy. This work was supported by the Australian Research Council Centre of Excellence for Engineered Quantum Systems (EQUS, CE170100009), and partially supported by NSFC (No.92476206).
IX Appendix
Throughout this appendix, we work in the field $\mathbb{F}_{2}$. Thus, all addition operations are modulo $2$ except for the computation of the weight of vectors or matrices, i.e., the function $\mathrm{wt}(\cdot)$.
IX.1 Vector reshaping for 3D matrices
Consider a basis $\mathcal{B}$ of the vector space $\mathbb{F}_{2}^{n_{1}}\otimes\mathbb{F}_{2}^{n_{2}}$:
$\mathcal{B}=\left\{a_{i}\otimes b_{j}\mid i=\{1,\ldots,n_{1}\}\text{ and }j=\{1,\ldots,n_{2}\}\right\}.$ (164)
Then any vector $v\in\mathbb{F}_{2}^{n_{1}}\otimes\mathbb{F}_{2}^{n_{2}}$ can be written as
$v=\sum_{a_{i}\otimes b_{j}\in\mathcal{B}}v_{ij}\left(a_{i}\otimes b_{j}\right),$ (165)
for some $v_{ij}\in\mathbb{F}_{2}$. We call the $n_{1}\times n_{2}$ matrix $V$ with entries $v_{ij}$ the reshaping of the vector $v$. By this definition, if $A,B$ are respectively $m_{1}\times n_{1}$ and $m_{2}\times n_{2}$ matrices, then
$(A\otimes B)v\Rightarrow AVB^{T}.$ (166)
Define $\mathrm{wt}(M)$ as the number of ones in the vector (or matrix) $M$. We have $\mathrm{wt}((A\otimes B)v)=\mathrm{wt}(AVB^{T})$.
Now consider a vector $v$ lies in the tensor product space $\mathbb{F}_{2}^{n_{1}}\otimes\mathbb{F}_{2}^{n_{2}}\otimes\mathbb{F}_{2}^{n_{3}}$, we need to extend the reshaping and transformation concepts to accommodate this higher-order tensor structure.
The basis for the tensor product space $\mathbb{F}_{2}^{n_{1}}\otimes\mathbb{F}_{2}^{n_{2}}\otimes\mathbb{F}_{2}^{n_{3}}$ is:
$\mathcal{B}=\left\{a_{i}\otimes b_{j}\otimes c_{k}\right\},$ (167)
for $i=\{1,\ldots,n_{1}\};j=\{1,\ldots,n_{2}\};k=\{1,\ldots,n_{3}\}$.
Thus, any vector $v$ in this space can be written as
$v=\sum_{i=1}^{n_{1}}\sum_{j=1}^{n_{2}}\sum_{k=1}^{n_{3}}v_{ijk}\left(a_{i}\otimes b_{j}\otimes c_{k}\right),$ (168)
where $v_{ijk}\in\mathbb{F}_{2}$. Here, the coefficients $v_{ijk}$ form a 3D matrix $V$ with shape $n_{1}\times n_{2}\times n_{3}$.
To analyze the action of a Kronecker product transformation, say $(A\otimes B\otimes C)$, we extend the idea of reshaping $v$: The vector $v$ can be interpreted as a 3D matrix $V$ with entries $v_{ijk}$. Let $A,B$, and $C$ be matrices of dimensions $m_{1}\times n_{1},m_{2}\times n_{2}$, and $m_{3}\times n_{3}$, respectively. The action of $A\otimes B\otimes C$ on $v$ can be expressed as
$(A\otimes B\otimes C)v\equiv AVB^{T}C^{T}.$ (169)
Here, $V$ is treated as a 3D matrix, and the transformation applies $A,B$, and $C$ along the first, second, and third modes of $V$, respectively. The resulting transformed 3D matrix $AVB^{T}C^{T}$ is a new 3D matrix of shape $m_{1}\times m_{2}\times m_{3}$.
The weight of the 3D matrix $V$, denoted wt $(V)$, counts the total number of ones in $V$. The Kronecker product transformation preserves the weight
$\mathrm{wt}((A\otimes B\otimes C)v)=\mathrm{wt}\left(AVB^{T}C^{T}\right).$ (170)
IX.2 Proof of Lemma 14
Recall that a Z-type Pauli error of the 3D hypergraph product code can be expressed as a bit-string $z=\left(z^{(1)},z^{(2)},z^{(3)}\right)^{T}$. The corresponding energy $\epsilon(z)=\text{wt}(H_{X}z)$ is
$\epsilon(z)$ $=$ $\text{wt}\left((I_{r_{a}}\otimes\delta_{b}\otimes I_{n_{c}})z^{(1)}+(\delta_{a}\otimes I_{r_{b}}\otimes I_{n_{c}})z^{(2)}\right)$
$+\text{wt}\left((I_{r_{a}}\otimes I_{n_{b}}\otimes\delta_{c})z^{(1)}+(\delta_{a}\otimes I_{n_{b}}\otimes I_{r_{c}})z^{(3)}\right)$
$+\text{wt}\left((I_{n_{a}}\otimes I_{r_{b}}\otimes\delta_{c})z^{(2)}+(I_{n_{a}}\otimes\delta_{b}\otimes I_{r_{c}})z^{(3)}\right).$
By applying vector reshaping, the energy can be written as
$\epsilon(z)$ $=$ $\text{wt}\left(Z^{(1)}\delta_{b}^{T}+\delta_{a}Z^{(2)}\right)+\text{wt}\left(Z^{(1)}\delta_{c}^{T}+\delta_{a}Z^{(3)}\right)$ (172)
$+\text{wt}\left(Z^{(2)}\delta_{c}^{T}+Z^{(3)}\delta_{b}^{T}\right),$
where $Z^{(1)}$, $Z^{(2)}$ and $Z^{(3)}$ are the 3D matrices reshaped from $z^{(1)}$, $z^{(2)}$, and $z^{(3)}$, respectively. The $j$th a-slice of $Z^{(1)}$ is supported on qubit subset $r_{1}^{j}\otimes V_{2}\otimes V_{3}$.
Using Eq. (172), we aim to prove a lower bound on the energy $\epsilon(z)$ [Lemma 18]. To that end, we shall use the following convention. Let $L_{a}$ be a nontrivial codeword of $\delta_{a}^{T}$.
The set of a-slices of $Z^{(1)}$ associated with the nonzero entries of $L_{a}$ will play an important role. We define the index set of such columns as $\mathcal{C}(L_{a})$:
$\mathcal{C}(L_{a}):=\{j:(L_{a})_{j}=1\}.$ (173)
Given a codeword $L_{a}$ of $\delta_{a}^{T}$, one can construct a matrix $Z^{1,s}$ from the 3D matrix $Z^{(1)}$ by summing all the a-slices in the set $\mathcal{C}(L_{a})$. More formally,
$Z^{1,s}_{jk}=\sum_{i:i\in\mathcal{C}(L_{a})}Z^{(1)}_{ijk}$, (174)
where the addition is modulo $2$. For example, if $L_{a}=110100$, we would sum a-slices $1,2$, and $4$. Using Eq. (174), we can deform an arbitrary path to a path consisting of Paulis only supported on subset $r_{1}^{i}\otimes V_{2}\otimes V_{3}$ for some ii [^1] The precise choice of $i$ does not matter; any $i\in\mathcal{C}(L_{a})$ would suffice. In particular, we can prove an inequality between the energy of the original Pauli and the deformed Pauli, proved in Lemma 18.
Lemma 18.
$\mathrm{wt}(Z^{1,s}\delta_{b}^{T})\leqslant\mathrm{wt}\left(Z^{(1)}\delta_{b}^{T}+\delta_{a}Z^{(2)}\right).$ (175)
Proof.
We prove this by contradiction. Consider the elements of $Z^{1,s}\delta_{b}^{T}$ and $Z^{(1)}\delta_{b}^{T}+\delta_{a}Z^{(2)}$. If $\mathrm{wt}(Z^{1,s}\delta_{b}^{T})>\mathrm{wt}\left(Z^{(1)}\delta_{b}^{T}+\delta_{a}Z^{(2)}\right)$, then there exists $j,k$ such that
$\mathrm{wt}((Z^{1,s}\delta_{b}^{T})_{j,k})>\mathrm{wt}\left((Z^{(1)}\delta_{b}^{T}+\delta_{a}Z^{(2)})_{j,k}\right).$ (176)
We will prove that Eq. (176) cannot be satisfied, thereby proving the claim.
Without loss of generality, consider the $j,k$ elements. Notice that $(Z^{1,s}\delta_{b}^{T})_{jk}$ is a bit, thus the weight of it is either 0 or 1. If $\mathrm{wt}((Z^{1,s}\delta_{b}^{T})_{jk})=0$, Eq. (176) cannot be satisfied. Therefore, we consider the $\mathrm{wt}((Z^{1,s}\delta_{b}^{T})_{jk})=1$ case.
If $\mathrm{wt}((Z^{1,s}\delta_{b}^{T})_{jk})=1$, $(Z^{(1)}\delta_{b}^{T})_{jk}$ must contain an odd number of ones on the a-slice in the set $\mathcal{C}(L_{a})$. Otherwise, we would have had $\mathrm{wt}((Z^{1,s}\delta_{b}^{T})_{jk})=0$, which is a contradiction. On the other hand, we remark that $(\delta_{a}Z^{(2)})_{jk}$ consists of an even number of ones on the a-slice set $\mathcal{C}(L_{a})$. To see why, note that each $jk$-line of $\delta_{a}Z^{(2)}$ is a linear combination of the checks in $\delta_{a}^{T}$. Because $L_{a}$ is a codeword of $\delta_{a}^{T}$, $(\delta_{a}Z^{(2)})_{jk}L_{a}=0$. Therefore, in the $\mathrm{wt}((Z^{1,s}\delta_{b}^{T})_{jk})=1$ case, the number of ones in $Z^{(1)}\delta_{b}^{T}+\delta_{a}Z^{(2)}$ on the $jk$-line and the a-slice in $\mathcal{C}(L_{a})$ is odd.
Thus we conclude $\mathrm{wt}\left((Z^{(1)}\delta_{b}^{T}+\delta_{a}Z^{(2)})_{jk}\right)\geq 1$. As such, Eq. (176) cannot be satisfied. This completes the proof. ∎
With the same logic, one can prove that
Lemma 19.
$\mathrm{wt}(Z^{1,s}\delta_{c}^{T})\leqslant\mathrm{wt}\left(Z^{(1)}\delta_{c}^{T}+\delta_{a}Z^{(3)}\right).$ (177)
Now we are in a position to prove Lemma 14. We do so by identifying a path $r^{\prime}=\left\{P_{0}^{\prime},P_{1}^{\prime},\cdots,P_{F}^{\prime}\right\}$ that is only supported on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$ while ensuring that $\epsilon_{\max}\left(r^{\prime}\right)\leqslant\epsilon_{\max}(r)$, for any path $r$ supported on the entire qubit set. Lemma 18 and Lemma 19 suggest a way to deform the path $r$ to the one supported on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$.
Without loss of generality, let $r=\left\{P_{0},P_{1},\cdots,P_{F}\right\}$ be a path that $\Delta(L)=\epsilon_{\max}(r)$, with $P_{0}=I$ and $P_{F}=L$. We consider a Pauli $P_{i}$ in the path $r$. It will be convenient to work in its binary representation, written as $(p^{(1)},p^{(2)},p^{(3)})^{T}$, where $p^{(1)}$$p^{(2)}$ and $p^{(3)}$ represent the Paulis supported on $C_{1}\otimes V_{2}\otimes V_{3}$, $V_{1}\otimes C_{2}\otimes V_{3}$ and $V_{1}\times V_{2}\otimes C_{3}$, respectively. First, we remove all the Paulis supported on $V_{1}\otimes C_{2}\otimes V_{3}$ and $V_{1}\times V_{2}\otimes C_{3}$ by setting $p^{(2)}$ and $p^{(3)}$ as the zero vector. Next, we apply the following transformations to $p^{(1)}$. We reshape $p^{(1)}$ and refer to the reshaped 3D matrix as $P^{(1)}$. Let $L_{a}$ be a codeword of $\delta_{a}^{T}$ such that $(L_{a})_{\alpha}=1$. Consider a set of a-slices in $P^{(1)}$ corresponding to the index set $\mathcal{C}(L_{a})$. Denoting each a-slice as $r_{1}^{k}$, where $k\in\mathcal{C}(L_{a})$, we update the a-slice $r_{1}^{\alpha}$ in the following way:
$r_{1}^{\alpha}\to r_{1}^{\alpha}+\sum_{k\in\mathcal{C}(L_{a})\setminus\{\alpha\}}r_{1}^{k}.$ (178)
Afterward, the other a-slices of $P^{(1)}$ are set to the zero matrices. This yields the deformed Pauli operator $P_{i}^{\prime}$.
By construction, the resulting $P_{i}^{\prime}$ is supported on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$. Note that $r^{\prime}$ is a valid path because $\mathrm{wt}(P_{i}^{\prime}P_{i+1}^{\prime})\leq 1$ for every $i$. Also, because $P_{F}^{\prime}=P_{F}=L$, $r^{\prime}$ is still a path for $L$. Moreover, because $\epsilon\left(P_{i}^{\prime}\right)\leqslant\epsilon\left(P_{i}\right)$ for all $i$ [Lemma 18 and  19] as
$\mathrm{wt}(Z^{1,s}\delta_{c}^{T})+\mathrm{wt}(Z^{1,s}\delta_{b}^{T})\leqslant\epsilon(z)=\mathrm{wt}\left(Z^{(1)}\delta_{b}^{T}+\delta_{a}Z^{(2)}\right)$
$+\mathrm{wt}\left(Z^{(1)}\delta_{c}^{T}+\delta_{a}Z^{(3)}\right)+\text{wt}\left(Z^{(2)}\delta_{c}^{T}+Z^{(3)}\delta_{b}^{T}\right),$
we have $\epsilon_{\max}\left(r^{\prime}\right)\leqslant\epsilon_{\max}(r)$. Both $r^{\prime}$ and $r$ are paths for $L$, by definition $\epsilon_{\max}\left(r^{\prime}\right)\geqslant\epsilon_{\max}(r)$, we conclude $\epsilon_{\max}\left(r^{\prime}\right)=\epsilon_{\max}(r)$. Thus, by deforming $r$, we obtained a new path $r^{\prime}$ supported on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$ that yields the energy barrier $\Delta(L)$.
This argument can be applied to prove similar lower bounds for logical operators on $V_{1}\otimes r_{2}^{\beta}\otimes V_{3}$ and $V_{1}\otimes V_{2}\otimes r_{3}^{\gamma}$. To conclude, for any elementary canonical logical operator $L$ supported on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$ (resp. $V_{1}\otimes r_{2}^{\beta}\otimes V_{3}$ and $V_{1}\otimes V_{2}\otimes r_{3}^{\gamma}$), their energy barrier can be given by a path supported on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$ (resp. $V_{1}\otimes r_{2}^{\beta}\otimes V_{3}$ and $V_{1}\otimes V_{2}\otimes r_{3}^{\gamma}$).
IX.3 Proof of Lemma 15
Any nontrivial canonical logical-Z operator $L$ belongs to one of the following categories:
  • Case 1: $L$ is supported solely on the qubit subset $C_{1}\otimes V_{2}\otimes V_{3}$.
  • Case 2: $L$ is supported solely on the qubit subset $V_{1}\otimes C_{2}\otimes V_{3}$.
  • Case 3: $L$ is supported solely on the qubit subset $V_{1}\otimes V_{2}\otimes C_{3}$.
  • Case 4: $L$ is supported on multiple subsets.
We will focus solely on Case 1. Case 2 and Case 3 can be analyzed similarly by considering subsets $V_{1}\otimes C_{2}\otimes V_{3}$ and $V_{1}\otimes V_{2}\otimes C_{3}$, while case 4 can be treated as Case 1, 2 or 3.
Without loss of generality, let the energy barrier of $L$ be attained by a path $r=\{P_{0},P_{1},\cdots,P_{F}\}$, with $P_{0}=I$ and $P_{F}=L$. Similar to the approach taken in Lemma 14, we aim to deform the path $r$ to the one supported on $r_{1}^{k}\otimes V_{2}\otimes V_{3}$ for some $k$, such that the energy barrier of the deformed path lower bounds that of the $r$.
The deformation works in the same way as in the proof of Lemma 14. We describe this procedure again for the readers’ convenience. Let $L_{a}$ be a nontrivial codeword of $\delta_{a}^{T}$ and $\mathcal{C}(L_{a})$ be its corresponding column index set. We consider a binary representation of a Pauli $P_{i}$, written as $(p^{(1)},p^{(2)},p^{(3)})^{T}$. As in the proof of Lemma 14, we remove the Paulis supported on $V_{1}\otimes C_{2}\otimes V_{3}$ and $V_{1}\otimes V_{2}\otimes C_{3}$ by setting $p^{(2)},p^{(3)}$ as the zero vector. Next, reshape $p^{(1)}$ into a 3D matrix $P^{(1)}$ and update its columns in the following way. Choose $\alpha\in\mathcal{C}(L_{a})$. This column is updated as Eq. (178). The other columns of $P^{(1)}$ are converted to zero vectors.
Thanks to Lemma 18 and  19, we obtain a new path $r^{\prime}=\{P_{0}^{\prime},P_{1}^{\prime},\cdots,P_{F}^{\prime}\}$ supported on $r_{1}^{\alpha}\otimes V_{2}\otimes V_{3}$ with the property $\epsilon_{\max}\left(r^{\prime}\right)\leqslant\epsilon_{\max}\left(r\right)$. Note that $P_{F}^{\prime}$, in the binary representation, is of the form $r_{1}^{\alpha}\otimes\bar{y}\otimes\bar{z}$, where $\bar{y},\bar{z}$ are codewords of $\delta_{b}$ and $\delta_{c}$ respectively. Therefore, $P_{F}^{\prime}$ is either a nontrivial elementary canonical logical operator or the identity. In the latter case, $P_{F}^{\prime}$ is the trivial codeword (zero vector) in the binary representation. Henceforth, we denote this as $L^{\prime}=P_{F}^{\prime}$.
If $L^{\prime}$ is nontrivial, we can use the relation between the energy barriers of $L$ and $L^{\prime}$:
$\Delta(L)=\epsilon_{\max}(r)\geqslant\epsilon_{\max}(r^{\prime})\geqslant\Delta(L^{\prime}).$ (180)
Because $L^{\prime}$ is an elementary logical operator, $\Delta(L)$ is greater or equal to the minimum energy barrier of elementary canonical logical operators. Thus, if $L^{\prime}$ is nontrivial, the proof follows immediately.
If $L^{\prime}$ is an identity, the above argument does not work. Fortunately, it turns out that for any $L^{\prime}$, one can choose $L_{a}$ (the codeword of $\delta_{a}^{T}$ used in the current proof) such that $L^{\prime}$ is not an identity.
Without loss of generality, consider a canonical logical-Z operator $L$, expressed as
$L=\sum_{i,j,k}\alpha_{ijk}(x_{i}\otimes\bar{y}_{j}\otimes\bar{z}_{k},0,0)^{T},$ (181)
where (i) $\delta_{b}\bar{y}_{j}=\delta_{c}\bar{z}_{k}=0$ and (ii) $x_{i}\notin\operatorname{Im}\left(\delta_{a}\right)$ are unit vectors. If a given path ends with $L$, its deformation (using Eq. (178)) yields the following operator $L^{\prime}$:
$L^{\prime}=\sum_{j,k}c_{jk}(x_{\alpha}\otimes\bar{y}_{j}\otimes\bar{z}_{k},0,0)^{T},$ (182)
where $\alpha\in\mathcal{C}(L_{a})$ and $c_{jk}$ is defined as
$c_{jk}:=\sum_{i\in\mathcal{C}(L_{a})}\alpha_{ikj}.$ (183)
Note that $L^{\prime}$ is trivial if and only if $c_{jk}=0$ for all $j,k$. Therefore, we aim to prove that there exists a choice of $L_{a}$ such that $c_{jk}=1$ for at least one choice of $j,k$.
Let us prove the contrapositive. Suppose $c_{jk}=0$ for all $k$, for any choice of $L_{a}$. Consider the following vector:
$u_{jk}:=\sum_{i}\alpha_{ijk}x_{i}.$ (184)
Note that $c_{jk}=u_{jk}^{T}L_{a}$. By our assumption, $c_{jk}=0$ for any choice of $L_{a}$, so the inner product of $u_{jk}$ with any codeword of $\delta_{a}^{T}$ must be zero. On the other hand, $u_{jk}$, if it is nonzero, must lie outside of $\text{Im}(\delta_{a})$ by the definition of the $x_{i}$. Thus, $u_{jk}$ is not an element of the row space of $\delta_{a}^{T}$. However, this is a contradiction for the following reasons. For a linear code, let $H$ and $G$ be the parity check matrix and the generator matrix. Then $v^{T}G=0$ if and only if $v$ is a vector in the row space of $H$. In our setup, if $c_{jk}=0$ for any $L_{a}$, then $u_{jk}^{T}G=0$. This implies that $u_{jk}$ must be in the row space of $\delta_{a}^{T}$, which contradicts the fact that it lies outside of $\text{Im}(\delta_{a})$. To conclude, there must be at least one choice of $k,j$ such that $c_{jk}=1$. Thus, there is always a choice of $L_{a}$ such that $L^{\prime}$ is not an identity, thereby proving the claim.
Case 2 and Case 3 can be analyzed similarly to Case 1 by considering subset $V_{1}\otimes C_{2}\otimes V_{3}$ and $V_{1}\otimes V_{2}\otimes C_{3}$. For Case 4, one can treat it just as Case 1, Case 2, or Case 3. For example, when treating it as Case 1, the logical operator $L$ has a nontrivial part in the subset $C_{1}\otimes V_{2}\otimes V_{3}$. One can prove there exists a codeword $L_{a}$ of $\delta_{a}^{T}$, such that after the deformation, the resulting $L^{\prime}$ is a nontrivial elementary canonical logical operator.
IX.4 Discussion on Conjecture 12
Let $L_{a}$ and $L_{b}$ denote the logical operators of $\delta_{a}$ and $\delta_{b}$, with distances $d_{a}$, $d_{b}$ and energy barriers $E_{a}$, $E_{b}$, respectively. Our goal is to determine the energy barrier of the operator $L_{c}=L_{a}\otimes L_{b}$, which serves as a logical operator for $\delta_{c}$.
To compute the energy barrier of $L_{c}$, we consider a path $r=\{P_{0},P_{1},\cdots,P_{F}\}$ with $P_{0}=I$ and $P_{F}=L_{c}$. Then
$\Delta(L_{c})=\min\{\epsilon_{\max}(r):r\in w(0,L_{c})\}.$ (185)
Note that $r$ is a path that walks on the bit of code $\delta_{c}$. Given a $P_{\ell}\in r$, the energy (the number of violated checks) of $P_{\ell}$ is
$\epsilon(P_{\ell})$ $=$ $\text{wt}(\delta_{c}P_{\ell})$ (188)
$=$ $\text{wt}\left(\left(\begin{array}[]{c}\delta_{a}\otimes I_{n_{b}}\\ I_{n_{a}}\otimes\delta_{b}\end{array}\right)P_{\ell}\right)$
$=$ $\text{wt}((\delta_{a}\otimes I_{n_{b}})P_{\ell})+\text{wt}((I_{n_{a}}\otimes\delta_{b})P_{\ell}).$ (189)
To proceed, one can express the identity matrices $I_{n_{a}}$ and $I_{n_{b}}$ as
$I_{n_{a}}=\sum_{j=1}^{n_{a}}w_{j}w_{j}^{T},\quad I_{n_{b}}=\sum_{k=1}^{n_{b}}u_{k}u_{k}^{T}.$ (190)
Eq. (189) can then be further processed as
$\epsilon(P_{\ell})$ $=\mathrm{wt}\left(\delta_{a}\otimes I_{n_{b}}P_{\ell}\right)+\mathrm{wt}\left(I_{n_{a}}\otimes\delta_{b}P_{\ell}\right)$ (191)
$=\mathrm{wt}\left(\delta_{a}\otimes\sum_{k=1}^{n_{b}}u_{k}u_{k}^{T}P_{\ell}\right)+\mathrm{wt}\left(\sum_{j=1}^{n_{a}}w_{j}w_{j}^{T}\otimes\delta_{b}P_{\ell}\right)$
$=\mathrm{wt}\left(\sum_{k=1}^{n_{a}}(\delta_{a}\otimes u_{k})(I_{n_{a}}\otimes u_{k}^{T})P_{\ell}\right)$
$\ \ +\mathrm{wt}\left(\sum_{j=1}^{n_{a}}(w_{j}\otimes\delta_{b})(w_{j}^{T}\otimes I_{n_{b}})P_{\ell}\right)$
$=\mathrm{wt}\left(\sum_{k=1}^{n_{b}}(\delta_{a}\otimes u_{k})P_{\ell}^{k}\right)+\mathrm{wt}\left(\sum_{j=1}^{n_{a}}(w_{j}\otimes\delta_{b})P_{\ell}^{j}\right)$
$=\sum_{k=1}^{n_{b}}\mathrm{wt}\left(\delta_{a}P_{\ell}^{k}\right)+\sum_{j=1}^{n_{a}}\mathrm{wt}\left(\delta_{b}P_{\ell}^{j}\right),$
where
$P_{\ell}^{k}=(I_{n_{a}}\otimes u_{k}^{T})P_{\ell},\quad P_{\ell}^{j}=(w_{j}^{T}\otimes I_{n_{b}})P_{\ell}.$ (192)
To calculate the energy barrier of $L_{c}$, we minimize $\epsilon_{\max}\left(P_{\ell}\right)$ over all possible $r$. Here, $\epsilon\left(P_{\ell}\right)$ is the sum of two components: $\sum_{k=1}^{n_{b}}\mathrm{wt}\left(\delta_{a}P_{\ell}^{k}\right)$ and $\sum_{j=1}^{n_{a}}\mathrm{wt}\left(\delta_{b}P_{\ell}^{j}\right)$. Although these two components depend on each other, we first consider minimizing them independently.
We can first minimize $\sum_{k=1}^{n_{b}}\mathrm{wt}\left(\delta_{a}P_{\ell}^{k}\right)$, followed by minimizing $\sum_{j=1}^{n_{a}}\mathrm{wt}\left(\delta_{b}P_{\ell}^{j}\right)$. To minimize $\sum_{k=1}^{n_{b}}\mathrm{wt}\left(\delta_{a}P_{\ell}^{k}\right)$, we apply $L_{a}$ along the path $r_{a}$, which corresponds to the energy barrier of $L_{a}$ for each $k$ sequentially. The maximum energy contribution for this part is $E_{a}$. Subsequently, we minimize $\sum_{j=1}^{n_{a}}\mathrm{wt}\left(\delta_{b}P_{\ell}^{j}\right)$ by setting the order of $k$ according to the path $r_{b}$ that gives the energy barrier of $L_{b}$. The highest energy contribution for this part is $d_{a}E_{b}$. In particular, when the second part reaches $d_{a}E_{b}$, the first part has an energy of 0. Therefore, this path results in an energy barrier of $\epsilon_{\max}(r)=d_{a}E_{b}$. Alternatively, we can reverse the process by first minimizing $\sum_{j=1}^{n_{a}}\mathrm{wt}\left(\delta_{b}P_{\ell}^{j}\right)$ and then $\sum_{k=1}^{n_{b}}\mathrm{wt}\left(\delta_{a}P_{\ell}^{k}\right)$. Following a similar analysis, this approach leads to an energy barrier of $\epsilon_{\max}\left(r^{\prime}\right)=d_{b}E_{a}$.
To complete the argument, one needs to prove that there is no path with an energy barrier lower than $\min\left[d_{a}E_{b},d_{b}E_{a}\right]$. We developed a program to verify the conjecture  [^2] Github repository: https://github.com/Guangqi-Phys/Energy-Barrier-of-tensor-product-codes.git. This program generates random codes with their logical operators, computes their energy barriers and distances, constructs the tensor product code, and calculates its energy barrier.
Computing the energy barrier of a general code is hard. Our implementation uses a best-first search algorithm that may not always identify the absolute minimum. In our tests across many random codes, most results support the conjecture. In the rare cases where the energy barrier of the tensor product code differs from our product-based estimate, the discrepancy is minor and can be attributed to computational limitations.
While our numerical simulations strongly support the conjecture, one needs to develop alternative combinatorial methods for a rigorous proof. We believe that solving this problem will not only strengthen our current findings but also have broader applications for tensor product codes, particularly given their close relationship to locally testable codes.